### v2

# Permutation twisted cohomology, remixed

Abstract.
We generalize Balmer and Gallauer’s (permutation) twisted cohomology ring, working toward an alternative method of deducing the Balmer spectrum of the derived category of permutation modules for any finite $p$-group. The construction comes equipped with a canonical comparison map from the Balmer spectrum to the homogeneous spectrum of the twisted cohomology ring. We show the comparison map is injective for any finite $p$-group and furthermore, an open immersion when the twisted cohomology ring is noetherian. For elementary abelian $p$-groups, the twisted cohomology ring coincides with Balmer and Gallauer’s original construction.
To perform this construction, we utilize endotrivial complexes (i.e. the invertible objects of the derived category of permutation modules) arising up to a shift from Bredon homology of representation spheres. This topological structure allows us to construct certain $p$-local isomorphisms, from which we build a refined open cover of the Balmer spectrum indexed by conjugacy classes of subgroups of $G$. Under this open cover, every endotrivial is isomorphic to a shift of the tensor unit in each localization, thus verifying that all endotrivials are line bundles. When the twisted cohomology ring is noetherian, this open cover endows the Balmer spectrum with Dirac scheme structure.
Key words and phrases:
permutation module, twisted cohomology, comparison map, Balmer spectrum, tensor-triangular geometry
2020 Mathematics Subject Classification:
18G80, 18G90, 18M05, 20C20, 20J05
Introduction
Let $k$ be a field of prime characteristic $p$ and $G$ be a finite group. Permutation $kG$-modules (modules which admit a $G$-stable $k$-basis), and their direct summands, $p$-permutation kGkG-modules, are perhaps the second easiest class of representations of $G$ over $k$ to define. In a wildly vast representation-theoretic cosmos of $kG$-modules, our $p$-permutation modules may seem tiny and insignificant, drowning in a sea of symmetries. However, our protagonists, or more precisely, the bounded homotopy category of $kG$-modules $\mathscr{K}(G):=\operatorname{K}_{b}(p\operatorname{-perm}(kG))$, exhibits significant control over the representation theory of $kG$ in general. For instance, if $G$ has an abelian Sylow $p$-subgroup $S$, Broué’s abelian defect group conjecture predicts a triangulated equivalence $\operatorname{K}_{b}(p\operatorname{-perm}(kGb_{0}))\cong\operatorname{K}_{b}(p\operatorname{-perm}(kN_{G}(S)c_{0}))$ where $b_{0}$ and $c_{0}$ denote the principal blocks of $kG$ and $kN_{G}(S)$ respectively. Perhaps more surprisingly, Balmer and Gallauer recently demonstrated that $\mathscr{K}(G)$ admits the bounded derived category $\operatorname{D}_{b}(\text{mod}(kG))$ as a Verdier quotient [BG23a], strengthening an unpublished result of Rouquier [Rou06]. Although these $p$-permutation are few, they are mighty; together, they recover the entire representation theory of $kG$.
If the reader is less representation-theoretically inclined, fear not! Our $p$-permutation category $\mathscr{K}(G)$ is the compact part of a “big” (i.e. rigidly-compactly generated) tt-category $\mathscr{D}(G):=\operatorname{DPerm}(G;k)$, the derived category of permutation modules. This category $\mathscr{D}(G)$ is equivalent to numerous big tt-categories arising from seemingly disparate areas of mathematics, such as:
  1. (a)
    the derived category of cohomological $k$-linear Mackey functors over $G$;
  2. (b)
    the homotopy category of modules over the constant Green functor $\operatorname{H}\underline{k}$ in genuine $G$-spectra;
  3. (c)
    the category of $k$-linear Artin motives generated by motives of intermediate fields in a Galois extension with Galois group $G$.
See [BG23b] for an overview of these correspondences; we remark Fuhrmann recently ascended equivalence (b) to the $\infty$-categorical level [Fuh25]. Hence any results concerning $\mathscr{D}(G)$ and $\mathscr{K}(G)$ translate over to the aforementioned categories and their compact parts respectively.
Continuing their foray into the land of permutation modules, Balmer-Gallauer deduced the Balmer spectrum $\operatorname{Spc}(\mathscr{K}(G))$ of $\mathscr{K}(G)$ in [BG25], a landmark achievement. Their construction of permutation twisted cohomology, a modification of the usual cohomology ring, plays a key role in this classification. The authors consider morphisms from the tensor unit to shifts of certain invertible objects (i.e. endotrivial complexes) corresponding to subgroups $N\triangleleft G$ of index $p$. The collection of these morphisms forms a multigraded ring $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$, the twisted cohomology ring, extending the usual notion of a cohomology ring of a triangulated category, and comes with a canonical comparison map,
$\operatorname{comp}_{G}\colon\operatorname{Spc}(\mathscr{K}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)),$
which was previously constructed in [Bal10a]. A twist of the cohomology ring is necessary - since we are considering a homotopy category, the usual cohomology ring $\operatorname{End}^{\scriptscriptstyle\bullet}_{\mathscr{K}(G)}(k)$ is a tad drab on its own. The miracle of this construction is:
Theorem.
[BG25, Theorem 10.5] Let $E$ be a finite elementary abelian $p$-group. The comparison map $\operatorname{comp}_{E}\colon\operatorname{Spc}(\mathscr{K}(E))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(E))$ identifies $\operatorname{Spc}(\mathscr{K}(E))$ with an open subspace of the homogeneous spectrum of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(E)$.
This remarkable fact, along with a colimit theorem [BG25, Theorem 11.10] reducing to the case of $G$ elementary abelian á la Quillen stratification, completes the description of $\operatorname{Spc}(\mathscr{K}(G))$ as a topological space, and as a Dirac scheme in the sense of [HP23]. As a set, $\operatorname{Spc}(\mathscr{K}(G))\cong\bigsqcup_{H\in s_{p}(G)/G}\operatorname{Spc}(\operatorname{D}_{b}(kG))$ ([BG25, Theorem 7.16]); we set $V_{G}:=\operatorname{Spc}(\operatorname{D}_{b}(kG))$. Hausmann and Schwede further considered the twisted cohomology ring for elementary abelian $2$-groups [HS25] (in the Mackey functor setting), which they entitled “Representation-graded Bredon homology.” The authors determined a minimal generating set and proved that for $p=2$, $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(E)$ is nilpotent-free.
Separately, the author deduced the Picard group of $\mathscr{K}(G)$, i.e. the group of endotrivials, in a sequence of papers [Mil24, Mil25b, Mil25a]. Significant parts of the classification were also completed by Bachmann in his dissertation [Bac16] in the context of Artin motives. When $G$ is a finite $p$-group, morally the endotrivials arise from Bredon homology of representation spheres, and therefore have additional topological structure. This was also observed by Yalçin in the context of $G$-Moore spaces [Yal17], where an analogous classification result was determined. Therefore, one obtains a canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K}(G))$ associated to real representations modulo Adams operations. Interestingly, for elementary abelian $p$-groups, this basis corresponds exactly to the endotrivials used by Balmer-Gallauer in their twisted cohomology ring. The natural question for one to ask is:
Question.
Can we recover the Balmer spectrum $\operatorname{Spc}(\mathscr{K}(G))$ by further twisting (or if you will, re-mixing) $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ by the endotrivial complexes arising (up to shifts) from representation spheres?
We consider twisting by shifts of complexes coming from genuine $\mathbb{R}G$-modules as opposed to virtual representations, as if we were to twist by all endotrivials, the graded ring becomes unmanageably large (c.f. [BG25, Remark 12.24]). The analogous comparison to make here is that the group cohomology ring $\operatorname{H}^{\scriptscriptstyle\bullet}(G;k)$ is noetherian and recovers the Balmer spectrum of the derived category $\operatorname{D}_{b}(kG)$, whereas the Tate cohomology ring is too large. We call such endotrivials effective, and their corresponding h-mark homomorphisms are monotonically decreasing superclass functions. Similarly, Hausmann-Schwede call the corresponding “quadrant” of twisted cohomology associated to twists by effective endotrivials the effective cone.
Denote the subset of the canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K}(G))$ arising from nontrivial irreducible real representations of $G$ by $\mathcal{B}(G)$. Let $\mathbb{N}^{\mathcal{B}(G)}=\{q\colon\mathcal{B}(G)\to\mathbb{N}\}$ be the monoid of twists, i.e. tuples of non-negative integers indexed by the set $\mathcal{B}(G)$. The monoid of twists is equivalently nothing more than the submonoid of $\operatorname{Pic}(\mathscr{K}(G))$ generated by $\mathcal{B}(G)$. Then the $(\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)})$-graded ring
$\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)=\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G;k):=\bigoplus_{s\in\mathbb{Z}}\bigoplus_{q\in\mathbb{N}^{\mathcal{B}(G)}}\operatorname{Hom}_{\mathscr{K}(G)}\left(k,\bigotimes_{C\in\mathcal{B}(G)}C^{\otimes q(C)}[s]\right)$
is the (re-)twisted cohomology ring; see Definition 5.1. We note that only non-positive shifts $s\leq 0$ play a nontrivial role.
We take the first steps towards answering this question for finite $p$-groups by proving injectivity of the comparison map. The $p$-group case is perhaps the most fundamental: given a finite group $G$ with $p$-Sylow $S$, one can determine $\operatorname{Spc}(\mathscr{K}(G))$ from $\operatorname{Spc}(\mathscr{K}(S))$ via restriction.
Theorem A. (Theorem 6.1).
Let $G$ be a finite $p$-group. The comparison map $\operatorname{Spc}(\mathscr{K}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ is injective.
Generally, injectivity of the comparison map can be a harder condition to prove (if it holds at all, which it frequently may not!) than surjectivity (after a categorical localization), which general tt-geometry usually resolves. Circling back to Balmer-Gallauer’s original construction, another feature of $\operatorname{Spc}(\mathscr{K}(E))$ for $E$ elementary abelian is a spectral open cover $\{U(H)\}_{H\leq E}$ of $\operatorname{Spc}(\mathscr{K}(E))$ indexed by subgroups of $E$. This cover has two remarkable features:
  1. (a)
    Over each open $U(H)$, all endotrivials for $E$ are trivial, i.e. isomorphic to a shift of the tensor unit in the localization $\mathscr{L}(H):=\mathscr{K}(E)|_{U(H)}$;
  2. (b)
    There is a homeomorphism between each $U(H)$ and the homogeneous spectrum of the $\mathbb{Z}$-graded endomorphism ring $\operatorname{End}_{\mathscr{L}(H)}^{\scriptscriptstyle\bullet}(k)$ in the localization $\mathscr{L}(H)$. In particular, $\operatorname{Spc}(\mathscr{K}(E))$ is a Dirac scheme.
In addition, under this open cover, there is an identification $\mathscr{K}(E)|_{U(1)}\cong\operatorname{D}_{b}(kE)$ - in other words, $U(1)$ is nothing more than the “cohomological open,” corresponding to the image under $\operatorname{Spc}$ of the localization $\mathscr{K}(E)\twoheadrightarrow\operatorname{D}_{b}(kE)$. We construct a refined open cover with these same properties.
Theorem B.
Let $G$ be a finite $p$-group. Then there exists an open cover
$\{U(H)\}_{H\in s_{p}(G)/G}$ (Construction 4.5) indexed by conjugacy classes of subgroups for which the following holds.
  1. (a)
    Each open $U(H)$ contains a unique closed point of $\operatorname{Spc}(\mathscr{K})$, $\mathfrak{m}_{H}$, corresponding to the unique closed point of $V_{{G}/\!\!/{H}}$; (Corollary 4.11)
  2. (b)
    The open $U(1)$ is equivalently the cohomological open $V_{G}=\operatorname{Spc}(\operatorname{D}_{b}(kG))$ of $\operatorname{Spc}(\mathscr{K}(G))$. In other words, we have an equivalence of categories $\mathscr{K}(G)|_{U(1)}\cong\operatorname{D}_{b}(kG)$; (Theorem 4.17)
  3. (c)
    Every endotrivial $C$ is a line bundle, i.e. over each open $U(H)$ is isomorphic to a shift of the tensor unit in the localization $\mathscr{K}(G)|_{U(H)}$. In particular, we have an isomorphism $C\cong k[h_{C}(H)]$ where $h_{C}$ denotes the unique h-mark function associated to $C$. (Theorem 4.13)
Finally, if $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian, we completely recover the Balmer spectrum.
Theorem C. (Corollaries 6.4, 6.5).
If $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian, the comparison map is an open immersion and for each subgroup $H\leq G$, restricts to a local homeomorphism
$\operatorname{comp}_{\mathcal{L}_{G}(H)}:U(H)\to\operatorname{Spec}^{h}(\operatorname{End}_{\mathscr{L}_{G}(H)}^{\scriptscriptstyle\bullet}(k)).$
In this case $(\operatorname{Spc}(\mathscr{K}(G)),\mathcal{O}_{G}^{\scriptscriptstyle\bullet})$ is a Dirac scheme.
To obtain these results, it is seemingly necessary to twist by all effective endotrivials (or at minimum, a full-rank subset); see for instance Remark 4.12. In this sense, we believe our construction for $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is the “correct” generalization, and not too big. The not-so-secret sauce to the above results is a generalization of Balmer-Gallauer’s maps $a_{N}$ and $b_{N}$ for any effective endotrivial. For any effective endotrivial $C$, we construct unique (up to scaling) maps $\iota_{C}^{H}$ for every subgroup $H\leq G$ for which $\Psi^{H}(\iota_{C}^{H})$ is a quasi-isomorphism. Proving these maps exist is rather technical and relies heavily on the topological origins of effective endotrivials. These maps may not exist for chain complexes not arising from chain complexes of free modules over the orbit category $\Gamma_{G}$ (in particular, non-effective endotrivials), see Remark 3.8.
Theorem D. (Theorem 4.2).
Let $G$ be a finite $p$-group. For every effective endotrivial complex (see Definition 2.1) and subgroup $H\leq G$, there exists a homomorphism $\iota^{H}_{C}:k\to C[h_{C}(H)]$ such that $\Psi^{H}(\iota^{H}_{C})$ is a quasi-isomorphism.
Note that there are numerous maps that satisfy the above property (see Remark 4.3); our maps $\iota^{H}_{C}$ satisfy additional technical conditions. Finally, we prove some technical results regarding these chain complexes. Although they are not strictly necessary for our constructions, they contribute to our understanding of effective endotrivials. These propositions may be known to experts, but we reprove them for completeness.
Theorem E. (Proposition 3.11, Corollary 3.12).
Let $G$ be a finite group. Suppose $C$ is a bounded chain complex of permutation $kG$-modules arising from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules. Then $C$ has a contractible direct summand if and only if $C^{?}$ does. In particular, if $C$ is an indecomposable effective endotrivial, then $C$ arises from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules.
Open questions
There is but one essential ingredient missing to this story, that of noetherianity of the remixed twisted cohomology ring $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$. At present, we conjecture that $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian, and in fact, generated by homomorphisms $k\to C[s]$ for $C$ an irreducible endotrivial. Sadly, the strategy of [BG25, Lemma 12.12] to prove noetherianity seems to not carry over to the general setting. We (sketchily) remark Balmer-Gallauer’s proof can be adapted to any finite abelian $p$-group, but we omit a proof, as it follows near identically as in the elementary abelian case.
A new possibility we propose is comparing twisted cohomology to the usual cohomology ring. We construct a homomorphism from the twisted cohomology ring to a direct product of $p$-local group cohomology rings (see Definition 5.5), and whose kernel is nilpotent. Any group cohomology ring is well-known to be noetherian [Ven59, Eve61], therefore if the image of $\prod\hat{\Psi}$ is finite in each $p$-local group cohomology ring, $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian modulo nilpotents. See Remark 6.6 for details.
Beyond this missing link, there remain numerous questions left to answer. First, it is quite reasonable to ask if a generalization of twisted cohomology exists for all finite groups. Not all endotrivials $C$ arise from a representation spheres (or any reasonable topological space), or can be expressed as (up to homotopy) a chain complex of permutation modules (see e.g. [BG23a, Corollary 5.6]). As a consequence, many of our maps $\iota^{H}_{C}$ cannot be constructed outside of the $p$-group setting; circumventing this issue is the main challenge. To go about constructing twisted cohomology, one may first have to determine $\operatorname{Pic}(\operatorname{K}_{b}(\operatorname{perm}(kG)))$, which at present is unknown (for non-$p$-groups). We ask if $\operatorname{Pic}(\operatorname{K}_{b}(\operatorname{perm}(kG)))$ and the subgroup of $\operatorname{Pic}(\mathscr{K}(G))$ generated by representation spheres are in fact equal, and if they are finite index subgroups of $\operatorname{Pic}(\mathscr{K}(G))$.
Second, while our twisted cohomology ring may be sufficient to capture the subtleties of $\operatorname{Spc}(\mathscr{K}(G))$, it is obviously more cumbersome. An explicit presentation in terms of generators and relations is at present seemingly beyond our reach. Even explicitly writing down endotrivials for larger groups (e.g. $p$-rank at least 3) is a significant challenge and relatively unexplored. Finally, although every endotrivial is a line bundle, it is currently unknown and an interesting question whether $\operatorname{Pic}(\mathscr{K}(G)|_{U(H)})\cong\mathbb{Z}$; we expect this to be the case. In correspondence, Gallauer gave a proof sketch of this for $p=2$ and $G$ elementary abelian, but not much is known beyond this case.
Organization
The paper is organized as follows. Section 1 covers some quick preliminaries regarding $p$-permutation modules, modular fixed points/Brauer quotients, and endotrivial complexes. Section 2 further reviews Borel-Smith functions. Section 3 covers chain complexes of free modules over the orbit category and effective endotrivials, and proves some technical lemmas. Section 4 constructs the local quasi-isomorphisms $\iota^{H}_{C}$ and the open cover $\{U(H)\}$ of the Balmer spectrum. Section 5 constructs the twisted cohomology ring. Section 6 proves injectivity of the comparison map and corollaries, generalizing the results of Balmer-Gallauer in the noetherian setting, and discusses noetherianity.
Notation
Our notation mostly follows [BG25], as opposed to [Mil25a]. Any group $G$ is assumed to be finite. The Weyl group G//H{G}/\!\!/{H} of $H\leq G$ is the subquotient $N_{G}(H)/H$. Given two subgroups $H,K\leq G$, we write $H=_{G}K$ if $H,K$ are $G$-conjugate, and $H\leq_{G}K$ to denote $H$ is a subgroup of a $G$-conjugate of $K$. We denote the set of all $p$-subgroups of $G$ by $s_{p}(G)$, and write $s_{p}(G)/G$ to denote a choice of conjugacy class representatives. We write $\operatorname{CF}(G)$ to denote the additive group of superclass functions, and if $G$ is not a $p$-group, we write $\operatorname{CF}(G,p)$ to denote the group of superclass functions valued on $p$-subgroups. We denote the trivial $kG$-module, the tensor unit of $\mathscr{K}(G)$, as $k$. Given a module $M$, $M[i]$ denotes the chain complex with $M$ in homological degree $i\in\mathbb{Z}$ and zero in all other degrees. [1]-[1] also denotes the shift functor on $\mathscr{K}(G)$. We denote by $\operatorname{H}^{\scriptscriptstyle\bullet}(G)=\operatorname{H}^{\scriptscriptstyle\bullet}(G;k)$ the $\mathbb{N}$-graded cohomology ring of the tensor unit $k$ in $\operatorname{D}_{b}(kG):=\operatorname{D}_{b}(kG\operatorname{-mod})$ (with both odd and even-degree shifts for $p$ odd).
Acknowledgments
I am indebted to Paul Balmer and Martin Gallauer for numerous conversations and unwieldy email threads about their work and tensor-triangular geometry in abstracto. I also am grateful to Robert Boltje for his mentorship and suggestions, those of which heavily influenced Section 3 in particular, and the UC Santa Cruz Department of Mathematics for their generous graduate student travel support, which enabled conversations leading to the ideas in this paper to occur.
1. Preliminaries
We expediently review some of the essentials of permutation modules from [BG25] and endotrivial complexes [Mil25a]. For an overview of tensor-triangular geometry, we refer the reader to [Bal05, Bal10b], and for an in-depth “classical” overview of $p$-permutation modules and the Brauer quotient, we refer the reader to [Lin18, Chapter 5] or [Las23].
Definition 1.1.
A $kG$-module $M$ is:
  1. (a)
    a permutation module if $M$ admits a $G$-stable basis, or equivalently, $M\cong kX$ for some $G$-set $X$;
  2. (b)
    a $p$-permutation module (where $p$ denotes the characteristic of $k$) if $M$ is a direct summand of a permutation module, or equivalently, $\operatorname{Res}^{G}_{S}M$ is a permutation module for some Sylow $p$-subgroup $S$ of $G$. In particular, if $G$ is a finite $p$-group, every $p$-permutation module is a permutation module.
We set $\mathscr{K}(G):=\operatorname{K}_{b}(p\operatorname{-perm})$, the bounded homotopy category of $p$-permutation modules.
Definition 1.2.
Let $K\leq H\leq G$ be subgroups of $G$. The augmentation homomorphism for $G$-sets $\operatorname{aug}\colon G/K\to G/H$ is the homomorphism induced by the assignment $1K\mapsto 1H$. The augmentation homomorphism between the permutation modules $\operatorname{aug}\colon k[G/K]\to k[G/H]$ is the induced map on vector spaces. We note the standard example of the augmentation homomorphism is when $H=G$ and $K=1$.
We have an analogous coaugmentation homomorphism $\operatorname{coaug}\colon k[G/H]\to k[G/K]$ induced from the assignment $1H\mapsto\sum_{g\in[H/K]gK}$. This homomorphism of permutation modules is in no way induced from any $G$-set homomorphism.
Remark 1.3.
By [BG23a], we have a surjective tt-functor $\mathscr{K}(G)\twoheadrightarrow\operatorname{D}_{b}(kG)$ with kernel $\mathscr{K}_{ac}(G)$, the tt-ideal consisting of all acyclic complexes. Set $V_{G}:=\operatorname{Spc}(\operatorname{D}_{b}(kG))$, then [BG25, Proposition 3.22] asserts that this functor induces an open inclusion $V_{G}\hookrightarrow\operatorname{Spc}(\mathscr{K}(G))$. We say $V_{G}\subseteq\operatorname{Spc}(\mathscr{K}(G))$ is the cohomological open. As noted in the introduction of [BG25], the crux of the matter is its closed complement, $\operatorname{supp}(\mathscr{K}_{ac})$.
The Brauer quotient is an explicit construction on the level of modules which mimics taking fixed points of $G$-sets. Balmer-Gallauer adapt this construction to work on the level of big categories, dubbing them modular fixed points functors.
Theorem 1.4.
[BG25, Proposition 2.7] For every $p$-subgroup $H\leq G$ there exists a coproduct-preserving tt-functor on the big derived category of permutation modules
$\Psi^{H}\colon\mathscr{D}(G)\to\mathscr{D}({G}/\!\!/{H})$
such that $\Psi^{H}(kX)\cong k[X^{H}]$ for every $G$-set $X$. In particular, this functor preserves compacts and restricts to a tt-functor $\Psi^{H}\colon\mathscr{K}(G)\to\mathscr{K}({G}/\!\!/{H})$.
Notation 1.5.
Each tt-functor $\Psi^{H}$ induces a continuous map on spectra
$\operatorname{Spc}(\Psi^{H}):=\psi^{H}\colon\operatorname{Spc}(\mathscr{K}({G}/\!\!/{H}))\to\operatorname{Spc}(\mathscr{K}(G)),$
and composing with the surjection $\hat{\Psi}^{H}\colon\mathscr{K}(G)\twoheadrightarrow\operatorname{D}_{b}(k[{G}/\!\!/{H}])$ induces another continuous map on spectra
$\hat{\psi}^{H}\colon V_{{G}/\!\!/{H}}\to\operatorname{Spc}(\mathscr{K}(G)).$
In fact, the Balmer spectrum, as a set, is built entirely from these “$p$-local” cohomological opens.
Theorem 1.6.
[BG25, Theorem 2.10] Every point of $\operatorname{Spc}(\mathscr{K}(G))$ is the image $\hat{\psi}^{H}(\mathfrak{p})$ of a point $\mathfrak{p}\in V_{{G}/\!\!/{H}}$ for some subgroup $H\leq G$ unique up to $G$-conjugation, i.e. we have $\hat{\psi}^{H}(\mathfrak{p})=\hat{\psi}^{H^{\prime}}(\mathfrak{p}^{\prime})$ if and only if there exists a $g\in G$ such that $^{g}H=H^{\prime}$ and $^{g}\mathfrak{p}=\mathfrak{p}^{\prime}$.
Notation 1.7.
We denote by $\mathscr{P}_{G}(H,\mathfrak{p})\in\operatorname{Spc}(\mathscr{K})$ the image of the prime $\mathfrak{p}\in V_{{G}/\!\!/{H}}$ under $\hat{\psi}^{H}:V_{{G}/\!\!/{H}}\to\operatorname{Spc}(\mathscr{K}(G))$. By [BG25, Theorem 2.10], every prime is, up to $G$-conjugation, uniquely expressible in this way.
Theorem 1.8.
[BG25, Theorem 2.11] The family of functors
$\{\mathscr{D}(G)\xrightarrow{\Psi^{H}}\mathscr{D}({G}/\!\!/{H})\twoheadrightarrow\operatorname{K}(\operatorname{Inj}(k[{G}/\!\!/{H}]))\}_{H\in s_{p}(G)/G}$
indexed by conjugacy classes of $p$-subgroups $H\leq G$ is conservative. This restricts to a conservative family of functors $\{\mathscr{K}(G)\xrightarrow{\Psi^{H}}\mathscr{K}({G}/\!\!/{H})\twoheadrightarrow\operatorname{D}_{b}(k[{G}/\!\!/{H}])\}_{H\in s_{p}(G)/G}$ on compacts.
Next, we briefly review the classification of endotrivial complexes.
Definition 1.9.
A chain complex $C\in\mathscr{K}(G)$ is endotrivial if it is an invertible object, i.e. $C^{*}\otimes_{k}C\simeq k[0]$.
Remark 1.10.
By the Künneth formula for $kG$-modules, for any $p$-subgroup $H\leq G$, $\Psi^{H}(C)$ has nonzero homology in exactly one homological degree, with that homology having $k$-dimension one. Let $h_{C}(H)\in\mathbb{Z}$ denote the unique integer for which $\operatorname{H}_{h_{C}(H)}(\Psi^{H}(C))\neq 0$.
Proposition 1.11.
[Mil24, Definition 3.5]
  1. (a)
    The function $h_{C}\colon s_{p}(G)\to\mathbb{Z}$ is a well-defined superclass function, i.e. a function constant on conjugacy classes of subgroups of $G$.
  2. (b)
    The assignment $h\colon\operatorname{Pic}(\mathscr{K}(G))\to\operatorname{CF}(G,p)$ given by $C\mapsto h_{C}$ is a well-defined group homomorphism.
Given an endotrivial $C$, we call $h_{C}$ the h-mark function of $C$, and call the homomorphism $h$ the h-mark homomorphism. Under the hood, $h$ is nothing more than a numerical avatar of the conservative family of functors $\{\hat{\Psi}^{H}\}_{H\in s_{p}(G)}$. Since the invertible objects of $\operatorname{D}_{b}(kG)$ consists of shifts of $k$-dimension one $kG$-modules, $h_{C}$ simply tracks the image of $C$ in $\operatorname{D}_{b}(k[{G}/\!\!/{H}])$ where $H$ runs through all $p$-subgroups of $G$.
Definition 1.12.
A Borel-Smith function is a superclass function $f\in\operatorname{CF}(G,p)$ satisfying the following three conditions, which we call the Borel-Smith conditions.
  1. (a)
    If $p$ is odd, then for any subquotient $T/S$ of $G$ of order $p$, $f(T)\equiv f(S)\mod 2$.
  2. (b)
    If $p=2$, then for any sequence of subgroups $H\trianglelefteq K\trianglelefteq L\leq N_{G}(H)$, with $[K:H]=2$, $f(K)\equiv f(H)\mod 2$ if $L/K$ is cyclic of order 4 and $f(K)\equiv f(H)\mod 4$ is $L/K$ is quaternion of order 8.
  3. (c)
    For any elementary abelian subquotient $T/S$ of $G$ of rank 2, the equality
    $f(S)-f(T)=\sum_{S<X<T}\big{(}f(X)-f(T)\big{)}$
    holds.
The collection of Borel-Smith functions forms an additive subgroup $\operatorname{CF}_{b}(G,p)$ of $\operatorname{CF}(G,p)$. In fact, if $G$ is a finite $p$-group, then under the identification $\operatorname{CF}(G)\cong B^{*}(G):=\operatorname{Hom}(B(G),\mathbb{Z})$, $\operatorname{CF}_{b}(G)$ forms a rational $p$-biset subfunctor of $\operatorname{CF}(G)$. See [BY07, Proposition 3.7] for details.
Theorem 1.13.
We have a complete characterization of endotrivials.
  1. (a)
    [Mil25a, Theorem 4.6] Let $G$ be a finite $p$-group. The h-mark homomorphism is an isomorphism onto the subgroup of Borel-Smith functions
    $h\colon\operatorname{Pic}(\mathscr{K}(G))\cong\operatorname{CF}_{b}(G).$
  2. (b)
    [Mil25a, Corollary 6.4] Let $G$ be a finite group. The h-mark homomorphism has image $\operatorname{CF}_{b}(G,p)$ and induces a split exact sequence
    $0\to\operatorname{Hom}(G,k^{\times})\to\operatorname{Pic}(\mathscr{K}(G))\to\operatorname{CF}_{b}(G,p)\to 0.$
Example 1.14.
Let $G$ be a finite $p$-group. The first nontrivial example of an endotrivial, and those which make the magic happen in the elementary abelian case, are as follows (c.f. [BG25, Definition 12.3]). If $p=2$, one has an endotrivial for $C_{2}$ given by $kC_{2}\to k,$ with $k$ in homological degree 0 and the nonzero differential given by the augmentation homomorphism. Otherwise if $p$ is odd, one has an endotrivial for $C_{p}$ by truncating a periodic resolution of the trivial $kC_{p}$-module $k$, $kC_{p}\to kC_{p}\to k$ with $k$ in homological degree 0. For any subgroup $N\triangleleft G$ of index $p$, inflation yields the following complex
$u_{N}:=k[G/N]\to k[G/N]\to k$
Here $p$ is assumed odd, if $p=2$ then the highest degree term can be deleted.
A less easy example of an endotrivial for $p=2$ and $G=D_{2^{n}}$ is as follows. Let $H_{1},H_{2}$ be nonconjugate, noncentral subgroups of order 2 (this choice is unique up to conjugacy and reordering). Then the following complex is endotrivial, and a small example of a nontrivial faithful, irreducible endotrivial, i.e. one arising from a faithful irreducible real representation (see Definition 2.5),
$kD_{2^{n}}\to k[D_{2^{n}}/H_{1}]\oplus k[D_{2^{n}}/H_{1}]\to k,$
where all maps between indecomposable permutation modules are augmentation homomorphisms.
2. Borel-Smith functions and representation spheres
We begin by considering the topological properties of the main characters of our story, the endotrivials arising from genuine real representations, i.e. representations spheres. Such endotrivials come from objects of higher structure; they arise from chain complexes of free modules over the orbit category $\Gamma_{G}$, and in particular, their differentials are augmentation homomorphisms which are well-behaved with respect to local-global considerations. We discuss this now.
Definition 2.1.
We say that a superclass function $f$ is effective or if the following holds: if $K\leq H$ are subgroups of $G$, then $f(K)\geq f(H)$. That is, $f$ is monotonically decreasing with respect to the post of subgroups of $G$. Similarly, we say that an endotrivial $C$ is effective if its corresponding h-mark function $h_{C}$ is effective.
This choice of terminology will become clear in the sequel. We first recall some important facts about Borel-Smith functions. Let $G$ be a nilpotent group and $F$ a field. Given a $FG$-module $V$, the dimension function associated to $V$ is the superclass function
$\dim\colon H\mapsto\dim_{F}V^{H}.$
If $F$ has characteristic 0, this induces a group homomorphism $R_{F}(G)\to\operatorname{CF}(G)$. When $F=\mathbb{R}$, topologists may write $RO(G)=R_{\mathbb{R}}(G)$, as a representation sphere is nothing more than a one-point compactification of a real representation.
Theorem 2.2.
[tD87, Theorem 5.4 and Theorem 5.13, pages 211 and 216] Let $G$ be a nilpotent group. The image of $\dim\colon RO(G)\to\operatorname{CF}(G)$ is the group of Borel-Smith functions $\operatorname{CF}_{b}(G)$. Moreover, if $f$ is a effective Borel-Smith function, there exists a real representation $V$ for which $\dim(V)=f$.
Corollary 2.3.
Let $G$ be a nilpotent group and suppose $f$ is a Borel-Smith function. Then $f$ can be expressed as the difference of two effective Borel-Smith functions. In particular, every endotrivial is isomorphic in $\mathscr{K}(G)$ to the product of an effective endotrivial and the dual of an effective endotrivial.
Proof.
This follows from [tD87, Theorem 5.4, page 211] since the dimension function of any representation is effective, and [Mil25a, Theorem 4.6], since the image of the h-mark homomorphism is precisely the group of Borel-Smith functions of $G$. ∎
The previous corollary tells us that it suffices to work with effective endotrivials, which are particularly well-behaved. In fact, when $G$ is a nilpotent group, the dimension function gives a canonical effective basis of $\operatorname{CF}_{b}(G)$, which arises from the real irreps of $G$. Therefore, if $G$ is a finite $p$-group, we have a corresponding canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K})$ consisting of effective endotrivials.
Theorem 2.4.
Let $V_{1},\dots,V_{n}$ denote the irreducible real representations of $G$. Then the set of associated dimension functions $\{f_{1},\dots,f_{n}\}$, after removing duplicates, forms a $\mathbb{Z}$-basis of $\operatorname{CF}_{b}(G)$. In particular, there is an associated canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K}(G))$.
Proof.
Since the image of the dimension homomorphism is precisely $\operatorname{CF}_{b}(G)$, it suffices to show that the set $\mathcal{B}:=\{f_{1},\dots,f_{n}\}$ is linearly independent after removing duplicates. [tD87, Proposition 5.9, page 213] asserts that $\ker(\dim)$ is generated by elements of the form $V-\psi^{k}(V)$, where $V$ is an irreducible real representation, $\psi^{k}$ is the $k$-th Adams operation, and $k$ is coprime to $|G|$. Therefore, the duplicates in $\mathcal{B}$ arise from Adams operation conjugates, and it follows that any set of real irreducible representations which are not Adams operation conjugates will correspond to a linearly independent set of Borel-Smith functions. ∎
Definition 2.5.
We call the Borel-Smith functions (resp. endotrivials) associated to the real irreps of $G$ the irreducible Borel-Smith functions (resp. endotrivials). These objects are necessarily effective.
Remark 2.6.
Given a real representation $V$ of $G$ with corresponding character $\chi$, we have an equality
$\dim_{\mathbb{R}}V^{H}=\frac{1}{|H|}\sum_{h\in H}\chi(h),$
a practical method of computing the basis of $\operatorname{CF}_{b}(G)$.
Checking character tables of $G$ a finite $p$-group of normal $p$-rank one shows the basis of $\operatorname{Pic}(\mathscr{K}(G))$ obtained in [Mil24, Section 6] coincides with the canonical $\mathbb{Z}$-basis of $\operatorname{CF}_{b}(G)$. (The author remarks that he finds this is rather surprising, as the computations of [Mil24] were performed entirely ad-hoc and were done prior to the classification of $\operatorname{Pic}(\mathscr{K}(G))$.)
Notation 2.7.
Let $G$ be a finite $p$-group. Given a $\mathbb{R}G$-module $V$, let $f_{V}$ and $C_{V}$ denote the corresponding effective Borel-Smith function and effective indecomposable endotrivial respectively. That is, $f_{V}:=\dim V$ and $C_{V}$ is the unique indecomposable endotrivial for which $h_{C_{V}}=\dim V$.
Per tom Dieck [tD87], Borel-Smith functions correspond to representation spheres, which is simply a manifestation of 2.2. If $X$ is a finite-dimensional $G$-complex which is a $\mathbb{F}_{p}$-cohomology sphere, the fixed-point set $X^{H}$ is a $\mathbb{F}_{p}$-cohomology-sphere of dimension $d(H)$. The assignment $H\mapsto d(H)+1$ is (unfortunately) also called the dimension function of $X$. The set of dimension functions of representation spheres is precisely the effective non-negative Borel-Smith functions. Given a real representation $V$, let $S^{V}$ denote the corresponding representation sphere.
3. Chain complexes over the orbit category
Since effective endotrivials identify with representation spheres, they have additional structural properties, as representation spheres (and more generally, any $G$-CW-complex) produce free chain complexes over the orbit category $\Gamma_{G}$. We first review this notion, following [Yal17], then prove some technical results which will be critical for the sequel.
Definition 3.1.
Let $\Gamma_{G}$ denote the orbit category of $G$. The objects of $\Gamma_{G}$ are transitive $G$-sets $G/H$ for subgroups $H\leq G$, and the morphisms from $G/H$ to $G/K$ are $G$-set homomorphisms $G/H\to G/K$. In particular, $\operatorname{Hom}_{\Gamma_{G}}(H,K)$ is empty unless a conjugate of $H$ is a subgroup of $K$.
A $k\Gamma_{G}$-module $M$ is a contravariant functor from the category $\Gamma_{G}$ to the category of $k$-modules. By identifying $\operatorname{Aut}_{\Gamma_{G}}(G/H)$ with G//H{G}/\!\!/{H}, $M(H)$ has k[G//H]k[{G}/\!\!/{H}]-module structure. The category of finitely generated $k\Gamma_{G}$-modules, denoted $\operatorname{mod}(\Gamma_{G})$ is abelian, so the usual categorical concepts apply.
Given a $G$-set $X$, we define the $k\Gamma_{G}$-module $k[X^{?}]$ as the module with value at $G/H$ given by $k[X^{H}]$, with the obvious induced maps. A module over $\Gamma_{G}$ is free if it is isomorphic to a direct sum of modules of the form $k[(G/K)^{?}]$. Let $\operatorname{proj}(\Gamma_{G})$ denote the full subcategory of $\operatorname{mod}(\Gamma_{G})$ consisting of free $\Gamma_{G}$-modules (by the Yoneda lemma, every projective $\Gamma_{G}$-module is free, see [Yal17, Definition 3.1]).
The next propositions follow immediately from definition of $k\Gamma_{G}$-modules. We refer to these facts as the stabilizers grow conditions. This property and name were first suggested to the author by Robert Boltje.
Proposition 3.2.
Let $X$ and $Y$ be two $G$-sets. Any $k\Gamma_{G}$-homomorphism $f\colon k[X^{?}]\to k[Y^{?}]$ satisfies the following property: given any $x\in X$ and subgroup $H\leq G$, if $x\in X^{H}$, then $f(x)\in k[Y^{H}]$. In particular, $\operatorname{Hom}_{\operatorname{mod}(\Gamma_{G})}(k[X^{?}],k[Y^{?}])$ has $k$-basis induced from the set $\operatorname{Hom}_{G-\operatorname{set}}(X,Y)$.
Proposition 3.3.
Let $k[X^{?}]$ and $k[Y^{?}]$ be two free $k\Gamma_{G}$-modules and let $f\colon k[X^{?}]\to k[Y^{?}]$ be a $k\Gamma_{G}$-module homomorphism. Let $K\leq_{G}H$ be two subgroups of $G$. For any $m\in k[X^{H}]$, if $f^{H}(m)=n\in k[Y^{H}]$, then regarding $m$ as an element of $k[X^{K}]\supseteq k[X^{H}]$ of k[G//H]k[{G}/\!\!/{H}]-modules, $f^{K}(m)=n\in k[Y^{K}]$.
Remark 3.4.
As a result, each free $k\Gamma_{G}$-module has a stratification by fixed points. That is, given a free $k\Gamma_{G}$-module $kX$, we may write $kX=kG^{\oplus a_{1}}\oplus k[G/H_{2}]^{\oplus a_{2}}\cdots\oplus k[G/G]^{a_{n}}$ for some $a_{1},\dots,a_{n}\in\mathbb{N}$, where each direct summand $k[G/H_{i}]^{\oplus a_{i}}$ is a uniquely determined submodule of $kX$. Therefore, we may fix a “canonical permutation basis” in accordance with this stratification. We will make considerable use of this fact in the sequel.
Given a $G$-set $X$ and two subgroups $K\leq H$ of $G$, we have a canonical inclusion homomorphism (of $k[N_{G}(H)\cap N_{G}(K)]$-modules) $i\colon k[X^{H}]\to k[X^{K}]$ associated to the $G$-set homomorphism $G/K\to G/H,1K\mapsto 1H$. This comes associated with a (not necessarily unique) projection map $p\colon k[X^{K}]\to k[X^{H}]$ satisfying $p\circ i=\operatorname{id}$ and $p(m)=0$ if $m\in\Psi^{H}(k[X^{K}])=0$. After choosing a canonical permutation basis, we obtain a corresponding projection map associated to the basis.
For any subgroup $H\leq G$, we have an obvious functor $(H)\colon\operatorname{proj}(\Gamma_{G})\to\operatorname{perm}(k[{G}/\!\!/{H}])$ from the assignment $k[X^{?}]\to k[X^{H}]$. Modular fixed points behave as one would hope.
Proposition 3.5.
Let $P$ be a $p$-subgroup of $G$. Then the following diagram commutes up to natural isomorphism.
$\operatorname{proj}(\Gamma_{G})$$\operatorname{perm}(kG)$$\operatorname{perm}(k[{G}/\!\!/{P}])$$\scriptstyle{-(1)}$$\scriptstyle{-(P)}$$\scriptstyle{\Psi^{P}}$
Proof.
This follows from the natural isomorphism $k[X^{P}]\cong\Psi^{P}(kX)$. ∎
Definition 3.6.
Let $X$ be a $G$-CW-complex. The reduced chain complex $X$ over the orbit category is the functor $\tilde{C}_{\ast}(X^{?};k)$ from the orbit category $\Gamma_{G}$ to the category of chain complexes of $k$-modules. This gives rise to a chain complex of free (by [Yal17, Lemma 3.2]) $k\Gamma_{G}$-modules
$\tilde{C}_{\ast}(X^{?};k):=\cdots\to C_{i}(X^{?};k)\xrightarrow{d_{i}}C_{i-1}(X^{?};k)\to\cdots\to C_{0}(X^{?};k)\xrightarrow{\epsilon}\underline{k}\to 0,$
where $\underline{k}$ denotes the constant functor with values $k(H)=k$ for all subgroups $H\leq G$, and $\epsilon$ denotes the augmentation homomorphism (this is essentially the construction of reduced Bredon homology, see [Ill73]). We denote chain complexes of free $k\Gamma_{G}$-modules by $C^{?}$, and their evaluations at the subgroup $H\leq G$ by $C^{H}$ to inspire an aura of fixed points.
If $G$ is a finite $p$-group and $S$ is a representation sphere of $G$, it may be realized as a $G$-CW-complex, hence producing a chain complex of free $k\Gamma_{G}$-modules. Therefore for any subgroup $H\leq G$, $\tilde{C}_{\ast}(S^{H};k)$ is an effective endotrivial complex of permutation k[G//H]k[{G}/\!\!/{H}]-modules. Conversely, after shifting, every effective endotrivial complex $C$ of $kG$-modules is homotopy equivalent to $\tilde{C}_{\ast}(S^{1};k)$ for some representation sphere $S$ of $G$ with dimension homomorphism $\dim(S)=h_{C}$ (see [tD87, Page 217, Theorem 5.16]), and hence up to homotopy arises from a chain complex of free $k\Gamma_{G}$-modules.
This proposition is critical for many of the constructions to follow: it allows us to perform lifts of $p$-local homomorphisms $k\to\Psi^{H}(C)$.
Proposition 3.7.
Let $G$ be a finite $p$-group, let $k[X^{?}]$ and $k[Y^{?}]$ be two free $k\Gamma_{G}$-modules with $X$ transitive, let $f\colon k[X^{?}]\to k[Y^{?}]$ be a $k\Gamma_{G}$-module homomorphism, and let $H\leq G$ be a subgroup. Suppose $k[X^{H}]\neq 0$, and let $M^{\prime}$ denote the unique minimal nonzero submodule of $k[X^{H}]$. If $M^{\prime}\subseteq\ker(f^{H})$, then there exists a submodule $M^{?}\subseteq k[X^{?}]$, minimal with respect to the property that $M^{L}\neq 0$ when $k[X^{L}]\neq 0$, such that $M^{H}=M^{\prime}$ and for all subgroups $K\leq_{G}H$, $M^{K}\subseteq\ker(f^{K})$.
Proof.
We define $M^{?}\subseteq k[X^{?}]$ as follows: for $L\leq G$, $M^{L}\subset k[X^{L}]$ is the unique minimal submodule if $X$ has any $K$-fixed points, and the zero module if not. It is straightforward that, by construction, $M^{?}$ is minimal with respect to the property that $M^{L}\neq 0$ when $k[X^{L}]\neq 0$. Now by minimality of $M^{?}$, it suffices to show $f^{K}$ is not an injective k[G//K]k[{G}/\!\!/{K}]-module homomorphism for all $K\leq_{G}H$, but this follows directly from Proposition 3.3. ∎
Remark 3.8.
Proposition 3.7 asserts that we can construct “local” inclusions $k\to C[s]$ for chain complexes arising from chain complexes over the orbit category. This does not hold for arbitrary objects of $\mathscr{K}(G)$. For instance, the two-term complex $C=k\xrightarrow{\operatorname{coaug}}kG$ (with $k$ in homological degree 0) has no nonzero global homomorphism $k\to C$, but there exist nonzero local homomorphisms $k\to\Psi^{H}(C)$ (in fact isomorphisms) for every nontrivial subgroup $H\leq G$, since $\Psi^{H}(kG)=0$.
3.1. Removing contractible summands
We prove a couple technical results for finite $p$-groups that show we can remove contractible summands from chain complexes of free $k\Gamma_{G}$-modules if and only if we can for the corresponding chain complexes of permutation $kG$-modules (in particular, for effective endotrivials). Though most of these results are not strictly necessary for future sections, they provide elucidation and may be of independent interest.
Remark 3.9.
A chain complex of free $k\Gamma_{G}$ modules has in each homological degree a canonical permutation basis, and each differential respects the stabilizers grow condition. Moreover, it is easy to see that given a chain complex $C$ of permutation $kG$-modules, if one can choose a canonical permutation basis of $G$-sets in each degree such that the stabilizers grow condition holds with respect to this basis, then $C$ may be realized as the image of a chain complex of free $k\Gamma_{G}$-modules evaluated at the subgroup $1\leq G$. In this sense, the chain complexes of permutation $kG$-modules satisfying the stabilizers grow condition on differentials are exactly those which arise from chain complexes of free $k\Gamma_{G}$-modules.
If $C$ is a chain complex of permutation modules, we say $C$ arises from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules if there exists a chain complex of free $k\Gamma_{G}$-modules $C^{?}$ such that $C^{1}\cong C$ as chain complexes.
Proposition 3.10.
Let $G$ be a finite group and suppose $C$ is a bounded chain complex of permutation $kG$-modules arising from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules. Let $i$ be an integer and suppose $T\subseteq C_{i}$ is a $k$-dimension one submodule. If $T\not\in\ker(d_{i})$, then there exists a contractible direct summand $K$ of $C$ consisting of permutation modules such that $T\subseteq K_{i}$.
Proof.
We have a canonical permutation basis associated to $C_{i}$ arising from the $k\Gamma_{G}$-module structure, $C_{i}\cong kX_{1}\oplus\cdots\oplus kX_{m}$ with each $X_{j}$ a transitive $G$-set. For $j\in\{1,\dots,m\}$, let $T_{j}$ denote the projection of $T$ into $kX_{j}$. Similarly, we have a canonical permutation basis $C_{i-1}\cong kY_{1}\oplus\cdots\oplus kY_{n}$ with each $Y_{j}$ a transitive $G$-set. For $j\in\{1,\dots,n\}$, let $U_{j}$ denote the projection of $T$ into $kY_{j}$. Finally, let $\{b_{1},\dots,b_{l}\}$ denote the indices for which the projection $p_{j}\circ d_{i}(T)$ onto $U_{j}$ is nonzero. By assumption this set is nonempty. Set $b$ equal to the index $b_{j}$ for which $Y_{b_{j}}$ has minimal stabilizers (this is well-defined since $Y_{b_{j}}$ is transitive).
It follows from Corollary 3.3 that there exists an index $a\in\{1,\dots,m\}$ such that the composition $p_{b}\circ d_{i}\circ i_{a}$ is an isomorphism, where $p_{b}$ denotes projection onto $kY_{b}$ and $i_{a}$ denotes inclusion into $kX_{a}$. In particular, $X_{a}$ and $Y_{b}$ are isomorphic $G$-sets. Note the choice of $a$ is not necessarily unique (for example, the two-term complex $k\oplus k\to k$ with differential $(\operatorname{id},\operatorname{id})$).
Set $C_{i}^{\prime}:=kX_{1}\oplus\cdots\oplus\widehat{kX}_{a}\oplus\cdots\oplus kX_{m}$ and $C_{i-1}^{\prime}:=kY_{1}\oplus\cdots\oplus\widehat{kY}_{b}\oplus\cdots\oplus kY_{n}$. We will now modify the canonical bases of $C_{i}$ and $C_{i-1}$ to construct the contractible chain complex $K$ that splits off of $C$. Choose any generator $t\in T$ of $T$, then we may write $t=p_{a}(t)+t^{\prime}$, with $t^{\prime}\in T_{1}\oplus\cdots\oplus\widehat{T}_{a}\oplus\cdots\oplus T_{m}$. The $kG$-modules $\langle p_{a}(t)\rangle$ and $\langle t^{\prime}\rangle$ are both isomorphic to $k$, i.e. are $G$-stable. We replace the $C_{i}$ basis elements $\{x_{1},\dots,x_{e}\}=X_{a}\subset kX_{a}\subseteq C_{i}$ with $\{x_{1}+t^{\prime},\dots,x_{e}+t^{\prime}\}=X_{a}^{\prime}\subset C_{i}$. $X_{a}^{\prime}$ is again a $G$-set, and we have $C_{i}=C_{i}^{\prime}\oplus kX^{\prime}_{a}$. Moreover, $T\subseteq kX_{a}^{\prime}$ by construction.
We replace the $C_{i-1}$ basis elements $\{y_{1},\dots,y_{f}\}=Y_{b}\subset kY_{b}\subseteq C_{i-1}$ with $\{d_{i}(x_{1}+t^{\prime}),\dots,d_{i}(x_{e}+t^{\prime})\}=X_{b}^{\prime}\subset C_{i-1}$. Again, $Y_{b}^{\prime}$ is a $G$-set. We claim that $p_{b}(X_{a}^{\prime})$ is a $k$-basis of $Y_{b}$. This follows because $\langle p_{b}\circ d_{i}(t^{\prime})\rangle$ is either 0 or the unique $k$-dimension one submodule of $kY_{b}$, and we chose $b$ such that both $p_{b}\circ d_{i}\circ i_{a}$ was an isomorphism and $p_{b}\circ d_{i}(T)$ was nonzero. Therefore $p_{b}\circ d_{i}(y_{j}+t^{\prime})$ has the same $G$-stabilizer as both $y_{j}$ and $y_{j}+t^{\prime}$. Hence $Y_{b}^{\prime}$ is isomorphic as a $G$-set to $Y_{b}$ and $X_{a}$, and it follows that $C_{i-1}=C^{\prime}_{i-1}\oplus kY^{\prime}_{b}$. Moreover, $d_{i}$ restricts to an isomorphism $kX_{a}^{\prime}\xrightarrow{\cong}kY_{b}^{\prime}$, and the result follows. ∎
The next proposition demonstrates that if a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules satisfies that $C^{1}$ has a contractible summand (as a chain complex of $kG$-modules), then $C^{?}$ itself has a contractible summand (as a chain complex of $k\Gamma_{G}$-modules).
Proposition 3.11.
Let $G$ be a finite group and suppose $C$ is a bounded chain complex of permutation $kG$-modules arising from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules. If $C$ contains a contractible direct summand $K$, then there exists a direct sum decomposition $C=K\oplus D$ such that $D$ also arises from a chain complex $D^{?}$ of free $k\Gamma_{G}$-modules.
In particular, $C$ has a contractible direct summand of permutation modules if and only if $C^{?}$ has a contractible direct summand.
Proof.
To prove this, it suffices to find a decomposition of $C$ into direct summands $K,D$ as stated and show that each homological degree of $D$ has a canonical permutation basis satisfying the stabilizers grow condition. First, since $C$ has a contractible direct summand, there exists an integer $i\in\mathbb{Z}$ and $k$-dimension one submodule $T\subseteq C_{i}$ such that $T\not\in\ker(d_{i})$. Let $C_{i}=kX_{1}\oplus\cdots\oplus kX_{m}$ and $C_{i-1}=kY_{1}\oplus\cdots\oplus kY_{n}$ be the direct sum decompositions into canonical bases, with each $X_{j},Y_{j}$ a transitive permutation module. After projecting $T$ onto each $kX_{j}$, it follows that there exists an index $l$ for which $d_{i}|_{X_{l}}$ is injective. Choose the index $l$ such that the stabilizer of $X_{l}$ is minimal with respect to subgroup inclusion.
We modify the canonical bases of $C_{i}$ and $C_{i-1}$ as follows. First, it follows from Corollary 3.3 that there exists a (non-unique) index $h\in\{1,\dots,n\}$ such that $p_{h}\circ d_{i}\circ i_{l}$ is an isomorphism, where $i_{l}$ denotes inclusion into $kX_{l}$ and $p_{h}$ denotes projection onto $kY_{h}$. We replace the basis elements of $Y_{h}$ with the basis elements $Y_{h}^{\prime}=\{d_{i}(x)\mid x\in X_{l}\}$. Then $Y_{h}^{\prime}$ is a $G$-set isomorphic to $X_{l}$, and since the projection of $kY_{h}^{\prime}$ onto $kY_{h}$ is an isomorphism, this forms a new basis of $C_{i-1}$. Finally, we have $d_{i-1}(kY_{h}^{\prime})=0$ since $d_{i-1}\circ d_{i}=0$, so the resulting basis still respects the stabilizers grow property. Set $D_{i-1}:=kY_{1}\oplus\cdots\oplus\widehat{kY}_{h}\oplus\cdots\oplus kY_{n}$ and $K_{i-1}:=kY_{h}^{\prime}$.
Set $K_{i}:=kX_{l}$, by construction $d_{i}|_{K_{i}}$ is an isomorphism and maps canonical permutation basis to canonical permutation basis. We next modify the canonical permutation basis of $C_{n}$ to obtain a direct summand $D_{i}$ for which $d_{i}(D_{i})\subseteq D_{i-1}$. Set $C_{i}^{\prime}:=kX_{1}\oplus\cdots\oplus\widehat{kX}_{l}\oplus\cdots\oplus kX_{m}$, so we have $C_{i}=C_{i}^{\prime}\oplus K_{i}$. For each canonical permutation basis element $x$ of $C_{i}$ (i.e. some $x\in X_{j}$ for $j\neq l$) replace $x$ with $x^{\prime}:=x-((d_{i}|_{kX_{l}})^{-1}\circ p_{h}^{\prime}\circ d_{i})(x)$, where $p_{h}^{\prime}$ denotes projection onto $kY_{h}^{\prime}$ and $(d_{i}|_{kX_{l}})^{-1}$ denotes the inverse of the isomorphic projection of $kX_{l}$ onto $kY_{h}^{\prime}$. A straightforward computation shows $d_{i}(x^{\prime})\in D_{i-1}$. By construction, $((d_{i}|_{kX_{l}})^{-1}\circ p_{h}^{\prime}\circ d_{i})(x)\in K_{i}$, and the collection $X^{\prime}$ of the $x^{\prime}$ forms a $G$-set, with $X^{\prime}\cong X_{1}\sqcup\cdots\sqcup\widehat{X}_{l}\sqcup\cdots\sqcup X_{m}$. Set $D_{i}:=kX^{\prime}$, then we have a direct sum decomposition $C_{i}=D_{i}\oplus K_{i}$. Finally, for all $j\neq i,i-1$, set $D_{j}=C_{j}$. It follows that we have a direct sum decomposition $C=K\oplus D$. Finally, observe that for each $x^{\prime}$, under the image of $d_{i}$, the canonical permutation basis representation of $x^{\prime}$ in $D_{i-1}$ is identical to the canonical permutation basis representation of $x$ in $D_{i-1}$, therefore the stabilizers grow property holds for the differential $d_{i}$.
It remains to show the stabilizers grow property holds for the differentials $d_{i-1}$ and $d_{i+1}$ with respect to the same canonical bases for $C_{i+1}$ and $C_{i-1}$. The condition holding for $d_{i-1}$ is straightforward, since the only modified basis elements in $C_{i-1}$ now belong to $\ker(d_{i-1})$. Similarly, $\operatorname{im}(d_{i+1})\subseteq D_{i}$, and under the decomposition $x^{\prime}:=x-((d_{i}|_{kX_{l}})^{-1}\circ p_{h}^{\prime}\circ d_{i})(x)$ into $C_{i}^{\prime}$ and $kY^{\prime}_{h}$ respectively, it follows by construction that if a canonical permutation basis element $z\in C_{i+1}$ satisfies $d_{i+1}(z)=\sum a_{j}x_{j}+m$ with $m\in kX_{l}$ and each $x_{j}$ a canonical permutation basis element of $C_{i}^{\prime}$, then $d_{i+1}(z)=\sum a_{j}x_{j}^{\prime}$ where $x_{j}^{\prime}$ is the refined basis element of $X^{\prime}$ corresponding to $x_{j}$. Thus $d_{i+1}$ also satisfies the stabilizers grow condition, as desired.
In particular, if $G$ is a finite $p$-group, then every indecomposable effective endotrivial arises from a free chain complex of $k\Gamma_{G}$-modules.
Corollary 3.12.
Let $G$ be a finite $p$-group and let $C$ be an indecomposable effective endotrivial of $kG$-modules. Then $C$ arises from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules.
Proof.
Recall every effective endotrivial (after a possible shift) corresponds to a real representation $V$ of $\mathbb{R}G$-modules. The representation sphere $S^{V}$ is a $G$-CW-complex, therefore produces a (reduced) chain complex $C^{?}$ for which $C^{1}$ is an endotrivial with the same h-marks as $C$. By Corollary 3.11, we can remove all contractible summands from $C^{?}$ until $C^{1}$ is indecomposable, and since two indecomposable endotrivials with the same h-marks are isomorphic, the result follows. ∎
Moreover, we can completely identify the homomorphisms $k\to C$ for an effective endotrivial $C$.
Corollary 3.13.
Let $G$ be a finite $p$-group and let $C^{?}$ be a chain complex of free $k\Gamma_{G}$-modules with no contractible summands (e.g. an indecomposable effective endotrivial). For every integer $s\in\mathbb{Z}$ and submodule $T\subseteq C^{1}_{i}$ isomorphic to $k$, there exist a chain complex homomorphism $f\colon k\to C[s]$ with $\operatorname{im}(f_{0})=T$. In particular, $\dim_{k}\operatorname{Hom}_{\mathscr{K}(G)}(k[0],C^{1}[s])$ is equal to the number of indecomposable direct summands of $C_{s}$.
Proof.
The existence of such a homomorphism $f$ is equivalent to the inclusion $T\subseteq\ker(d_{i}),$ and this inclusion holds since if not, Proposition 3.10 implies the existence of a contractible summand of $C^{1}$, hence a contractible summand of $C^{?}$, which cannot occur. The last statement is straightforward. ∎
4. An open cover of the Balmer spectrum via endotrivials
For the rest of the paper, we assume $G$ is a finite $p$-group.
4.1. $p$-local quasi-isomorphisms
Proposition 4.1.
Let $G$ be a finite $p$-group and let $V$ be a $\mathbb{R}G$-module with kernel $N\trianglelefteq G$. Then the associated descending indecomposable endotrivial $C_{V}$ satisfies the following property: $n:=h_{C_{V}}(1)$ is the maximal nonzero homological degree of $C_{V}$, and $(C_{V})_{n}\cong k[G/N]$.
Proof.
Set $C:=C_{V}$. First, $V$ is equivalently a faithful $\mathbb{R}[G/N]$-module, so it suffices to assume $N=1$ after replacing $G$ with $G/N$, and it suffices to show $C_{n}$ is indecomposable and projective. Since $C$ is effective, it follows by an inductive argument that for all subgroups $H\leq G$, $\Psi^{H}(C)$ is isomorphic to an indecomposable complex in $\mathscr{K}({G}/\!\!/{H})$ whose highest nonzero homological degree is $h_{C}(H)$ (see [Lin18, Proposition 5.8.11]). In particular, $n=h_{C}(1)$ is the maximal nonzero degree of $C$, since $C$ is indecomposable.
First, we show that there cannot exist any non-projective direct summands of $C_{n}$, and that more generally, the highest homological degree $i$ for which the permutation module $k[G/H]$ can occur as a direct summand of $C_{i}$ is $i=h_{C}(H).$ Since $V$ is a faithful representation, $h_{C}=\dim V$ satisfies $h_{C}(H)<h_{C}(1)$ for any nontrivial subgroup $H\leq G$. Let $j$ be the highest degree for which $\Psi^{H}(C)_{j}\neq 0$. If $j\leq h_{C}(H)$, there is nothing to show, as this implies that only permutation modules with stabilizers not contained in $H$ occur in degrees above $h_{C}(H)$. Otherwise, if $j>h_{C}(H)$, since $\Psi^{H}(C)$ is isomorphic in $\mathscr{K}({G}/\!\!/{H})$ to an indecomposable complex with highest nonzero degree $h_{C}(H)$, the differential $\Psi^{H}(C)_{j}\xrightarrow{\Psi^{H}(d_{j})}\Psi^{H}(C)_{j-1}$ is split injective. But now, we may write $C_{j}\xrightarrow{d_{j}}C_{j-1}$ as follows:
$M_{1}$$N_{1}$$M_{2}$$N_{2}$$\scriptstyle{d_{1}^{1}}$
Here, we have $C_{j}=M_{1}\oplus M_{2}$, $C_{j-1}=N_{1}\oplus N_{2}$, where $M_{1}$ and $N_{1}$ satisfy $\Psi^{H}(M_{1})=\Psi^{H}(C_{j})$ and $\Psi^{H}(N_{1})=\Psi^{H}(C)_{j-1}$ and $M_{2}$ and $N_{2}$ satisfy $\Psi^{H}(M_{2})=0$ and $\Psi^{H}(N_{2})=0$. With this setup, we have $\Psi^{H}(d_{1}^{1})=\Psi^{H}(d_{j})$. By applying [Mil24, Lemma 5.7] and its dual statement, $d_{1}^{1}$ is an isomorphism if and only if $\Psi^{K}(d_{1}^{1})$ is an isomorphism for all $K\geq H$. Since $\Psi^{H}(d_{1}^{1})=\Psi^{H}(d_{j})$ is an isomorphism, it follows by an inductive argument up the poset of subgroups $K$ of $G$ containing $H$ that $\Psi^{H}(d_{1}^{1})$ is an isomorphism. Now by a standard homological algebra argument (c.f. [BM23, Lemma 9.2]), $M_{1}\xrightarrow{d_{1}^{1}}N_{1}$ splits off as a contractible direct summand of $C$, a contradiction since $C$ was assumed to be indecomposable. Therefore, $j=h_{C}(H)$. We conclude the highest homological degree $i$ for which the permutation module $k[G/H]$ can occur as a direct summand of $C_{i}$ is $i=h_{C}(H).$ In particular, since $V$ is faithful, only projective modules can occur in $C_{n}$.
It remains to show $C_{n}$ is indecomposable. Suppose for contradiction $C_{n}\cong P_{1}\oplus P_{2}$ for projective $kG$-modules $P_{1},P_{2}$. Since $\dim_{k}H_{n}(C)=1$, either $d_{n}|_{P_{1}}$ or $d_{n}|_{P_{2}}$ is injective, hence split injective, and it follows that a contractible chain complex containing $P_{1}$ or $P_{2}$ splits off from $C$, contradicting indecomposability of $C$. Thus, $C_{n}$ is indecomposable projective, as desired. ∎
In the situation described in Proposition 4.1, $\ker(d_{n})$ is the unique submodule of $k[G/N]$ of $k$-dimension 1. We are now ready to generalize the $a_{N}$ and $b_{N}$ homomorphisms constructed in [BG25, Definition 12.3] for every effective endotrivial $C$ and ($p$-)subgroup $H\leq G$.
Theorem 4.2.
Let $G$ be a finite $p$-group and let $C$ be an indecomposable effective endotrivial. For every subgroup $H\leq G$, there exists a chain complex homomorphism
$\iota^{H}_{C}\colon k[h_{C}(H)]\to C$
unique up to scaling, such that $\Psi^{H}(\iota^{H}_{C})$ is an isomorphism $k[h_{C}(H)]\cong\Psi^{H}(C)$ in $\operatorname{D}_{b}(k[{G}/\!\!/{H}])$. We have $\iota^{H}_{C}=\iota^{{}^{g}H}_{C}$ for all $g\in G$. Moreover, the image of $\iota^{H}_{C}$ in $C_{h_{C}(H)}$ is contained in an indecomposable direct summand isomorphic to $k[G/K]$, for some subgroup $K\geq H$.
Proof.
Since $C$ is effective, we may assume (after possibly shifting $C$) by Corollary 3.12 that there exists an indecomposable chain complex of free $k\Gamma_{G}$-modules $C^{?}$ such that $C\cong C^{1}$. Denote the differentials of $C^{?}$ by $d_{i}^{?}$. By Proposition 4.1, the chain complex of permutation k[G//H]k[{G}/\!\!/{H}]-modules $C^{H}$, which by Proposition 3.5 is isomorphic in $\mathscr{K}({G}/\!\!/{H})$ to $\Psi^{H}(C)$, contains up to homotopy a unique indecomposable k[G//H]k[{G}/\!\!/{H}]-module $M$ in top homological degree $h_{C}(H)$, and $\dim_{k}\ker(d_{h_{C}(H)}^{H})=1$. It follows by the stabilizers grow condition that in the canonical permutation basis of $C^{?}$, $M$ corresponds to a unique direct summand in homological degree $h_{C}(H)$.
Therefore, we are in the situation of Proposition 3.7 after restricting $d_{i}^{?}$ to $M$. Choosing a nonzero $m\in\ker(d_{h_{C}(H)}^{H})$ (this choice is unique up to scaling), Proposition 3.7 asserts the existence of a $m^{\prime}\in\ker(d_{h_{C}(H)}^{1})$ generating a (unique) $k$-dimension one submodule of $C^{1}_{h_{C}(H)}$. We have a unique (up to scaling) nonzero homomorphism $\iota^{H}_{C}$ of chain complexes as follows.
[…]
We have that $\Psi^{H}(\iota^{H}_{C})$ is by construction a quasi-isomorphism. The statement $\iota^{H}_{C}=\iota^{{}^{g}H}_{C}$ for all $g\in G$ follows since $c_{g}\circ\Psi^{H}=\Psi^{H^{\prime}}$ for any $g\in G$ satisfying $^{g}H=H^{\prime}$. The final statement follows from the final statement of Proposition 3.7. ∎
Remark 4.3.
One has to take care in how the maps $\iota^{H}_{C}$ are constructed; such maps that become quasi-isomorphisms locally are non-unique (not even up to choice of identification as in [BG25, Remark 12.5]). For instance, let $G=D_{16}$, $C$ be the endotrivial
$C:=kD_{16}\xrightarrow{d_{2}}k[D_{16}/H_{1}]\oplus k[D_{16}/H_{2}]\xrightarrow{d_{1}}k,$
where $H_{1}$ and $H_{2}$ are non-conjugate non-central subgroups of order 2, $d_{2}$ is induced by the assignment $g\mapsto(gH_{1},gH_{2})$, and $d_{1}$ is induced by augmentation homomorphisms. In this case, we could choose $\iota^{H_{1}}_{C}$ and $\iota^{H_{2}}_{C}$ to be the inclusion
$\iota\colon k\to k[D_{16}/H_{1}]\oplus k[D_{16}/H_{2}],\quad 1\mapsto\left(\sum_{g\in[G/H_{1}]}gH_{1},\sum_{g\in[G/H_{2}]}gH_{2}\right).$
In this case, both $\hat{\Psi}^{H_{1}}(\iota)$ and $\hat{\Psi}^{H_{2}}(\iota)$ are isomorphisms, a desired property. However, from the construction in Proposition 3.7, it follows that $\iota^{H_{1}}_{C}$ and $\iota^{H_{2}}_{C}$ satisfy $\Psi^{H_{1}}(\iota^{H_{2}}_{C})=0$ and vice versa, since neither $H_{1}\geq_{G}H_{2}$ or vice versa.
Additionally, given two subgroups $K,H\leq G$, $\iota^{H}_{C}$ and $\iota^{K}_{C}$ may coincide. For instance, any shift of the tensor unit has $\iota^{H}_{C}=\iota^{G}_{C}$ for all subgroups $H\leq G$.
Proposition 4.4.
Let $H,K$ be subgroups of $G$ and let $C$ be an effective endotrivial of $kG$-modules. Then $\iota^{H}_{C}=\iota^{K}_{C}$ if and only if there exists a subgroup $B$ of $G$ such that $h_{C}(B)=h_{C}(H)=h_{C}(K)$ and $H,K\leq_{G}B$.
Proof.
The forward implication is straightforward by the construction in Proposition 4.2 - in particular, we may choose $B$ to be the vertex of the permutation $kG$-submodule containing the image of $\iota^{H}_{C}$. Conversely, assume $h_{C}(H)=h_{C}(K)$. Suppose there exists a subgroup $B$ satisfying $h_{C}(B)=h_{C}(H)=h_{C}(K)$ and $H,K\leq_{G}B$. From the construction in Proposition 4.2, it follows that the image of $\iota^{B}_{C}$ is the unique minimal submodule of a transitive permutation module isomorphic to $k[G/B^{\prime}]$ for some $B^{\prime}$ containing $B$. It follows that $\iota^{H}_{C}$ and $\iota^{K}_{C}$ also have image contained in this permutation module, and hence also have image the unique minimal submodule of $k[G/B^{\prime}]$, as desired. ∎
4.2. The open cover
Construction 4.5.
With the notation of [BG25, Definition 12.3], given any normal subgroup $N$ of $G$ of index $p$, the maps $a_{N}$ and $b_{N}$ are examples of such morphisms (however, $c_{N}$ is not!). For instance, the endotrivial
$u_{N}=k[G/N]\to k[G/N]\to k$
(assuming $p$ odd) satisfies $a_{N}=\iota_{u_{N}}^{H}$ for any $H\not\leq N$ and $b_{N}=\iota_{u_{N}}^{H}$ for any $H\leq N$.
Given this observation, we generalize the open cover of $\operatorname{Spc}(\mathscr{K})$ presented in [BG25, Proposition 13.11] as follows. Let $\mathcal{B}(G)$ denote the subset of the canonical $\mathbb{Z}$-basis of $\operatorname{CF}_{b}(G)$ not induced from the trivial $\mathbb{R}G$-module $\mathbb{R}$, i.e. the effective endotrivials arising from irreducible real representation spheres, excluding $k[1]$. Define an open of $\operatorname{Spc}(\mathscr{K})$ by
$U(H):=\bigcap_{C\in\mathcal{B}(G)}\operatorname{open}(\iota_{C}^{H}).$
Here,
$\operatorname{open}(f):=\operatorname{open}(\operatorname{cone}(f))=\{\mathscr{P}\mid f\text{ is invertible in }\mathscr{K}(G)/\mathscr{P}\}.$
Note that in defining the opens $U(H)$, the canonical $\mathbb{Z}$-basis element $k[1]\in\mathcal{B}(G)$ would play no significant role, so we may exclude $k[1]$ from $\mathcal{B}(G)$ without issue.
Remark 4.6.
Let us recall the closed points of $\operatorname{Spc}(\mathscr{K})$. Recall that closed points are exactly the minimal primes for inclusion, and every prime contains a minimal prime. [BG25, Corollary 7.31] asserts that the minimal primes of $\operatorname{Spc}(\mathscr{K})$ are those explicitly of the form $\mathfrak{m}_{H}:=\mathscr{P}(H,0)$ for some subgroup $H$, where 0 denotes the unique closed point of $V_{{G}/\!\!/{H}}$. These closed points are precisely the kernel of the residue tt-functor $\mathbb{F}^{H}:=\operatorname{Res}_{1}^{{G}/\!\!/{H}}\circ\Psi^{H}=\Psi^{H}\circ\operatorname{Res}^{G}_{H}$ (see [BG25, Definition 7.26]).
The collection $\{U(H)\}_{H\leq G}$ is an open cover of $\operatorname{Spc}(\mathscr{K})$, with the closed point $\mathfrak{m}_{H}$ belonging to $U(H)$.
Proposition 4.7.
Let $H$ be a subgroup of $G$ and $C$ be an effective endotrivial. Recall the residue tt-functor $\mathbb{F}^{H}\colon\operatorname{Res}^{{G}/\!\!/{H}}_{1}\circ\Psi^{H}\colon\mathscr{K}\to\operatorname{D}_{b}(k)$ of the closed point $\mathfrak{m}_{H}$. $\mathbb{F}^{H}(\iota^{H}_{C})$ is an isomorphism. Moreover, if $K$ is a subgroup of $G$ satisfying $h_{C}(K)\neq h_{C}(H)$, $\mathbb{F}^{K}(\iota^{H}_{C})$ is not an isomorphism.
In particular, the open $U(H)$ contains $\mathfrak{m}_{H}$, therefore the set of opens $\{U(H)\}_{H\leq G}$ is an open cover of $\operatorname{Spc}(\mathscr{K})$.
Proof.
The fact that $\mathbb{F}^{H}(\iota^{H}_{C})$ is an isomorphism follows immediately since $\Psi^{H}(\iota^{H}_{C})$ is an isomorphism. Moreover, if $h_{C}(K)\neq h_{C}(H)$, $\mathbb{F}^{K}(\iota^{H}_{C})$ cannot be an isomorphism since $k[h_{C}(H)]$ is not isomorphic to $\Psi^{K}(C)$ in $\operatorname{D}_{b}(k[{G}/\!\!/{K}])$.
It follows by conservativity that for any effective endotrivial $C$, $\iota^{H}_{C}$ is an isomorphism in $\mathscr{K}/\mathfrak{m}_{H}$ since $\mathbb{F}^{H}$ is the residue functor of $\mathfrak{m}_{H}$, so $U(H)$ contains $\mathfrak{m}_{H}$. Therefore by general tt-geometry, $\{U(H)\}_{H\leq G}$ is an open cover of $\operatorname{Spc}(\mathscr{K})$, as every prime specializes to some $\mathfrak{m}_{H}$, which are precisely the closed points of $\operatorname{Spc}(\mathscr{K}(G)$. ∎
We next show that if $K,H$ are non-conjugate subgroups of $G$, then $\mathfrak{m}_{K}\not\in U(H)$. This takes a bit more work. It follows from the previous proposition that if there exists a Borel-Smith function $f$ for which $f(K)\neq f(H)$, then $\mathfrak{m}_{K}\not\in U(H)$. However, this may not necessarily occur!
Definition 4.8.
Let $K,H$ be a pair of non-conjugate subgroups of $G$. Say $K$ and $H$ are indistinguishable if for all Borel-Smith functions $f$, $f(K)=f(H)$. If no such indistinguishable pairs exist in $G$, say non-conjugacy is detected in $G$.
Remark 4.9.
If $K$ and $H$ are indistinguishable, we have $|K|=|H|$, since $\dim(\mathbb{R}G)^{H}=[G:H]$, the index of $H$ in $G$.
Many $p$-groups have non-conjugacy detected, such as abelian groups trivially, all finite $p$-groups of normal $p$-rank one (this is computed indirectly in [Mil24, Section 6]), and all groups of order at most $p^{3}$. However, indistinguishable pairs of subgroups exist. For instance, set $G:=C_{8}\ltimes(C_{2}\times C_{2})$, with generators $a,b,c$ satisfying
$a^{8}=b^{2}=c^{2}=1,\quad{}^{b}a=a^{-1},\quad{}^{c}a=a^{3},\quad bc=cb.$
This group has a GAP implementation of SmallGroup(32,43); it is the holomorph of $\mathbb{Z}/8\mathbb{Z}$. In this case, $C_{2}\times C_{2}$ acts on $D_{8}$ faithfully. $G$ has two nonconjugate subgroups isomorphic to $V_{4}$,
$H:=\{1,b,c,bc\},\quad K:=\{1,a^{2}b,a^{2}c,bc\},$
whose individual elements are all conjugate:
$b\sim_{G}a^{2}b,\quad c\sim_{G}a^{2}c.$
Therefore, given any real representation $V$ of $G$ with character $\mathscr{P}_{V}$, we have $\dim V^{H}=\dim V^{K}$, since
$\dim V^{H}=\frac{1}{|H|}\sum_{h\in H}\chi_{V}(h).$
Since $\operatorname{im}(\dim)=\operatorname{CF}_{b}(G)$, every Borel-Smith function $f$ satisfies $f(H)=f(K)$. The author thanks Math.SE user testaccount for this example.
Lemma 4.10.
Let $H,K$ be non-conjugate subgroups of $G$ with the same order, let $L$ be the smallest subgroup of $G$ containing both a conjugate of $H$ and a conjugate of $K$, and let $C$ be an effective endotrivial for which $h_{C}(L)<h_{C}(H)$. Then $\mathbb{F}^{K}(\iota^{H}_{C})$ and $\mathbb{F}^{H}(\iota^{K}_{C})$ are 0.
Proof.
It suffices to assume $h_{C}(H)=h_{C}(K)$. From Proposition 4.1 and the construction of $\iota^{H}_{C}$ in Theorem 4.2, the image of $\iota^{H}_{C}$ in homological degree $h_{C}(H)$ is contained in an indecomposable permutation $kG$-module $M$ with stabilizer containing $H$ but strictly contained in $L$. Therefore, $\Psi^{K}(M)=0$, so the image of $\mathbb{F}^{K}(\iota^{H}_{C})$ in degree $h_{C}(H)$ is the zero map, as desired. An analogous argument demonstrates $\mathbb{F}^{H}(\iota^{K}_{C})$ is also the zero map. ∎
Corollary 4.11.
The prime $\mathfrak{m}_{H}$ is the only closed point of $\operatorname{Spc}(\mathscr{K})$ contained in the open $U(H)$.
Proof.
We previously showed in Proposition 4.7 that $\mathfrak{m}_{H}\in U(H)$. Since $U(H)$ is defined by iterating over the irreducible endotrivials, if $h_{C}(H)\neq h_{C}(K)$ for some irreducible endotrivial $C$, then $\mathfrak{m}_{K}\not\in U(H)$. Therefore, it suffices to consider the case where $G$ has a pair of indistinguishable subgroups $H,K$. If this occurs, clearly $H,K<G$, so $H,K<_{G}L\leq G$, where $L$ is the smallest subgroup of $G$ containing both a conjugate of $H$ and $K$. Since $|L|>|H|$, there exists a Borel-Smith function $f$ for which $f(L)\neq f(H)$ (see Remark 4.9). Therefore, there exists a canonical $\mathbb{Z}$-basis element $b$ of $\operatorname{CF}_{b}(G)$ which also satisfies $b(L)\neq b(H)$. Hence, there exists an element $C$ of the canonical $\mathbb{Z}$-basis $\mathcal{B}(G)\cup\{k[1]\}$ of $\operatorname{Pic}(\mathscr{K}(G))$ with $h_{C}(H)\neq h_{C}(L)$. The previous lemma implies $\mathbb{F}^{K}(\iota^{H}_{C})$ is not an isomorphism, and it follows by definition that $\mathfrak{m}_{K}\not\in U(H)$, as desired. ∎
Warning: Corollary 4.11 does not imply that the open $U(H)$ has a unique closed point viewed as a subspace of $\operatorname{Spc}(\mathscr{K})$. The open $U(H)$ has a minimal closed point if and only if $\mathscr{K}(G)|_{U(H)}\cong\mathscr{K}(G)/\mathfrak{m}_{H}$, i.e. $U(H)$ consists of all points specializing to $\mathfrak{m}_{H}$. In general this may not occur.
Remark 4.12.
Corollary 4.11 result does not hold for the open cover of [BG25, Section 13] for non elementary abelian $p$-groups. For instance, [BG25, Proposition 13.14] states the closed complement of the open $U^{\prime}(1)$ (we write $U^{\prime}(1)$ to denote the open $U_{G}(1)$ of [BG25, Proposition 13.14] in order to distinguish the covers apart) is the support of $\operatorname{kos}_{G}(F)$, where $F$ denotes the Frattini subgroup of $G$ (i.e. the intersection of all maximal subgroups of $G$). By [BG25, Corollary 7.17], we have $\mathfrak{m}_{H}\in\operatorname{supp}(\operatorname{kos}_{G}(F))$ if and only if $H\not\leq_{G}F$, hence $\mathfrak{m}_{H}\in U^{\prime}(1)$ if and only if $H\leq_{G}F$. Therefore, $U^{\prime}(1)$ contains a unique closed point of $\mathscr{K}(G)$ if and only if $F=1$ if and only if $G$ is an elementary abelian $p$-group.
Corollary 4.13.
Every element $C\in\operatorname{Pic}(\mathscr{K}(G))$ is a line bundle under the open cover $\{U(H)\}_{H\leq G}$. In particular, for every subgroup $H\leq G$, we have an isomorphism $C\cong k[h_{C}(H)]$ in the localization $\mathscr{K}(G)|_{U(H)}$.
Proof.
By construction of $U(H)$, every irreducible endotrivial corresponding to an element of the canonical $\mathbb{Z}$-basis of $\operatorname{CF}_{b}(G)$ is isomorphic to $k[h_{C}(H)]$ in $\mathscr{K}(G)|_{U(H)}$. As the set of irreducible endotrivials is a basis of $\operatorname{Pic}(\mathscr{K}(G))$, it follows that every endotrivial $C$ is isomorphic to $k[h_{C}(H)].$
Remark 4.14.
Corollary 4.13 does not imply the existence of a homomorphism $f:k[h_{C}(H)]\to C$ which is an isomorphism in $\mathscr{K}(G)|_{U(H)}$. Such homomorphisms may not exist in general for non-effective endotrivials. For example, let $p=2$ and $G=C_{2}$, then the two-term endotrivial $C:=k\xrightarrow{\operatorname{coaug}}kC_{2}$ with $k$ in homological degree 0 satisfies $\operatorname{Hom}_{\mathscr{K}(G)}(k,C)=0$, but we have a local isomorphism $k[0]\cong C$ in the localization $\mathscr{K}(G)|_{U(C_{2})}$.
We turn to our open $U(1).$ First, we need a lemma about faithful endotrivials, i.e. endotrivials that arise from a faithful real representation $V$. Equivalently, these are endotrivials whose h-marks are all 0 for any subgroup $H$ containing a nonzero normal subgroup of $G$, see [Mil24]. In this case, we recover a generalization of [BG25, Proposition 13.14].
Proposition 4.15.
Let $C$ be an effective endotrivial arising from a faithful real representation $V$. Then $\operatorname{cone}(\iota^{1}_{C})$, generates $\mathscr{K}_{ac}(G)$ as a tt-ideal, and we have an equality
$\operatorname{supp}(\operatorname{cone}(\iota^{1}_{C}))=\operatorname{supp}(\ker(\operatorname{Res}^{G}_{1}))=\operatorname{supp}(\mathscr{K}_{ac}(G))=\operatorname{supp}(\operatorname{kos}_{G}(1)).$
Proof.
We have that $\ker(V)=1$, so by Proposition 4.1 if $n$ is the highest nonzero homological degree of $C$, $C_{n}\cong kG$. Now consider the complex $\operatorname{cone}(\iota^{1}_{C})^{*}$. After shifting, this complex satisfies $C_{1}\cong kG$ and $C_{0}\cong k$. Therefore, we are in the situation of [BG25, Corollary 3.20], and it follows that $\operatorname{cone}(\iota^{1}_{C})^{*}$, hence $\operatorname{cone}(\iota^{1}_{C})$, generates $\mathscr{K}_{ac}(G)$ as a tt-ideal, and $\operatorname{supp}(\mathscr{K}_{ac}(G))=\operatorname{supp}(\operatorname{cone}(\iota^{1}_{C}))$, as desired. ∎
Remark 4.16.
Analogously, if $V$ is a real representation with kernel $N\trianglelefteq G$, then it is a faithful real representation of $G/N$, and the associated endotrivial $C$ is a faithful endotrivial complex of $k[G/N]$-modules. It follows that in this case, $\operatorname{cone}(\iota^{1}_{C})=\operatorname{cone}(\iota^{N}_{C})$ (as they are the same map) and $\operatorname{supp}(\ker(\operatorname{Res}^{G}_{N}))=\operatorname{supp}(\operatorname{kos}_{G}(N))$, recovering [BG25, Lemma 13.2].
Theorem 4.17.
Let $G$ be a finite $p$-group. The closed complement of the open $U(1)$ is the support of $\operatorname{kos}_{G}(1)$, i.e. the closed support of the tt-ideal $\mathscr{K}_{ac}(G)=\ker(\operatorname{Res}^{G}_{1})$. In particular, $U(1)$ is equal to the cohomological open $V_{G}=\operatorname{Spc}(\operatorname{D}_{b}(kG)).$
Proof.
Since $\mathcal{B}(G)\cup\{k[1]\}$ is a basis of $\operatorname{Pic}(\mathscr{K}(G))$, for any $\mathscr{P}\in U(1)$, all endotrivials are isomorphic to a shift of the tensor unit in $\mathscr{K}(G)/\mathscr{P}$. In particular, there exists a faithful real representation $V$ (for instance the regular representation $\mathbb{R}G$) and a corresponding endotrivial $C_{V}$, for which $\iota_{C_{V}}^{1}\colon k[h_{C_{V}}(1)]\to C_{V}$ is an isomorphism in $\mathscr{K}(G)/\mathscr{P}$ (in fact, an isomorphism in $\mathscr{K}|_{U(1)}$ by Corollary 4.13). Therefore, $\operatorname{open}(\operatorname{cone}(\iota^{1}_{C_{V}}))\supseteq U(1)$. By Proposition 4.15, we have $\operatorname{supp}(\operatorname{kos}_{G}(1))=\operatorname{supp}(\operatorname{cone}(\iota_{C_{V}}^{1}))\subseteq U(1)^{c}$.
Conversely, by definition we have
$U(1)=\bigcap_{C\in\mathcal{B}(G)}\operatorname{open}(\iota^{1}_{C}).$
Therefore by Remark 4.16,
$U(1)^{c}=\bigcup_{C\in\mathcal{B}(G)}\operatorname{supp}(\operatorname{cone}(\iota^{1}_{C}))=\bigcup_{C\in\mathcal{B}(G)}\operatorname{supp}(\operatorname{kos}_{G}(\ker(C))),$
where $\ker(C)$ denotes the kernel of the real representation corresponding to $C$. For any subgroup $H\leq G$, one has by [BG25, Corollary 7.17] $\operatorname{supp}(\operatorname{kos}_{G}(H))\subseteq\operatorname{supp}(\operatorname{kos}_{G}(1))$, thus $U(1)^{c}\subseteq\operatorname{supp}(\operatorname{kos}_{G}(1))$, thus equality holds. ∎
Corollary 4.18.
We have an tt-equivalence $\mathscr{K}(G)|_{U(1)}\cong\operatorname{D}_{b}(kG)$.
Proof.
This follows immediately from Proposition 4.17 and general tt-geometry. ∎
In particular, the open $U(1)$ has a unique closed point as a subspace of $\mathscr{K}(G)$. However, we stress that in general the other opens $U(H)$ need not have unique closed points.
5. The (re-)twisted cohomology ring
We are now ready to define the twisted cohomology ring for finite $p$-groups. When $G$ is a finite elementary abelian $p$-group, we recover the previously constructed twisted cohomology ring, since in this case the set of endotrivials $\{u_{N}\}_{N\in\mathcal{N}}$ as defined in [BG25, Example 12.1] and the shift of the tensor unit $k[1]$ form precisely the canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K}(G))$.
Definition 5.1.
Let $\mathbb{N}^{\mathcal{B}(G)}=\{q\colon\mathcal{B}(G)\to\mathbb{N}\}$ be the monoid of twists, i.e. tuples of non-negative integers indexed by the $\mathcal{B}(G)$. Equivalently, the monoid of twists identifies with the submonoid of $\operatorname{Pic}(\mathscr{K}(G))$ generated by $\mathcal{B}(G)$. Consider the $(\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)})$-graded ring
$\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)=\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G;k):=\bigoplus_{s\in\mathbb{Z}}\bigoplus_{q\in\mathbb{N}^{\mathcal{B}(G)}}\operatorname{Hom}_{\mathscr{K}(G)}\left(k,\bigotimes_{C\in\mathcal{B}(G)}C^{\otimes q(C)}[s]\right).$
Its multiplication is induced by the tensor product in $\mathscr{K}(G)$. Note that only non-positive shifts $s\leq 0$ produce non-zero homomorphisms. We call $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ the twisted cohomology ring of $G$. It is convenient to write
$k(q):=\bigotimes_{C\in\mathcal{B}(G)}C^{q(C)}$
for every twist $q\in\mathbb{N}^{\mathcal{B}(G)}$ and thus abbreviate $\operatorname{H}^{s,q}(G)=\operatorname{Hom}_{\mathscr{K}(G)}(k,k(q)[s])$.
This ring is commutative for $p=2$ and graded-commutative for $p$ odd (see [BG25, Remark 12.8]) - note that only the shift $s$ plays a role in the graded commutativity, and not the twist $k(q)$.
Remark 5.2.
This construction may not immediately adapt if $G$ is not a $p$-group, since in this setting it is not known if one has a canonical $\mathbb{Z}$-basis for $\operatorname{Pic}(\mathscr{K}(G))$. A possible replacement could be the subgroup of $\mathscr{K}(G)$ of endotrivials arising from representation spheres for $G$. We propose that this subgroup has finite index in $\mathscr{K}(G),$ hence the analogous construction of an open cover should also hold.
To examine the $k$-vector space $\operatorname{H}^{s,q}(G)$, it suffices to choose an indecomposable representative due to the stabilizers grow property.
Proposition 5.3.
Let $G$ be a finite $p$-group, $k(q)=C$ an effective endotrivial, and $s$ an integer. Let $f\colon k\to C[s]$ be a nonzero chain complex homomorphism. We have that $f\in\operatorname{Hom}_{\mathscr{K}(G)}(k,C[s])=0$ if and only if $\operatorname{im}(f)\subset C_{s}$ is contained in a contractible direct summand of $C$. In particular, if $C$ is indecomposable, then
$\operatorname{H}^{s,q}(G):=\operatorname{Hom}_{\mathscr{K}(G)}(k,C[s])=\operatorname{Hom}_{\operatorname{Ch}(p-\operatorname{perm}(kG))}(k,C[s]).$
Proof.
The converse implication is straightforward. Suppose $f\in\operatorname{Hom}_{\mathscr{K}(G)}(k,C[s])=0$, then there exists a homotopy $h\colon k\to C_{s+1}$ such that $f=d_{s+1}\circ h$. Choose a canonical $\mathbb{Z}$-basis of $C$, and consider the projection of the image of $f$ onto each indecomposable summand. Necessarily on each summand $k[G/H]$ whose projection is nonzero, the image must be the unique minimal nonzero submodule $k\subseteq k[G/H]$. Since $f=d_{s+1}\circ h$, there exists an indecomposable direct summand $M$ of $C_{s+1}$ such that $p_{k[G/H]}\circ f=d_{s+1}\circ p_{M}\circ h$. Since the stabilizers grow condition holds for $C$, the only possible homomorphisms $M\to k[G/H]$ with $k\subset M$ not in the kernel are isomorphisms, hence $M\cong k[G/H]$ and we may split off the contractible summand. An induction argument demonstrates $\operatorname{im}(f)$ is contained in a contractible direct summand of $C$ as desired. ∎
Just like in [Bal10a, BG25], we have a canonical comparison map.
Proposition 5.4.
There is a continuous comparison map
$\operatorname{comp}_{G}\colon\operatorname{Spc}(\mathscr{K}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$
mapping a tt-prime $\mathscr{P}$ to the ideal generated by the homogeneous $f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ whose cone does not belong to $\mathscr{P}$. It is characterized by the fact that for all $f$,
$\operatorname{comp}_{G}^{-1}(Z(f))=\operatorname{supp}(\operatorname{cone}(f))=\{\mathscr{P}\mid f\text{ is not invertible in }\mathscr{K}(G)/\mathscr{P}\}$
where $Z(f)=\{\mathfrak{p}\mid f\in\mathfrak{p}\}$ is the closed subset of $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ defined by $f$.
Proof.
Same as [BG25, Proposition 13.4]. ∎
Recall the conservative functor of Theorem 1.8, the collection of modular fixed points functors $\{\hat{\Psi}^{H}\}_{H\leq G}$. For shorthand, we denote this functor by $\prod\hat{\Psi}:\mathscr{K}(G)\to\prod_{H\leq G}\operatorname{D}_{b}({G}/\!\!/{H})$. We’ll describe the “twisted” analogue of the induced homomorphism on the respective cohomology rings.
Definition 5.5.
Let $C\in\mathcal{B}(G)$, $s\in\mathbb{Z}$, and $f\colon k\to C[s]$ be a homogeneous element of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$. For every subgroup $H\leq G$, $\hat{\Psi}^{H}(C)\in\operatorname{Hom}_{\operatorname{D}_{b}({G}/\!\!/{H})}(k,\Psi^{H}(C)[s])$, and since we have an isomorphism $\Psi^{H}(C)[s]\cong k[h_{C}(H)+s]$ (determined by $\iota^{H}_{C}$), this determines an element in group cohomology $f_{H}\in\operatorname{H}^{h_{C}(H)+s}({G}/\!\!/{H},k)=\operatorname{Ext}^{h_{C}(H)+s}_{k[{G}/\!\!/{H}]}(k,k)$ determined by the fraction
$k\xrightarrow{\Psi^{H}(f)}C[s]\xleftarrow{\Psi^{H}(\iota^{H}_{C})}k[h_{C}(H)+s].$
We consider $\hat{\Psi}^{H}(f)$ an element of $\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H})$, i.e. as a morphism in $\operatorname{D}_{b}(k[{G}/\!\!/{H}])$. Note that for the endotrivial $u_{N}$ defined in [BG25, Definition 12.3] and $H=1$, the fraction $\hat{\Psi}^{1}(\iota^{G}_{u_{N}})$ is precisely the fraction $\zeta_{N}^{+}$ of [BG25, Remark 12.7]. If $H=G$, the fraction $\hat{\Psi}^{G}(\iota^{1}_{u_{N}})$ is precisely the fraction $\zeta_{N}^{-}$.
Performing this construction over all subgroups $H\leq G$ up to conjugacy determines an element $\left(\prod\hat{\Psi}\right)(f)$ in the $\operatorname{CF}(G)$-graded ring
$\operatorname{dH}^{{\scriptscriptstyle\bullet}}(G):=\prod_{H\in s_{p}(G)/G}\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H},k).$
For $\prod\hat{\Psi}$ to be a well-defined ring homomorphism, the denominator associated to an arbitrary effective endotrivial $C$ is defined as the product of $\iota^{H}_{C}$s corresponding to the unique tensor product factorization of $C$ into a product of irreducible endotrivials. This product of $\iota^{H}$s remains a quasi-isomorphism. Note that this product is well-defined up to reordering, since $h_{C}$ returns only even values for $p$ odd.
The assignment described describes the ring homomorphism $\prod\hat{\Psi}\colon\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{dH}^{\scriptscriptstyle\bullet}(G)$ induced by the conservative functor of [BG25, Theorem 7.2]. It is clear that this homomorphism is linear with respect to addition. Linearity over multiplication follows from the following roof; let $C$ and $D$ be endotrivials, $H$ a subgroup of $G$, $i$ and $j$ integers, and $f\colon k\to C[i]$ and $g\colon k\to D[j]$ homogeneous elements of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$. For shorthand, we write $\iota^{H}_{C}$ and $\iota^{H}_{D}$ for the products of $\iota^{H}_{C^{\prime}}$ and $\iota^{H}_{D^{\prime}}$ running over all irreducible endotrivial $C^{\prime}$ and $D^{\prime}$ occurring in the tensor decompositions of $C$ and $D$ respectively. Then, the following roof commutes.
[…]
Note that $\prod\hat{\Psi}$ sends tensor products of homomorphisms to Yoneda products of cohomology classes, sums of homomorphisms to Baer sums, and sends a homogeneous element in homological degree $(s,k(q))$ to a homogeneous element in degree $\underline{s}+h_{k(q)}$, where $\underline{s}$ denotes the constant superclass function returning only $s$.
Finally, we of course have a homeomorphism
$\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))\cong\bigsqcup_{H\in s_{p}(G)/G}V_{{G}/\!\!/{H}}.$
We denote the induced map on spectrum by $\prod\hat{\psi}\colon\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$, similar to the categorical case.
Proposition 5.6.
The kernel of $\prod\hat{\Psi}$ is $\otimes$-nilpotent, i.e. elements $f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ for which $f^{\otimes n}=0$ for some $n\geq 1$.
Proof.
This follows directly from [BG25, Theorems 7.1, 7.2], since $\prod\hat{\Psi}$ is induced from the conservative family of functors $\{\hat{\Psi}^{H}\}$. ∎
Proposition 5.7.
The following diagram in $\operatorname{Top}$ commutes.
[…]
In particular, the image of the comparison map $\operatorname{comp}_{G}$ is as a set, precisely the image of $\prod\hat{\psi}$. Moreover, the image of $\prod\hat{\psi}$ is dense in $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$.
Proof.
Every prime $\mathscr{P}(H,\mathfrak{p})\in\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))$ (in the notation of [BG25]) is explicitly the product of the prime $\mathfrak{p}$ of $\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H})$ and the full rings $\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{K})$ with $K$ different from $H$. The map $\prod\hat{\psi}$ sends the prime ideal $\mathscr{P}(H,\mathfrak{p})\in V_{{G}/\!\!/{H}}\subseteq\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))$ to the ideal $\{f\in\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))\mid\left(\prod\hat{\psi}\right)(f)\in\mathscr{P}(H,\mathfrak{p})\}$, and we have $\left(\prod\hat{\psi}\right)(f)\in\mathscr{P}(H,\mathfrak{p})$ if and only if $\hat{\Psi}^{H}(f)\in\mathfrak{p}$. Verification of commutativity now follows from the fact that given any $f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$, $H\leq G$, and $\mathfrak{p}\in V_{{G}/\!\!/{H}}$, we have $\operatorname{cone}(f)\not\in\mathscr{P}(H,\mathfrak{p})$ if and only if $\hat{\Psi}^{H}(f)\in\mathfrak{p}$. Indeed, it is well-known that the comparison map on the tt-category $\operatorname{D}_{b}({G}/\!\!/{H})$ is a homeomorphism $\operatorname{Spc}(\operatorname{D}_{b}({G}/\!\!/{H}))\xrightarrow{\cong}\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H}))$, sending a prime $\mathscr{P}(H,\mathfrak{p})$ to the prime $\mathfrak{p}=\{\zeta\mid\operatorname{cone}(\zeta)\not\in\mathfrak{p}\}\subset\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H},k)$. Thus, $\zeta\in\mathfrak{p}$ if and only if $\operatorname{cone}(\zeta)\not\in\mathscr{P}(H,\mathfrak{p})$, and setting $\zeta=\hat{\Psi}^{H}(f)$ shows the result.
For the final statement, Proposition 5.6 implies the kernel of $d$ is nilpotent, so the result follows by a standard algebraic geometry result. ∎
Remark 5.8.
In particular, the image of the prime $\mathscr{P}(H,\mathfrak{p})$ under the comparison map $\operatorname{comp}_{G}$ is equivalently
$\left(\prod\hat{\psi}\right)(\mathscr{P}(H,\mathfrak{p}))=\{f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\mid\hat{\Psi}^{H}(f)\in\mathfrak{p}\}.$
5.1. Twisted cohomology under localization and tt-functors
Next, we’ll extend the results of [BG25, Section 14] regarding localization.
Definition 5.9.
Let $H$ be a subgroup of $G$. Let $S_{H}\subset\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ be the multiplicative subset generated by all $\iota^{H}_{C}$, where $C$ runs over the irreducible endotrivials $C\in\mathcal{B}(G)$. We define a $\mathbb{Z}$-graded ring
$\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H):=\big{(}\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)[S_{H}^{-1}]\big{)}_{\text{0-twist}}$
as the twist-zero part of the localization of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ with respect to $S_{H}$. Explicitly, the homogeneous elements of $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$ consist of fractions $\frac{f}{g}$ where $f,g\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ with the same $\mathcal{B}(G)$-twist $q$, and $g$ is a product of “$H$-local isomorphisms” $g^{\prime}\colon k\to C[s]$ with $C\in\mathcal{B}(G)$. Thus, $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$ is $\mathbb{Z}$-graded by the shift only. The homological degree of $\frac{f}{g}$ is the difference sts-t between the shifts of $f$ and $g$. In particular, $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(1)\cong\operatorname{H}^{\scriptscriptstyle\bullet}(G)$, following directly from Definition 5.5.
Remark 5.10.
Recall for a morphism $f$ in $\mathscr{K}(G)$ we write
$\operatorname{open}(f):=\operatorname{open}(\operatorname{cone}(f))=\{\mathscr{P}\mid f\text{ is invertible in }\mathscr{K}(G)/\mathscr{P}\}.$
We have that $\operatorname{open}(f)$ is the preimage of $\operatorname{comp}_{G}$ of the principle open $Z(f)^{c}=\{\mathfrak{p}\mid f\not\in\mathfrak{p}\}$, and is the open locus of $\operatorname{Spc}(\mathscr{K}(G))$ where $f$ is invertible.
Construction 5.11.
We can perform the central localization (see [Bal10a]) of the whole category $\mathscr{K}(G)$
$\mathcal{L}_{G}(H):=\mathscr{K}(G)[S_{H}^{-1}].$
In fact, this localization has idempotent completion $\mathscr{K}(G)|_{U(H)}$ by the same argument as in [BG25, Construction 14.12] or [Bal10a, Theorem 3.6]. Explicitly, the category $\mathcal{L}_{G}(H)$ is the Verdier quotient of $\mathscr{K}(G)$ by the tt-ideal $\langle\{\operatorname{cone}(g)\mid g\in S_{H}\}\rangle$. It has the same objects as $\mathscr{K}(G)$ and morphisms $x\to y$ of the form $\frac{f}{g}$ where $g\colon k\to C$ belongs to $S_{H}$ for $C$ an effective endotrivial and $f\colon x\to C\otimes y$ is any morphism in $\mathscr{K}(G)$ with the same twist $C$ as the denominator $g$. Moreover, the $\mathbb{Z}$-graded endomorphism ring $\operatorname{End}^{\scriptscriptstyle\bullet}_{\mathcal{L}_{G}(H)}(k)$ of the unit in $\mathcal{L}_{G}(H)$ is the $\mathbb{Z}$-graded ring $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$.
Construction 5.12.
Twisted cohomology $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is graded over a monoid over the form $\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)}$. The ring homomorphisms induced by tt-functors or localization will be homogeneous with respect to a certain homomorphism $\gamma$ on the corresponding grading monoids.
Let $H\leq G$ be a subgroup and consider the central localization $(-)_{U(H)}\colon\mathscr{K}(G)\twoheadrightarrow\mathcal{L}_{G}(H)$. Here, the morphisms $\iota^{H}_{C}$ become isomorphisms, yielding a homomorphism on the grading
$\gamma=\gamma_{U(H)}\colon\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)}\to\mathbb{Z}$
defined by $\gamma(s,q)=s+h_{k(q)}(H)$ and we obtain a ring homomorphism
$(-)_{U(H)}\colon\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{End}^{{\scriptscriptstyle\bullet}}_{\mathcal{L}_{G}(H)}(k)=\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H),$
homogeneous with respect to the homomorphism $\gamma_{U(H)}$.
We may also consider the effects of restriction and modular fixed points on twisted cohomology. Note $\mathcal{B}(G)$ is more complicated than $\mathcal{N}\subseteq\mathcal{B}(G)$, where $\mathcal{N}$ consists of endotrivials inflated from normal subgroups of $G$ of index $p$, so in general, the shifts $\gamma$ may be more difficult to compute in the case of restriction.
Construction 5.13.
Let $H\trianglelefteq G$ be a normal subgroup. The tt-functor $\Psi^{H}\colon\mathscr{K}(G)\to\mathscr{K}(G/H)$ maps effective endotrivials with h-mark at $G$ equal to 0 to effective endotrivials at $G/H$ equal to 0. In fact, since deflation preserves irreducibility of a representation, $\Psi^{H}$ sends elements of $\mathcal{B}(G)$ to $\mathcal{B}(G/H)$.
This defines a homomorphism of graded monoids
$\gamma=\gamma_{\Psi^{H}}\colon\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)}\to\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G/H)}$
given by $\gamma(s,q)=(s,\overline{q})$ where $\overline{q}$ is given by the surjection $\mathbb{N}^{\mathcal{B}(G)}\twoheadrightarrow\mathbb{N}^{\mathcal{B}(G/H)}$ along the inclusion $\mathcal{B}(G/H)\hookrightarrow\mathcal{B}(G)$ induced from inflation. Therefore, modular fixed points defines a ring homomorphism $\Psi^{H}:\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G/H)$ homogeneous with respect to $\gamma_{\Psi^{H}}$.
Given a group homomorphism $\alpha:G^{\prime}\to G$, restriction along $\alpha$ also defines a tt-functor $\alpha^{*}:\mathscr{K}(G)\to\mathscr{K}(G^{\prime})$. This again defines a corresponding ring homomorphism $\alpha^{*}:\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G^{\prime})$ homogeneous with respect to $\gamma_{\alpha^{*}}$.
Note in this case if $\alpha$ is not surjective, then $\gamma_{\alpha^{*}}$ may not be constant on the shift $s$ and in general $\overline{q}$ will be difficult to compute. Conversely, if $\alpha$ is surjective, then $\alpha^{*}$ is simply inflation, which may be easily computed, as it is section of modular fixed points. It follows that the homomorphism $\Psi^{H}$ on twisted cohomology is split surjective.
Remark 5.14.
Suppose $F\colon\mathscr{K}(G)\to\mathscr{K}(G^{\prime})$ is a tt-functor and the induced homomorphism $F\colon\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G^{\prime})$ is homogeneous with respect to $\gamma=\gamma_{F}\colon\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)}\to\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G^{\prime})}$ (for instance, modular fixed points or restriction). Then the following square commutes, since $F(\operatorname{cone}(f))=\operatorname{cone}(F(f))$.
$\operatorname{Spc}(\mathscr{K}(G^{\prime}))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G^{\prime}))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{F^{*}}$$\scriptstyle{\operatorname{comp}_{G^{\prime}}}$$\scriptstyle{\operatorname{comp}_{G}}$$\scriptstyle{F^{*}}$
In particular, setting the functor $F$ equal to the localization functor $(-)|_{U(H)}$ obtains the following commutative square.
$\operatorname{Spc}(\mathcal{L}_{G}(H))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\operatorname{comp}_{\mathcal{L}_{G}(H)}}$$\scriptstyle{\operatorname{comp}_{G}}$
Here, the left hand vertical map is the usual twist-free comparison map from [Bal10a] for the tt-category $\mathcal{L}_{G}(H)$ and the $\otimes$-invertible $k[1]$.
We combine the functors as before to strengthen [BG25, Proposition 14.21].
Proposition 5.15.
Let $H\trianglelefteq G$ be a normal subgroup. Then we have a commutative square
$V_{G/H}=\operatorname{Spc}(\operatorname{D}_{b}(k[G/H]))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}(G/H))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\hat{\psi^{H}}}$$\scriptstyle{\operatorname{comp}_{\operatorname{D}_{b}(k[G/H])}}$$\scriptstyle{\operatorname{comp}_{G}}$
In particular, the diagonal is injective.
Proof.
The proof follows essentially the same as the proof of [BG25, Proposition 14.21], except without the added assumption that $G/H$ is elementary abelian. For this, we use Proposition 4.17 which asserts the quotient $\mathscr{K}(G/H)\to\operatorname{D}_{b}(k[G/H])$ is the central localization $(-)|_{U(1)}$ for any group $G/H$. ∎
6. Injectivity of the comparison map
The crux of injectivity of the comparison map is showing that, the diagonal of Proposition 5.15 is injective when $H$ is not necessarily normal in $G$ as well.
Theorem 6.1.
Let $G$ be a finite $p$-group. The comparison map
$\operatorname{comp}_{G}\colon\operatorname{Spc}(\mathscr{K}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$
of Proposition 5.4 is injective.
Proof.
Let $\mathscr{P}=\mathscr{P}(H_{1},\mathfrak{p})$ and $\mathscr{Q}=\mathscr{P}(H_{2},\mathfrak{q})$ in $\operatorname{Spc}(\mathscr{K}(G))$ satisfy $\operatorname{comp}_{G}(\mathscr{P})=\operatorname{comp}_{G}(\mathscr{Q})$ in $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$. This implies that $\mathscr{P}\in\operatorname{open}(f)$ if and only if $Q\in\operatorname{open}(f)$ for every $f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$. In particular, for every effective endotrivial $C\in\mathcal{B}(G)$ and subgroup $H\leq G$, we have that $\mathscr{P}\in\operatorname{open}(\iota^{H}_{C})$ if and only if $\mathscr{Q}\in\operatorname{open}(\iota^{H}_{C})$.
Suppose for contradiction that $H_{2}\not\leq_{G}H_{1}$. By Corollary 4.11, there exists an effective endotrivial $C\in\mathcal{B}(G)$ for which $\Psi^{H_{2}}(\iota_{C}^{H_{1}})$ is the zero map, and vice versa. Now, consider the map $\hat{\psi}^{H_{2}}\colon V_{{G}/\!\!/{H_{2}}}\hookrightarrow\operatorname{Spc}(\mathscr{K}(G))$. We have:
$(\hat{\psi}^{H_{2}})^{-1}(\operatorname{open}(\iota^{H_{1}}_{C}))$ $=\operatorname{open}(\operatorname{cone}(\hat{\Psi}^{H_{2}}(\iota^{H_{1}}_{C})))$
$=\operatorname{open}(0\colon k\to\hat{\Psi}^{H_{2}}(C))$
$=\emptyset$
Therefore, $V_{{G}/\!\!/{H_{2}}}\cap\operatorname{open}(\iota^{H_{1}}_{C})=\emptyset$ in $\operatorname{Spc}(\mathscr{K}(G))$. On the other hand, $V_{{G}/\!\!/{H_{1}}}\subseteq\operatorname{open}(\iota^{H_{1}}_{C})$, since $\iota^{H_{1}}_{C}$ is invertible in $\mathscr{K}(G)/\mathfrak{m}_{H_{1}}$. Therefore, $\mathscr{P}\in\operatorname{open}(\iota^{H_{1}}_{C})$ but $\mathscr{Q}\not\in\operatorname{open}(\iota^{H_{1}}_{C})$, a contradiction. Thus $H_{2}\leq_{G}H_{1}$, and by symmetry $H_{1}=_{G}H_{2}$. Set $H:=H_{1}.$
If $H\trianglelefteq G$, then we have two points $\mathfrak{p},\mathfrak{q}\in V_{{G}/\!\!/{H}}$ that go to the same image under
$V_{{G}/\!\!/{H}}\xrightarrow{\hat{\psi}^{H}}\operatorname{Spc}(\mathscr{K}(G))\xrightarrow{\operatorname{comp}_{G}}\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$
and this map is injective from Proposition 5.15. Otherwise, we have the following commutative diagram.
$V_{N_{G}(H)/H}$$\operatorname{Spc}(\mathscr{K}(N_{G}(H))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(N_{G}(H)))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\operatorname{Res}^{*}}$$\scriptstyle{\operatorname{comp}_{N_{G}(H)}}$$\scriptstyle{\operatorname{comp}_{G}}$$\scriptstyle{\operatorname{Res}^{*}}$
Here, the diagonal arrow is injective since $H\trianglelefteq N_{G}(H)$, and the top row is injective since $\operatorname{Spc}(\mathscr{K}(G))=\bigsqcup_{H\in s_{p}(G)/G}V_{{G}/\!\!/{H}}$, and the composition of the above maps is precisely $\hat{\psi}^{H}$. Therefore, to show injectivity, it suffices to show the bottom composition of maps
$V_{N_{G}(H)/H}\hookrightarrow\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(N_{G}(H)))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$
is injective. We establish this as a separate lemma.
Lemma 6.2.
Let $H$ be a subgroup of $G$. The composition of continuous maps
$V_{N_{G}(H)/H}\hookrightarrow\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(N_{G}(H)))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$
is injective.
Proof.
Since $G$ is a finite $p$-group, we may choose a subnormal sequence $N_{G}(H)=H_{0}\triangleleft H_{1}\triangleleft\cdots\triangleleft H_{n}=G$ with $H_{i+1}/H_{i}=C_{p}$. We have the following setup depicted by the commutative diagram below.
$V_{N_{G}(H)/H}$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{0}))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{1}))$$\cdots$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{n}))$}
Here, the horizontal arrows are induced by restriction and the downwards arrows are $\hat{\psi}^{H}$. We will prove this statement inductively by assuming that the map $V_{N_{G}(H)}\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))$ is injective and showing that the composition $V_{N_{G}(H)}\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1}))$ is injective, which then of course implies the next map $V_{N_{G}(H)}\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1}))$ is injective as well. Injectivity of the base case $H_{0}=N_{G}(H)$ is established by Proposition 5.15.
Let $\mathfrak{p},\mathfrak{q}\in\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))$. We claim that $\operatorname{Spec}(\operatorname{Res}^{H_{i+1}}_{H_{i}})(\mathfrak{p})=\operatorname{Spec}(\operatorname{Res}^{H_{i+1}}_{H_{i}})(\mathfrak{q})$ if and only if the following property $(\ast)$ is satisfied:
$(\ast)$  For every $f\in\mathfrak{p}$, there exists a $g\in H_{i+1}$ such that $^{g}f\in\mathfrak{q}$.
The converse is straightforward. Suppose there exists an $f\colon k\to C[s]\in\mathfrak{p}$ for which $^{g}f\not\in\mathfrak{q}$ for all $g\in H_{i+1}$. Then, the trace product of $H_{i+1}$-conjugates
$f^{\prime}:=\prod_{g\in H_{i+1}/H_{i+1}}{}^{g}f$
is a morphism $f^{\prime}\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1})$. Indeed, the tensor product of chain complexes
[…]
has $kH_{i+1}$-module structure (one may see for instance that $C^{\prime}$ is the restriction of the tensor induced complex $\operatorname{Ten}^{H_{i+1}}_{H_{i}}C$ via the Mackey formula - while tensor induction is not defined in general up to homotopy, in this case the restriction is well-behaved up to homotopy), and the image of $k$ in $C^{\prime}$ is $H_{i+1}$-stable. By primality of $\mathfrak{p}$ and $\mathfrak{q}$, we have $f^{\prime}\in\mathfrak{p}$ but $f^{\prime}\not\in\mathfrak{q}$, as desired.
Now, we show that if $\mathfrak{p},\mathfrak{q}\in\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))$ satisfy $(\ast)$ and belong to the image of $V_{N_{G}(H)/H}$, then $\mathfrak{p}=\mathfrak{q}$. The statement is trivial if $H_{i+1}/H_{i}$ acts on $H_{i}$ trivially, so assume the action is nontrivial. Recall from Proposition 5.7 that in this case, $\mathfrak{p}=\rho(H,\overline{\mathfrak{p}})$ with $\overline{\mathfrak{p}}\in V_{N_{G}(H)/H}$, and we have $\mathfrak{p}=\{f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i})\mid\hat{\Psi}^{H}(f)\in\overline{\mathfrak{p}}\}$. Let $f\in\mathfrak{p}$. By Lemma 4.10, there exists an effective endotrivial chain complex $C$ of $kH_{i}$-modules and morphism $\iota^{H}_{C}$ such that $\hat{\Psi}^{H}(\iota^{H}_{C})=\operatorname{id}_{k}$ and $\hat{\Psi}^{H}({}^{g}(\iota^{H}_{C}))=0$ for all $g\in H_{i+1}\setminus H_{i}$. Therefore, $\iota^{H}_{C}\not\in\mathfrak{p}$ (since $\overline{\mathfrak{p}}\neq V_{N_{G}(H)/H}$), but for all $g\in H_{i+1}\setminus H_{i}$, we have $^{g}(\iota^{H}_{C})\in\varphi(H,0)\subseteq\mathfrak{p}$. We construct a morphism $f^{\prime\prime}\in\mathfrak{p}$ as follows,
$f^{\prime\prime}:=f+\left(\sum_{g\in[(H_{i+1}/H_{i})-H_{i}]}{}^{g}(\iota^{H}_{C})\cdot(\operatorname{id}_{k}-f)\right)\in\mathfrak{p}.$
By construction $\hat{\Psi}^{({}^{g}H)}(f^{\prime\prime})$ is an isomorphism for all $g\in H_{i+1}\setminus H_{i}$, therefore $^{g}(f^{\prime\prime})\not\in\mathfrak{q}$. But since $\mathfrak{p},\mathfrak{q}$ satisfy $(\ast)$, it follows that $f^{\prime\prime}\in\mathfrak{q}$ as well. Since each $^{g}(\iota^{H}_{C})\in\varphi(H,0)\subseteq\mathfrak{q}$, cancellation implies $f\in\mathfrak{q}$ as well. Thus $\mathfrak{p}\subseteq\mathfrak{q}$, and symmetry implies $\mathfrak{p}=\mathfrak{q}$, and the composition $V_{N_{G}(H)}\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1}))$ is injective, as desired. ∎
As an immediate corollary, we obtain a strengthening of Proposition 5.15.
Corollary 6.3.
Let $H$ be any subgroup of $G$. Then we have a commutative square
$V_{{G}/\!\!/{H}}=\operatorname{Spc}(\operatorname{D}_{b}(k[{G}/\!\!/{H}]))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H}))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\hat{\psi^{H}}}$$\scriptstyle{\operatorname{comp}_{\operatorname{D}_{b}(k[{G}/\!\!/{H}])}}$$\scriptstyle{\operatorname{comp}_{G}}$
In particular, the diagonal is injective, and the product
$\prod\hat{\psi}\colon\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$
is injective.
Corollary 6.4.
If $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian, then for any $H\leq G$, the comparison map restricts to a homeomorphism
$\operatorname{comp}_{G}\colon U(H)\cong\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)).$
In particular, $\operatorname{comp}_{G}\colon\operatorname{Spc}(\mathscr{K}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ is an open immersion, with image the open subspace of $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ with closed points (maximal primes)
$\operatorname{comp}_{G}(\mathfrak{m}_{H})=\{f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\mid\Psi^{H}(f)\text{ is not a quasi-isomorphism}\}.$
Proof.
We have already an injective map by Remark 5.14 and Theorem 6.1
$\operatorname{comp}_{\mathscr{L}_{G}(H)}:\operatorname{Spc}(\mathscr{L}_{G}(H))=U(H)\to\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H))).$
Noetherianity of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ implies $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$ is noetherian as well, hence [Bal10a, Theorem 7.3] implies $\operatorname{comp}_{\mathscr{L}_{G}(H)}$ is a continuous bijection. It remains to show this map is a homeomorphism; we show it is closed. Let $\mathscr{L}_{G}(H)_{\langle k\rangle}$ denote the unitization of $\mathscr{L}_{G}(H)$, i.e. the tt-subcategory compactly generated by the tensor unit. Of course, the unitization has the same cohomology ring $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$. Then [San25, Corollary 10.2] asserts that the inclusion $\mathscr{L}_{G}(H)_{\langle k\rangle}\to\mathscr{L}_{G}(H)$ induces a homeomorphism $\operatorname{Spc}(\mathscr{L}_{G}(H)_{\langle k\rangle})\cong\operatorname{Spc}(\mathscr{L}_{G}(H))$. The comparison map factors through $\operatorname{Spc}(\mathscr{L}_{G}(H))_{\langle k\rangle}$ as follows:
$\operatorname{Spc}(\mathscr{L}_{G}(H))\to\operatorname{Spc}(\mathscr{L}_{G}(H))_{\langle k\rangle}\to\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)).$
Now, $\mathscr{L}_{G}(H)_{\langle k\rangle}$ is (by definition) generated by its tensor unit. By noetherianity of $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$, $\mathscr{L}_{G}(H)_{\langle k\rangle}$ is therefore End-finite in the sense of [Lau23, Definition 2.6], hence [Lau23, Proposition 2.7] asserts the induced map on spectra $\operatorname{Spc}(\mathscr{L}_{G}(H))_{\langle k\rangle}\to\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H))$ is a homeomorphism, thus $\operatorname{comp}_{\mathscr{L}_{G}(H)}:U(H)\to\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H))$ is a homeomorphism as well, as desired.
Therefore, $\operatorname{comp}_{G}$ is a homeomorphism onto its image in $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$. It is easy to check the closed points $\mathfrak{m}_{H}$ are mapped to the corresponding maximal primes in the image as described in the theorem statement. ∎
The following corollary is immediate.
Corollary 6.5.
Let $\mathcal{O}_{G}^{\scriptscriptstyle\bullet}$ denote the sheaf of $\mathbb{Z}$-graded rings on $\operatorname{Spc}(\mathscr{K}(G))$ obtained by sheafifying $U\mapsto\operatorname{End}^{\scriptscriptstyle\bullet}_{\mathscr{K}(G)|_{U}}(k)$. If $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian, $(\operatorname{Spc}(\mathscr{K}(G)),\mathcal{O}_{G}^{\scriptscriptstyle\bullet})$ is a Dirac scheme.
Remark 6.6.
We informally say some words about noetherianity to close, to hopefully conjure hope within the reader that for any finite $p$-group $G$, $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian. First, one can show $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian when $G$ satisfies the following property: the indecomposable representative of each irreducible endotrivial $C$ in $\mathcal{B}(G)$ has at most one indecomposable module in every homological degree. This holds for instance (and possibly only when) $G$ is abelian, where the endotrivials correspond to subgroups $N\leq G$ such that $G/N$ is cyclic (see [Mil24, Section 6]). With this property, the proof follows essentially identically as [BG25, Lemma 12.12], with that the maps $\iota^{1}_{C}$ and $\iota^{G}_{C}$ replacing the maps $b_{N}$ and $a_{N}$ respectively for each irreducible endotrivial $C\in\mathcal{B}(G)$.
One crux of the matter is that if an irreducible endotrivial has a homological degree with two or more indecomposables, the exact sequence [BG25, 12.14] sends a term $f\in\operatorname{Hom}_{\mathscr{K}(G)}(k,v\otimes u^{\otimes q})$ to a direct sum of morphisms $(f_{1},\dots,f_{n})$ with $f_{i}\in\operatorname{Hom}_{\mathscr{K}(H_{i})}(k,\operatorname{Res}^{G}_{H_{i}}(v)[q])$, where the subgroups $H_{i}\leq G$ are the corresponding stabilizers of the indecomposable permutation modules. It is unclear if one can inductively lift this sum of morphisms to a morphism in $\operatorname{Hom}_{\mathscr{K}(G)}(1,v\otimes u^{\otimes q})$, polynomial in morphisms $k\to C[s]$ with $C\in\mathcal{B}(G)$, such that its image in every $\operatorname{Hom}_{\mathscr{K}(H_{i})}(k,\operatorname{Res}^{G}_{H_{i}}(v)[q])$ corresponds to the image of $f$. Additionally, the lengths of the irreducible endotrivials $C$ can be arbitrarily long, but an inductive argument may be used to resolve this issue.
A potential roadmap to proving noetherianity modulo nilpotents is to verify that the image of the map $\prod\hat{\Psi}:\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\prod_{H\in s_{p}(G)/G}\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H})$ is finite, as the kernel of $\prod\hat{\Psi}$ is nilpotent (Proposition 5.6). One can further reduce this to the elementary abelian case by postcomposing by restriction to all elementary abelian subquotients of $G$, as the restriction $\operatorname{H}^{\scriptscriptstyle\bullet}(G)\to\prod_{E\in\operatorname{elemab}}(G)\operatorname{H}^{\scriptscriptstyle\bullet}(E)$ detects nilpotents as well (see [Ben98b, Proposition 5.2.2]). It is not too unreasonable to expect this to hold - for instance [Eve61, Theorem 7.1] states that for any subgroup $H\leq G$, $\operatorname{H}^{\scriptscriptstyle\bullet}(H)$ is finite in $\operatorname{H}^{\scriptscriptstyle\bullet}(G)$ via restriction. Hence $\operatorname{H}^{\scriptscriptstyle\bullet}(H)^{G}$ (i.e. the image of restriction, see [Ben98a, Proposition 3.8.2]), is finite in $\operatorname{H}^{\scriptscriptstyle\bullet}(G)$. For this reason, it is quite pertinent and interesting to determine the image of $\hat{\Psi}^{H}:\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H})$ for any subgroup $H\leq G$. We note for instance that one can show the image of $\prod\hat{\Psi}$ is finite when $p=2$ and $G=Q_{8}$ exploiting the fact that every subgroup of $Q_{8}$ is normal. However, it is not clear if an analogue of [Bal10a, Theorem 7.3] under the hypothesis that the cohomology ring is noetherian modulo nilpotents can be utilized.

### v3

@@
 Abstract.
-We generalize Balmer and Gallauer’s (permutation) twisted cohomology ring, working toward an alternative method of deducing the Balmer spectrum of the derived category of permutation modules for any finite $p$-group.
-The construction comes equipped with a canonical comparison map from the Balmer spectrum to the homogeneous spectrum of the twisted cohomology ring.
-We show the comparison map is injective for any finite $p$-group and furthermore, an open immersion when the twisted cohomology ring is noetherian.
-For elementary abelian $p$-groups, the twisted cohomology ring coincides with Balmer and Gallauer’s original construction.
-To perform this construction, we utilize endotrivial complexes (i.e.
-the invertible objects of the derived category of permutation modules) arising up to a shift from Bredon homology of representation spheres.
-This topological structure allows us to construct certain $p$-local isomorphisms, from which we build a refined open cover of the Balmer spectrum indexed by conjugacy classes of subgroups of $G$.
-Under this open cover, every endotrivial is isomorphic to a shift of the tensor unit in each localization, thus verifying that all endotrivials are line bundles.
-When the twisted cohomology ring is noetherian, this open cover endows the Balmer spectrum with Dirac scheme structure.
+For each endotrivial complex arising from Bredon homology of a representation sphere, we construct $p$-local quasi-isomorphisms, called forerunners, enabling us to extend Balmer–Gallauer’s results in [BG25, Part II] concerning the tensor-triangular geometry of permutation modules for elementary abelian $p$-groups to all $p$-groups.
+We construct an open cover of the Balmer spectrum under which all endotrivials are line bundles, that is, every endotrivial is locally isomorphic to a shifted tensor unit.
+We define a remixed permutation twisted cohomology ring for which the canonical comparison map from the Balmer spectrum to the homogeneous spectrum of the twisted cohomology ring is injective.
+If the twisted cohomology ring is Noetherian, the comparison map is an open immersion, and the open cover endows the Balmer spectrum with Dirac scheme structure.
+We prove Noetherianity holds for Dedekind groups and all $p$-groups of order at most $p^{3}$, and conjecture Noetherianity holds for all finite $p$-groups.
 Key words and phrases:
-permutation module, twisted cohomology, comparison map, Balmer spectrum, tensor-triangular geometry
+permutation module, twisted cohomology, comparison map, Balmer spectrum, tensor-triangular geometry, endotrivial
 2020 Mathematics Subject Classification:
 18G80, 18G90, 18M05, 20C20, 20J05
+For the duration of this paper, $k$ denotes a field of prime characteristic $p$ and $G$ denotes a finite group.
 Introduction
-Let $k$ be a field of prime characteristic $p$ and $G$ be a finite group.
-Permutation $kG$-modules (modules which admit a $G$-stable $k$-basis), and their direct summands, $p$-permutation kGkG-modules, are perhaps the second easiest class of representations of $G$ over $k$ to define.
-In a wildly vast representation-theoretic cosmos of $kG$-modules, our $p$-permutation modules may seem tiny and insignificant, drowning in a sea of symmetries.
-However, our protagonists, or more precisely, the bounded homotopy category of $kG$-modules $\mathscr{K}(G):=\operatorname{K}_{b}(p\operatorname{-perm}(kG))$, exhibits significant control over the representation theory of $kG$ in general.
-For instance, if $G$ has an abelian Sylow $p$-subgroup $S$, Broué’s abelian defect group conjecture predicts a triangulated equivalence $\operatorname{K}_{b}(p\operatorname{-perm}(kGb_{0}))\cong\operatorname{K}_{b}(p\operatorname{-perm}(kN_{G}(S)c_{0}))$ where $b_{0}$ and $c_{0}$ denote the principal blocks of $kG$ and $kN_{G}(S)$ respectively.
-Perhaps more surprisingly, Balmer and Gallauer recently demonstrated that $\mathscr{K}(G)$ admits the bounded derived category $\operatorname{D}_{b}(\text{mod}(kG))$ as a Verdier quotient [BG23a], strengthening an unpublished result of Rouquier [Rou06].
-Although these $p$-permutation are few, they are mighty; together, they recover the entire representation theory of $kG$.
-If the reader is less representation-theoretically inclined, fear not!
-Our $p$-permutation category $\mathscr{K}(G)$ is the compact part of a “big” (i.e.
-rigidly-compactly generated) tt-category $\mathscr{D}(G):=\operatorname{DPerm}(G;k)$, the derived category of permutation modules.
-This category $\mathscr{D}(G)$ is equivalent to numerous big tt-categories arising from seemingly disparate areas of mathematics, such as:
+It is oft claimed that permutation $kG$-modules, modules which admit a $G$-stable $k$-basis, and their direct summands, $p$-permutation kGkG-modules, are at minimum the second easiest class of representations of $G$ over $k$ to define.
+In a wildly (c.f.
+[BD77]) vast representation-theoretic cosmos of $kG$-modules, our $p$-permutation modules appear tiny and insignificant, drowning in the endless sea of symmetries.
+However, our protagonist, the bounded homotopy category of $kG$-modules $\mathscr{K}(G):=\operatorname{K}_{b}(p\operatorname{-perm}(kG))$, exhibits significant control over said cosmos.
+For a classical example, if $G$ has an abelian Sylow $p$-subgroup $S$, Broué’s abelian defect group conjecture [BRO90, RIC96] predicts a particular triangulated equivalence $\operatorname{K}_{b}(p\operatorname{-perm}(kGb_{0}))\cong\operatorname{K}_{b}(p\operatorname{-perm}(kN_{G}(S)c_{0}))$ where $b_{0}$ and $c_{0}$ denote the principal blocks of $kG$ and $kN_{G}(S)$ respectively; such an equivalence induces a corresponding derived equivalence.
+More recently, Balmer–Gallauer demonstrated that every $kG$-module has a finite resolution by $p$-permutation modules [BG23a].
+Consequently, $\mathscr{K}(G)$ admits the bounded derived category $\operatorname{D}_{b}(\text{mod}(kG))$ as a Verdier quotient, strengthening an unpublished result of Rouquier [ROU06].
+Although these $p$-permutation are few, they are mighty; together, they recover the entire representation theory of kGkG!
+In a landmark series of papers [BG22, BG23a, BG23b, BG25], Balmer–Gallauer elucidated the tensor-triangular geometry of $\mathscr{K}(G)$.
+In particular, in [BG25], the authors deduced its Balmer spectrum $\operatorname{Spc}(\mathscr{K}(G))$, the universal support of the category.
+[^1] This classification has ramifications beyond representation theory: to the geometrically or topologically inclined reader, this classification can be interpreted in terms of cohomological Mackey functors, Artin motives, or the homotopy category of modules over the constant Green functor, see [BG23b, FUH25] As a set, the Balmer spectrum is built from $p$-local cohomology:
+$\operatorname{Spc}(\mathscr{K}(G))=\bigsqcup_{H\in s_{p}(G)/G}\operatorname{Spc}(\operatorname{D}_{b}(k[{G}/\!\!/{H}]))\cong\bigsqcup_{H\in s_{p}(G)/G}\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H};k)).$
+This stratification is obtained from certain tt-functors classically known as Brauer quotients for modules, and in [BG25] are termed modular fixed points, functors $\Psi^{H}\colon\mathscr{K}(G)\to\mathscr{K}({G}/\!\!/{H})$, where ${G}/\!\!/{H}:=N_{G}(H)/H$ is the Weyl group of $G$ at $H$.
+These functors mimic the fixed-point operation for $G$-sets: given any $G$-set $X$ and $p$-subgroup $H\leq G$, one has a natural isomorphism $\Psi^{H}(kX)\cong k[X^{H}]$.
+We’ll discuss the nitty-gritty of the classification, but first take a narrative detour to the domain of endotrivials.
+Endotrivial complexes and forerunners
+Separately, in a sequence of papers [MIL24, MIL25a, MIL25b], we deduced the Picard group $\operatorname{Pic}(\mathscr{K}(G))$, i.e., the group of endotrivial complexes for all finite groups.
+Endotrivials are by definition the invertible objects of $\mathscr{K}(G)$, the chain complexes $C\in\mathscr{K}(G)$ satisfying $C^{*}\otimes C\simeq k[0]$.
+This classification was also partially determined in Bachmann’s dissertation [BAC16] in the context of Artin motives.
+Independently, a similar computation was performed by Yalçin [YAL17] for $p$-groups in the context of $G$-Moore spaces, which was partially extended to finite groups in joint work with Gelvin [GY21].
+There are multiple alternative ways to think about endotrivial that lead to a classification.
+In [MIL25a, MIL25b], we deduced $\operatorname{Pic}(\mathscr{K}(G))$ by considering endotrivials as a special class of more generally “relatively” endotrivial complexes akin to endopermutation modules (see [BOU10, Chapter 12]); this parallel to endopermutation modules is developed further in [MIL26].
+Another characterization considered by Bachmann and Gelvin–Yalçin is topological in nature: endotrivials arise from Bredon homology of (virtual) representation spheres, i.e., one-point compactifications of real representations.
+In other words, one has a homomorphism $\operatorname{RO}(G)\to\operatorname{Pic}(\mathscr{K}(G))$.
+In fact, for $p$-groups (but not in general), the homomorphism is surjective, and its kernel is characterized by the Adams operation or the $j$-homomorphism.
+In particular, for $G$ a $p$-group, $\operatorname{Pic}(\mathscr{K}(G))$ has a canonical $\mathbb{Z}$-basis associated to the irreducible real representations of $G$.
+In this paper, we adopt this topological point of view for $p$-groups, with a particular emphasis on effective endotrivials, the endotrivials arising from Bredon homology of (genuine) representation spheres.
+Such endotrivials have additional structure: they are chain complexes of free modules over the orbit category $\Gamma_{G}$.
+This structure allows us to construct certain $p$-local quasi-isomorphisms which we term forerunners, as they foretell quasi-isomorphisms prior to applying modular fixed points.
+Theorem A.
+Let $G$ be a finite $p$-group.
+For every effective endotrivial complex (see Definition 2.9) and subgroup $H\leq G$, there exists a forerunner homomorphism $\iota^{H}_{C}:k\to C[h_{C}(H)]$ such that $\Psi^{H}(\iota^{H}_{C})$ is a quasi-isomorphism (Theorem 4.3).
+Here, $h_{C}$ denotes the h-mark homomorphism of $C$, see Definition 2.1.
+Let us justify to the inquisitive reader why the existence of forerunners is noteworthy.
+First, the existence of such morphisms is by no means a given: it is a special property of effective endotrivials.
+Proving their existence requires developing the theory of chain complexes of permutation $kG$-modules arising from the orbit category $\Gamma_{G}$; we do not list the results here, but believe they will be of independent interest.
+Second, Gallauer has already utilized these morphisms in his recent initiation of the study of periodicity for symmetric monoidal $\infty$-categories [GAL25]; we remark on this work in the sequel.
+But the real justification, the true power of these morphisms, is that they enable us to explain the geometry of permutation modules in a much more general light, as we’ll now explain.
+A new open cover of the spectrum
+In [BG25], Balmer–Gallauer split their deduction of $\operatorname{Spc}(\mathscr{K})$ into two parts: Part I concerns deducing $\operatorname{Spc}(\mathscr{K}(G))$ as a set and constructing the modular fixed-point functors, and Part II deduces the topology of $\operatorname{Spc}(\mathscr{K}(G))$.
+In Part II, the authors prove a colimit theorem a la Quillen stratification reducing classification to the case of $G$ elementary abelian, then investigate this case in greater depth.
+Using the classification of endotrivials, the existence of forerunners, and the topological origins of effective endotrivials, we extend these investigations in the context of any finite $p$-group.
+Our hope is that $\operatorname{Spc}(\mathscr{K}(G))$ can be completely deduced without any gluing.
+The $p$-group case is the crux: given a finite group $G$ with $p$-Sylow $S$, one can determine $\operatorname{Spc}(\mathscr{K}(G))$ from $\operatorname{Spc}(\mathscr{K}(S))$ by restriction via [BG25, Propositions 3.8, 4.7].
+We first review Balmer–Gallauer’s methods and results.
+Let $E$ be an elementary abelian $p$-group.
+Balmer–Gallauer constructed an open cover $\{U(H)\}_{H\leq E}$ of $\operatorname{Spc}(\mathscr{K}(E))$ indexed by subgroups of $E$.
+Explicitly, these opens are obtained by considering the open loci of certain morphisms from the tensor unit to certain endotrivials.
+This cover has some remarkable features:
   1. (a)
-    the derived category of cohomological $k$-linear Mackey functors over $G$;
+    All endotrivials for $E$ are isomorphic to a shift of the tensor unit in the localization $\mathscr{K}(E)|_{U(H)}$; in other words, all endotrivials for are line bundles [BG25, Remark 13.10];
   2. (b)
-    the homotopy category of modules over the constant Green functor $\operatorname{H}\underline{k}$ in genuine $G$-spectra;
+    Each open $U(H)$ contains the closed point $\mathscr{M}(H)\in\operatorname{Spc}(\mathscr{K}(E))$.
+Hence $\{U(H)\}_{H\leq E}$ is an open cover of $\operatorname{Spc}(\mathscr{K}(E))$; [BG25, Proposition 13.11]
   3. (c)
-    the category of $k$-linear Artin motives generated by motives of intermediate fields in a Galois extension with Galois group $G$.
-See [BG23b] for an overview of these correspondences; we remark Fuhrmann recently ascended equivalence (b) to the $\infty$-categorical level [Fuh25].
-Hence any results concerning $\mathscr{D}(G)$ and $\mathscr{K}(G)$ translate over to the aforementioned categories and their compact parts respectively.
-Continuing their foray into the land of permutation modules, Balmer-Gallauer deduced the Balmer spectrum $\operatorname{Spc}(\mathscr{K}(G))$ of $\mathscr{K}(G)$ in [BG25], a landmark achievement.
-Their construction of permutation twisted cohomology, a modification of the usual cohomology ring, plays a key role in this classification.
-The authors consider morphisms from the tensor unit to shifts of certain invertible objects (i.e.
-endotrivial complexes) corresponding to subgroups $N\triangleleft G$ of index $p$.
-The collection of these morphisms forms a multigraded ring $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$, the twisted cohomology ring, extending the usual notion of a cohomology ring of a triangulated category, and comes with a canonical comparison map,
+    Under this open cover, there is an identification $\mathscr{K}(E)|_{U(1)}\cong\operatorname{D}_{b}(kE)$.
+In other words, $U(1)$ is nothing more than the cohomological open, corresponding to the image under $\operatorname{Spc}$ of the localization $\mathscr{K}(E)\twoheadrightarrow\operatorname{D}_{b}(kE)$; [BG25, Proposition 13.14]
+This open cover is later used for to endow $\operatorname{Spc}(\mathscr{K}(E))$ with Dirac scheme structure in the sense of [HP23]; a Dirac scheme is to a usual scheme as a graded ring is to a non-graded one.
+The morphisms used to construct these opens are examples of forerunners - this observation suggests that such an open cover in fact exists in greater generality.
+Theorem B.
+Let $G$ be a finite $p$-group. There exists an open cover $\{U(H)\}_{H\in s_{p}(G)/G}$ (4.6) indexed by conjugacy classes of subgroups of $G$ for which the following holds.
+  1. (a)
+    Every endotrivial $C$ is a line bundle, i.e., over each open $U(H)$ is isomorphic to a shift of the tensor unit in the localization $\mathscr{K}(G)|_{U(H)}$.
+Explicitly, $C\cong k[h_{C}(H)]$ in $\mathscr{K}(G)|_{U(H)}$, where $h_{C}$ denotes the h-mark homomorphism associated to $C$; (Theorem 4.13)
+  2. (b)
+    Each open $U(H)$ contains a unique closed point $\mathscr{M}(H)\in\operatorname{Spc}(\mathscr{K}(G))$, corresponding to the unique closed point of $\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H};k))$; (Theorem 4.12)
+  3. (c)
+    The open $U(1)$ is the cohomological open $\operatorname{Spc}(\operatorname{D}_{b}(kG))\subseteq\operatorname{Spc}(\mathscr{K}(G))$.
+Therefore, we have a tt-equivalence $\mathscr{K}(G)|_{U(1)}\cong\operatorname{D}_{b}(kG)$.
+(Theorem 4.19)
+(Remixed) permutation twisted cohomology and Dirac geometry
+To completely deduce the topology of $\operatorname{Spc}(\mathscr{K}(E))$, Balmer–Gallauer define the permutation twisted cohomology ring $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$.
+Explicitly, the authors consider morphisms from the tensor unit to shifts of certain endotrivials corresponding to subgroups $N\triangleleft G$ of index $p$.
+The collection of these morphisms defines the multigraded ring $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$, and extends the usual notion of the cohomology ring of a triangulated category.
+The ring comes with a canonical comparison map,
 $\operatorname{comp}_{G}\colon\operatorname{Spc}(\mathscr{K}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)),$
-which was previously constructed in [Bal10a].
-A twist of the cohomology ring is necessary - since we are considering a homotopy category, the usual cohomology ring $\operatorname{End}^{\scriptscriptstyle\bullet}_{\mathscr{K}(G)}(k)$ is a tad drab on its own.
+extending the usual comparison map of [BAL10a].
+A twisted variant of the cohomology ring is necessary - since $\mathscr{K}(G)$ is a homotopy category of chain complexes, the usual cohomology ring $\operatorname{End}^{\scriptscriptstyle\bullet}_{\mathscr{K}(G)}(k)$ is a tad drab on its own.
 The miracle of this construction is:
 Theorem.
-[BG25, Theorem 10.5] Let $E$ be a finite elementary abelian $p$-group.
+[BG25, Theorem 10.5] Let $E$ be an elementary abelian $p$-group.
 The comparison map $\operatorname{comp}_{E}\colon\operatorname{Spc}(\mathscr{K}(E))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(E))$ identifies $\operatorname{Spc}(\mathscr{K}(E))$ with an open subspace of the homogeneous spectrum of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(E)$.
-This remarkable fact, along with a colimit theorem [BG25, Theorem 11.10] reducing to the case of $G$ elementary abelian á la Quillen stratification, completes the description of $\operatorname{Spc}(\mathscr{K}(G))$ as a topological space, and as a Dirac scheme in the sense of [HP23].
-As a set, $\operatorname{Spc}(\mathscr{K}(G))\cong\bigsqcup_{H\in s_{p}(G)/G}\operatorname{Spc}(\operatorname{D}_{b}(kG))$ ([BG25, Theorem 7.16]); we set $V_{G}:=\operatorname{Spc}(\operatorname{D}_{b}(kG))$.
-Hausmann and Schwede further considered the twisted cohomology ring for elementary abelian $2$-groups [HS25] (in the Mackey functor setting), which they entitled “Representation-graded Bredon homology.” The authors determined a minimal generating set and proved that for $p=2$, $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(E)$ is nilpotent-free.
-Separately, the author deduced the Picard group of $\mathscr{K}(G)$, i.e.
-the group of endotrivials, in a sequence of papers [Mil24, Mil25b, Mil25a].
-Significant parts of the classification were also completed by Bachmann in his dissertation [Bac16] in the context of Artin motives.
-When $G$ is a finite $p$-group, morally the endotrivials arise from Bredon homology of representation spheres, and therefore have additional topological structure.
-This was also observed by Yalçin in the context of $G$-Moore spaces [Yal17], where an analogous classification result was determined.
-Therefore, one obtains a canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K}(G))$ associated to real representations modulo Adams operations.
-Interestingly, for elementary abelian $p$-groups, this basis corresponds exactly to the endotrivials used by Balmer-Gallauer in their twisted cohomology ring.
-The natural question for one to ask is:
-Question.
-Can we recover the Balmer spectrum $\operatorname{Spc}(\mathscr{K}(G))$ by further twisting (or if you will, re-mixing) $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ by the endotrivial complexes arising (up to shifts) from representation spheres?
-We consider twisting by shifts of complexes coming from genuine $\mathbb{R}G$-modules as opposed to virtual representations, as if we were to twist by all endotrivials, the graded ring becomes unmanageably large (c.f.
-[BG25, Remark 12.24]).
-The analogous comparison to make here is that the group cohomology ring $\operatorname{H}^{\scriptscriptstyle\bullet}(G;k)$ is noetherian and recovers the Balmer spectrum of the derived category $\operatorname{D}_{b}(kG)$, whereas the Tate cohomology ring is too large.
-We call such endotrivials effective, and their corresponding h-mark homomorphisms are monotonically decreasing superclass functions.
-Similarly, Hausmann-Schwede call the corresponding “quadrant” of twisted cohomology associated to twists by effective endotrivials the effective cone.
+In particular, $\operatorname{Spc}(\mathscr{K}(E))$ is a Dirac scheme.
+Hausmann–Schwede have further considered the twisted cohomology ring for elementary abelian $2$-groups in the context of Mackey functors [HS25].
+The authors determined a minimal generating set and proved that for $p=2$, $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(E)$ is nilpotent-free.
+Interestingly, for elementary abelian $p$-groups, the endotrivials involved in the construction of twisted cohomology are precisely the effective endotrivials we’ve considered, motivating our following “remixed” construction.
+Definition (Definition 5.1).
+Let $G$ be a $p$-group.
 Denote the subset of the canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K}(G))$ arising from nontrivial irreducible real representations of $G$ by $\mathcal{B}(G)$.
-Let $\mathbb{N}^{\mathcal{B}(G)}=\{q\colon\mathcal{B}(G)\to\mathbb{N}\}$ be the monoid of twists, i.e.
-tuples of non-negative integers indexed by the set $\mathcal{B}(G)$.
-The monoid of twists is equivalently nothing more than the submonoid of $\operatorname{Pic}(\mathscr{K}(G))$ generated by $\mathcal{B}(G)$.
+Let $\mathbb{N}^{\mathcal{B}(G)}=\{q\colon\mathcal{B}(G)\to\mathbb{N}\}$ be the monoid of twists, i.e., tuples of non-negative integers indexed by the set $\mathcal{B}(G)$.
 Then the $(\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)})$-graded ring
 $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)=\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G;k):=\bigoplus_{s\in\mathbb{Z}}\bigoplus_{q\in\mathbb{N}^{\mathcal{B}(G)}}\operatorname{Hom}_{\mathscr{K}(G)}\left(k,\bigotimes_{C\in\mathcal{B}(G)}C^{\otimes q(C)}[s]\right)$
-is the (re-)twisted cohomology ring; see Definition 5.1. We note that only non-positive shifts $s\leq 0$ play a nontrivial role.
-We take the first steps towards answering this question for finite $p$-groups by proving injectivity of the comparison map.
-The $p$-group case is perhaps the most fundamental: given a finite group $G$ with $p$-Sylow $S$, one can determine $\operatorname{Spc}(\mathscr{K}(G))$ from $\operatorname{Spc}(\mathscr{K}(S))$ via restriction.
-Theorem A. (Theorem 6.1).
+is the (remixed) twisted cohomology ring.
+We twist only by shifts of effective endotrivials rather than all endotrivials, as otherwise, the graded ring becomes unmanageably large (c.f.
+[BG25, Remark 12.24]).
+Analogously, the group cohomology ring $\operatorname{H}^{\scriptscriptstyle\bullet}(G;k)$ is Noetherian and recovers the Balmer spectrum of the derived category $\operatorname{D}_{b}(kG)$, whereas the Tate cohomology ring $\hat{\operatorname{H}}{}^{\scriptscriptstyle\bullet}(G;k)$ is, for sufficiently interesting $G$, non-Noetherian and contains only nilpotents in negative degrees [BC92].
+Theorem C.
+Let $G$ be a finite $p$-group. The comparison map
+$\operatorname{comp}_{G}\colon\operatorname{Spc}(\mathscr{K}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)),\,\mathscr{P}\mapsto\left\langle f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\text{ homog.}\mid\operatorname{cone}(f)\not\in\mathscr{P}\right\rangle$
+is injective (Theorem 6.2) with dense image (Corollary 5.13).
+If $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is Noetherian, the comparison map is an open immersion.
+(Corollary 6.6)
+The proof of this theorem is highly technical and relies on our forerunner homomorphisms and an induction argument up a subnormal series of $p$-subgroups.
+It is seemingly necessary to twist by all effective endotrivials (or at minimum, a full-rank subset) to obtain these properties; see for instance Remark 4.16.
+In this sense, we believe our construction for $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is the “correct” generalization.
+Since every endotrivial is a line bundle with respect to our open cover $\{U(H)\}$ of Theorem B, for every open $U(H)$, the twisted cohomology ring locally identifies with the usual cohomology ring in the localization $\mathscr{K}(G)|{U(H)}$.
+This allows us to use the classical comparison map of [BAL10a] locally in each open.
+Morally, twisted cohomology globally describes the local data in each open.
+If $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is Noetherian, then we can endow $\operatorname{Spc}(\mathscr{K}(G))$ with Dirac scheme structure, as was the case for elementary abelian $p$-groups.
+Theorem D.
 Let $G$ be a finite $p$-group.
-The comparison map $\operatorname{Spc}(\mathscr{K}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ is injective.
-Generally, injectivity of the comparison map can be a harder condition to prove (if it holds at all, which it frequently may not!) than surjectivity (after a categorical localization), which general tt-geometry usually resolves.
-Circling back to Balmer-Gallauer’s original construction, another feature of $\operatorname{Spc}(\mathscr{K}(E))$ for $E$ elementary abelian is a spectral open cover $\{U(H)\}_{H\leq E}$ of $\operatorname{Spc}(\mathscr{K}(E))$ indexed by subgroups of $E$.
-This cover has two remarkable features:
+If $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is Noetherian, the comparison map is an open immersion and for each subgroup $H\leq G$, restricts to a local homeomorphism
+$\operatorname{comp}_{\mathcal{L}_{G}(H)}\colon U(H)\to\operatorname{Spec}^{h}(\operatorname{End}_{\mathscr{L}_{G}(H)}^{\scriptscriptstyle\bullet}(k)).$
+In this case $(\operatorname{Spc}(\mathscr{K}(G)),\mathcal{O}_{G}^{\scriptscriptstyle\bullet})$ is a Dirac scheme. (Corollary 6.6, Corollary 6.7) Noetherianity holds for the following groups:
   1. (a)
-    Over each open $U(H)$, all endotrivials for $E$ are trivial, i.e. isomorphic to a shift of the tensor unit in the localization $\mathscr{L}(H):=\mathscr{K}(E)|_{U(H)}$;
+    Res-onto $p$-groups (see Definition 7.1). In particular, all Dedekind $p$-groups, groups with every subgroup normal; (Theorem 7.4)
   2. (b)
-    There is a homeomorphism between each $U(H)$ and the homogeneous spectrum of the $\mathbb{Z}$-graded endomorphism ring $\operatorname{End}_{\mathscr{L}(H)}^{\scriptscriptstyle\bullet}(k)$ in the localization $\mathscr{L}(H)$.
-In particular, $\operatorname{Spc}(\mathscr{K}(E))$ is a Dirac scheme.
-In addition, under this open cover, there is an identification $\mathscr{K}(E)|_{U(1)}\cong\operatorname{D}_{b}(kE)$ - in other words, $U(1)$ is nothing more than the “cohomological open,” corresponding to the image under $\operatorname{Spc}$ of the localization $\mathscr{K}(E)\twoheadrightarrow\operatorname{D}_{b}(kE)$.
-We construct a refined open cover with these same properties.
-Theorem B.
-Let $G$ be a finite $p$-group. Then there exists an open cover
-$\{U(H)\}_{H\in s_{p}(G)/G}$ (Construction 4.5) indexed by conjugacy classes of subgroups for which the following holds.
-  1. (a)
-    Each open $U(H)$ contains a unique closed point of $\operatorname{Spc}(\mathscr{K})$, $\mathfrak{m}_{H}$, corresponding to the unique closed point of $V_{{G}/\!\!/{H}}$; (Corollary 4.11)
-  2. (b)
-    The open $U(1)$ is equivalently the cohomological open $V_{G}=\operatorname{Spc}(\operatorname{D}_{b}(kG))$ of $\operatorname{Spc}(\mathscr{K}(G))$.
-In other words, we have an equivalence of categories $\mathscr{K}(G)|_{U(1)}\cong\operatorname{D}_{b}(kG)$; (Theorem 4.17)
-  3. (c)
-    Every endotrivial $C$ is a line bundle, i.e.
-over each open $U(H)$ is isomorphic to a shift of the tensor unit in the localization $\mathscr{K}(G)|_{U(H)}$.
-In particular, we have an isomorphism $C\cong k[h_{C}(H)]$ where $h_{C}$ denotes the unique h-mark function associated to $C$.
-(Theorem 4.13)
-Finally, if $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian, we completely recover the Balmer spectrum.
-Theorem C. (Corollaries 6.4, 6.5).
-If $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian, the comparison map is an open immersion and for each subgroup $H\leq G$, restricts to a local homeomorphism
-$\operatorname{comp}_{\mathcal{L}_{G}(H)}:U(H)\to\operatorname{Spec}^{h}(\operatorname{End}_{\mathscr{L}_{G}(H)}^{\scriptscriptstyle\bullet}(k)).$
-In this case $(\operatorname{Spc}(\mathscr{K}(G)),\mathcal{O}_{G}^{\scriptscriptstyle\bullet})$ is a Dirac scheme.
-To obtain these results, it is seemingly necessary to twist by all effective endotrivials (or at minimum, a full-rank subset); see for instance Remark 4.12.
-In this sense, we believe our construction for $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is the “correct” generalization, and not too big.
-The not-so-secret sauce to the above results is a generalization of Balmer-Gallauer’s maps $a_{N}$ and $b_{N}$ for any effective endotrivial.
-For any effective endotrivial $C$, we construct unique (up to scaling) maps $\iota_{C}^{H}$ for every subgroup $H\leq G$ for which $\Psi^{H}(\iota_{C}^{H})$ is a quasi-isomorphism.
-Proving these maps exist is rather technical and relies heavily on the topological origins of effective endotrivials.
-These maps may not exist for chain complexes not arising from chain complexes of free modules over the orbit category $\Gamma_{G}$ (in particular, non-effective endotrivials), see Remark 3.8.
-Theorem D. (Theorem 4.2).
-Let $G$ be a finite $p$-group.
-For every effective endotrivial complex (see Definition 2.1) and subgroup $H\leq G$, there exists a homomorphism $\iota^{H}_{C}:k\to C[h_{C}(H)]$ such that $\Psi^{H}(\iota^{H}_{C})$ is a quasi-isomorphism.
-Note that there are numerous maps that satisfy the above property (see Remark 4.3); our maps $\iota^{H}_{C}$ satisfy additional technical conditions.
-Finally, we prove some technical results regarding these chain complexes.
-Although they are not strictly necessary for our constructions, they contribute to our understanding of effective endotrivials.
-These propositions may be known to experts, but we reprove them for completeness.
-Theorem E. (Proposition 3.11, Corollary 3.12).
-Let $G$ be a finite group.
-Suppose $C$ is a bounded chain complex of permutation $kG$-modules arising from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules.
-Then $C$ has a contractible direct summand if and only if $C^{?}$ does.
-In particular, if $C$ is an indecomposable effective endotrivial, then $C$ arises from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules.
+    $p$-groups of order up to $p^{3}$. (Corollary 7.9)
+We briefly remark on the connection to [GAL25].
+Gallauer considers twisted cohomology in an abstract context, and shows that when things behave particularly well, these constructions recover the Balmer spectrum and can be used to relate periodicity data between a tt-category and an associated twisted cohomology ring.
+The difficulty however, is showing such things behave well, and that is what we do here.
+In the language of [GAL25], finite generation of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ implies that the maps into the effective endotrivials, which Gallauer terms global sections, generate the topology of $\operatorname{Spc}(\mathscr{K})$.
+This determines a finite ample family of line bundles, which can be taken as the gradings which arise for a set of generators of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$.
+This notion of ampleness almost recovers the notion of ampleness in algebraic geometry.
+If this holds, $\mathscr{K}(G)$ is a divisorial tt-category.
+We characterize the global sections for any effective endotrivial in 5.3.
 Open questions
-There is but one essential ingredient missing to this story, that of noetherianity of the remixed twisted cohomology ring $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$.
-At present, we conjecture that $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian, and in fact, generated by homomorphisms $k\to C[s]$ for $C$ an irreducible endotrivial.
-Sadly, the strategy of [BG25, Lemma 12.12] to prove noetherianity seems to not carry over to the general setting.
-We (sketchily) remark Balmer-Gallauer’s proof can be adapted to any finite abelian $p$-group, but we omit a proof, as it follows near identically as in the elementary abelian case.
-A new possibility we propose is comparing twisted cohomology to the usual cohomology ring.
-We construct a homomorphism from the twisted cohomology ring to a direct product of $p$-local group cohomology rings (see Definition 5.5), and whose kernel is nilpotent.
-Any group cohomology ring is well-known to be noetherian [Ven59, Eve61], therefore if the image of $\prod\hat{\Psi}$ is finite in each $p$-local group cohomology ring, $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian modulo nilpotents.
-See Remark 6.6 for details.
-Beyond this missing link, there remain numerous questions left to answer.
-First, it is quite reasonable to ask if a generalization of twisted cohomology exists for all finite groups.
-Not all endotrivials $C$ arise from a representation spheres (or any reasonable topological space), or can be expressed as (up to homotopy) a chain complex of permutation modules (see e.g.
-[BG23a, Corollary 5.6]).
-As a consequence, many of our maps $\iota^{H}_{C}$ cannot be constructed outside of the $p$-group setting; circumventing this issue is the main challenge.
-To go about constructing twisted cohomology, one may first have to determine $\operatorname{Pic}(\operatorname{K}_{b}(\operatorname{perm}(kG)))$, which at present is unknown (for non-$p$-groups).
-We ask if $\operatorname{Pic}(\operatorname{K}_{b}(\operatorname{perm}(kG)))$ and the subgroup of $\operatorname{Pic}(\mathscr{K}(G))$ generated by representation spheres are in fact equal, and if they are finite index subgroups of $\operatorname{Pic}(\mathscr{K}(G))$.
-Second, while our twisted cohomology ring may be sufficient to capture the subtleties of $\operatorname{Spc}(\mathscr{K}(G))$, it is obviously more cumbersome.
-An explicit presentation in terms of generators and relations is at present seemingly beyond our reach.
-Even explicitly writing down endotrivials for larger groups (e.g.
-$p$-rank at least 3) is a significant challenge and relatively unexplored.
-Finally, although every endotrivial is a line bundle, it is currently unknown and an interesting question whether $\operatorname{Pic}(\mathscr{K}(G)|_{U(H)})\cong\mathbb{Z}$; we expect this to be the case.
-In correspondence, Gallauer gave a proof sketch of this for $p=2$ and $G$ elementary abelian, but not much is known beyond this case.
+The most pertinent question presently is Noetherianity of the remixed twisted cohomology ring $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$. We conjecture this holds.
+Conjecture.
+For any $p$-group $G$, $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is Noetherian. In particular, Theorem D holds in full generality for all $p$-groups.
+Balmer–Gallauer showed this holds for $G$ elementary abelian [BG25, Lemma 12.12], and in Section 7 we adapt their techniques to show it holds in a wider berth of cases.
+However for certain groups of order $p^{3}$, we obtain the first example of a group for which twisted cohomology is not generated by its minimally-graded elements, suggesting that a list of generators of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ in general may grow rapidly as $|G|$ does.
+Our proofs rely very heavily on the topological origin of effective endotrivials, but unfortunately also rely on information about the classification of endotrivials specific to these groups.
+We propose a potential strategy, Remark 7.10, to prove Noetherianity in general, including a lemma which should be sufficient, but at present the full proof appears beyond our grasp.
+However, we have laid out the rest of the groundwork: if one can show that $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is Noetherian, then the analogues to the results of [BG25, Part II] hold for $G$.
+While our twisted cohomology ring may be sufficient to capture the subtleties of $\operatorname{Spc}(\mathscr{K}(G))$, it is obviously cumbersome.
+An explicit presentation in terms of generators and relations, akin to the computations in [BG25, Section 17] are at present beyond our reach for arbitrary groups, but should be more feasible for the groups for which we’ve shown Noetherianity of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$, as one now has an explicit list of generators.
+Beyond twisted cohomology, there remain numerous questions.
+Perhaps the most immediate concerns the forerunner maps: our construction is ad-hoc and depend on a choice of direct sum decomposition, but we expect the existence of a canonical construction or description, perhaps arising from a topological point of view.
+Such a construction would resolve whether certain expected properties for forerunners hold, and elucidate the open covers we construct.
+We remark on this further, see Remark 4.4.
+Our understanding of endotrivials, effective or not, is still incomplete, and this seems to be the limiting factor for proving Noetherianity.
+Though endotrivials are classified, explicitly writing them down for larger groups (e.g.
+non-abelian with $p$-rank at least 3) is a significant challenge.
+One fairly elementary question is as follows: in [MIL25b] it was shown that $\operatorname{Pic}(\mathscr{K}(G))$ is a biset functor in the sense of Bouc [BOU10], and “induction” of endotrivials is induced on the topological level by the norm map.
+How this presents on the level of $\mathscr{K}(G)$ seems difficult to determine, as it is (perhaps surprisingly) not given by tensor induction of chain complexes.
+It is reasonable to ask if a generalization of twisted cohomology exists for all finite groups.
+Not all endotrivials $C$ arise from a representation spheres, so our forerunners cannot always be constructed outside of the $p$-group setting.
+Moreover, if $G/[G,G]$ is not a $p$-group, then not all endotrivials for $kG$ are line bundles.
+Explicitly, any shift of a $k$-dimension 1 $kG$-module $k_{\omega}$, with $\omega\in\operatorname{Hom}(G,k^{\times})$, is endotrivial.
+Related, we do not know a description of the subgroup of $\operatorname{Pic}(\mathscr{K}(G))$ generated by endotrivials induced from representation spheres; we expect this subgroup should be close to full-rank.
+We expect that one can construct an analogous open cover of $\operatorname{Spc}(\mathscr{K}(G))$ from this subgroup, and that some of the techniques of this paper carry over without much difficulty.
+Finally, one may ask questions regarding descent with respect to the open cover $\{U(H)\}$ of $\operatorname{Spc}(\mathscr{K}(G))$ - in particular, there should be a descent spectral sequence associated to $\operatorname{Pic}(\mathscr{K}(G))$, which we will examine in forthcoming work.
+Although we prove every endotrivial is a line bundle, it is currently unknown and an interesting question whether $\operatorname{Pic}(\mathscr{K}(G)|_{U(H)})\cong\mathbb{Z}$ for all $H\leq G$; we expect this to be the case.
+In correspondence, Gallauer gave a proof sketch of this for $p=2$ and $G=V_{4}=C_{2}\times C_{2}$, but not much is known beyond this case.
 Organization
 The paper is organized as follows.
-Section 1 covers some quick preliminaries regarding $p$-permutation modules, modular fixed points/Brauer quotients, and endotrivial complexes.
-Section 2 further reviews Borel-Smith functions.
-Section 3 covers chain complexes of free modules over the orbit category and effective endotrivials, and proves some technical lemmas.
-Section 4 constructs the local quasi-isomorphisms $\iota^{H}_{C}$ and the open cover $\{U(H)\}$ of the Balmer spectrum.
-Section 5 constructs the twisted cohomology ring.
-Section 6 proves injectivity of the comparison map and corollaries, generalizing the results of Balmer-Gallauer in the noetherian setting, and discusses noetherianity.
+Section 1 reviews tensor-triangular geometry, $p$-permutation modules and modular fixed points.
+Section 2 reviews endotrivial complexes and Borel-Smith functions.
+Section 3 covers chain complexes of free modules over the orbit category, effective endotrivials, and proves some technical results.
+Section 4 constructs the forerunners $\iota^{H}_{C}$ and the open cover $\{U(H)\}$ of the Balmer spectrum $\operatorname{Spc}(\mathscr{K}(G))$.
+Section 5 constructs the twisted cohomology ring $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ and proves compatabilities.
+Section 6 proves injectivity of the comparison map and corollaries.
+Finally, Section 7 proves Noetherianity of twisted cohomology for certain cases.
 Notation
-Our notation mostly follows [BG25], as opposed to [Mil25a].
-Any group $G$ is assumed to be finite.
-The Weyl group G//H{G}/\!\!/{H} of $H\leq G$ is the subquotient $N_{G}(H)/H$.
+Our notation mostly follows [BG25], which we warn the reader differs considerably from [MIL24, MIL25a, MIL25b].
+The Weyl group G//H{G}/\!\!/{H} of $H\leq G$ is the subquotient $N_{G}(H)/H$ of $G$.
 Given two subgroups $H,K\leq G$, we write $H=_{G}K$ if $H,K$ are $G$-conjugate, and $H\leq_{G}K$ to denote $H$ is a subgroup of a $G$-conjugate of $K$.
-We denote the set of all $p$-subgroups of $G$ by $s_{p}(G)$, and write $s_{p}(G)/G$ to denote a choice of conjugacy class representatives.
-We write $\operatorname{CF}(G)$ to denote the additive group of superclass functions, and if $G$ is not a $p$-group, we write $\operatorname{CF}(G,p)$ to denote the group of superclass functions valued on $p$-subgroups.
-We denote the trivial $kG$-module, the tensor unit of $\mathscr{K}(G)$, as $k$.
-Given a module $M$, $M[i]$ denotes the chain complex with $M$ in homological degree $i\in\mathbb{Z}$ and zero in all other degrees.
-[1]-[1] also denotes the shift functor on $\mathscr{K}(G)$.
-We denote by $\operatorname{H}^{\scriptscriptstyle\bullet}(G)=\operatorname{H}^{\scriptscriptstyle\bullet}(G;k)$ the $\mathbb{N}$-graded cohomology ring of the tensor unit $k$ in $\operatorname{D}_{b}(kG):=\operatorname{D}_{b}(kG\operatorname{-mod})$ (with both odd and even-degree shifts for $p$ odd).
+We denote the set of all subgroups of $G$ by $s(G)$, the set of all $p$-subgroups of $G$ by $s_{p}(G)$, and write $s_{p}(G)/G$ to denote a choice of conjugacy class representatives.
+We write $\operatorname{CF}(G)$ to denote the additive group of superclass functions, i.e., functions $s(G)\to\mathbb{Z}$ constant on conjugacy classes, and if $G$ is not a $p$-group, we write $\operatorname{CF}(G,p)$ to denote the group of superclass functions valued on $p$-subgroups of $G$.
+Throughout, $\mathscr{K}(G):=\operatorname{K}_{b}(p\operatorname{-perm}(kG))$.
+We denote the tensor unit of $\mathscr{K}(G)$ by $k:=k[0]$ for shorthand (as opposed to $\mathbb{1}$) and the tensor product by $\otimes$ (as opposed to $\otimes_{k}$).
+Given a module $M$, $M[i]\in\mathscr{K}(G)$ denotes the chain complex with $M$ in homological degree $i\in\mathbb{Z}$ and zero in all other degrees.
+We write $M:=M[0]$ if the context is clear.
+The shift functor on $\mathscr{K}(G)$ is also denoted [1]-[1].
+We denote by $\operatorname{H}^{\scriptscriptstyle\bullet}(G)=\operatorname{H}^{\scriptscriptstyle\bullet}(G;k)$ the $\mathbb{N}$-graded cohomology ring over $k$ in $\operatorname{D}_{b}(kG):=\operatorname{D}_{b}(\operatorname{mod}(kG))$ (including both odd and even-degree shifts for $p$ odd).
 Acknowledgments
-I am indebted to Paul Balmer and Martin Gallauer for numerous conversations and unwieldy email threads about their work and tensor-triangular geometry in abstracto.
-I also am grateful to Robert Boltje for his mentorship and suggestions, those of which heavily influenced Section 3 in particular, and the UC Santa Cruz Department of Mathematics for their generous graduate student travel support, which enabled conversations leading to the ideas in this paper to occur.
-1. Preliminaries
-We expediently review some of the essentials of permutation modules from [BG25] and endotrivial complexes [Mil25a].
-For an overview of tensor-triangular geometry, we refer the reader to [Bal05, Bal10b], and for an in-depth “classical” overview of $p$-permutation modules and the Brauer quotient, we refer the reader to [Lin18, Chapter 5] or [Las23].
+The author is indebted to Paul Balmer and Martin Gallauer for numerous conversations and unwieldy email threads about their work, comments on earlier versions of this work, and introducing the author to the world of tensor-triangular geometry.
+He also is grateful to Robert Boltje for his mentorship and suggestions during the course of the author’s Ph.D., and the UC Santa Cruz Department of Mathematics for their generous graduate student travel support, which enabled numerous conversations leading to the ideas in this paper to happen.
+Finally, he thanks Cat & Cloud Coffee of Santa Cruz, CA (on Swift St.) for their warm atmosphere, strong coffee, and impeccable breakfast burritos, countless of which were consumed during the lengthy inception of this project.
+During the course of this paper, the author was partially funded by a UC Santa Cruz dissertation year fellowship and an AMS-Simons travel grant.
+1. Preliminaries: Tensor-triangular geometry and permutation modules
+For the first two sections, we review the essential prerequisites of the paper.
+Section 1 is more on tensor-triangular geometric, and Section 2 concerns endotrivial complexes and their topological connections.
+1.1. Tensor-triangular geometry
+We begin by reviewing the basic concepts of tensor-triangular geometry. For more detailed overviews of tensor-triangular geometry, we refer the reader to [BAL05, BAL10b].
+Definition 1.1.
+Let $\mathscr{K}$ be an essentially small tensor-triangulated category.
+That is, $\mathscr{K}$ is an essentially small triangulated category with a symmetric monoidal structure such that for any $x\in\mathscr{K}$, the functor $x\otimes$- is an exact functor.
+  1. (a)
+    A triangulated subcategory $\mathscr{I}\subseteq\mathscr{K}$ is thick if $\mathscr{I}$ is closed under direct summands, and is a thick $\otimes$-ideal if in addition, for all $x\in\mathscr{I}$ and $y\in\mathscr{K}$, $x\otimes y\in\mathscr{I}$.
+  2. (b)
+    A proper thick $\otimes$-ideal $\mathscr{P}\subset\mathscr{K}$ is prime if for all $x,y\in\mathscr{K}$, $x\otimes y\in\mathscr{P}$ implies $x\in\mathscr{P}$ or $y\in\mathscr{P}$.
+  3. (c)
+    Given a subset of objects $S\subset\mathscr{K}$, we let $\langle S\rangle$ denote the thick $\otimes$-ideal generated by $S$.
+This is equivalently the intersection of all thick $\otimes$-ideals containing $S$.
+The following construction is due to Balmer [BAL05].
+Let $\operatorname{Spc}(\mathscr{K})$ denote the set of prime thick $\otimes$-ideals of $\mathscr{K}$.
+For any $x\in\mathscr{K}$, let $\operatorname{supp}(x)\subseteq\operatorname{Spc}(\mathscr{K})$ denote the subset
+$\operatorname{supp}(x):=\{\mathscr{P}\in\operatorname{Spc}(\mathscr{K})\mid x\not\in\mathscr{P}\},$
+the support of $x\in\mathscr{K}$ Then the collection of all supports induces a closed basis on $\operatorname{Spc}(\mathscr{K})$, and a subset of $Z\subseteq\operatorname{Spc}(\mathscr{K})$ is of the form $Z=\operatorname{supp}(x)$ for some $x\in\mathscr{K}$ if and only if $Z$ is closed with quasi-compact complement [BAL05, Proposition 2.14].
+We call $\operatorname{Spc}(\mathscr{K})$ the Balmer spectrum of $\mathscr{K}$.
+It is well-known that $\operatorname{Spc}(-)$ is contravariantly functorial [BAL05, Proposition 3.6] and the universal support datum for $\mathscr{K}$ in a precise sense [BAL05, Theorem 3.2].
+In particular, any tt-functor (i.e., an exact $\otimes$-functor) induces a map on Balmer spectra.
+Definition 1.2.
+Recall that given any category $\mathscr{C}$, one can form its idempotent completion $\mathscr{C}^{\natural}$ (also called its Karoubi envelope), which roughly speaking, closes the category $\mathscr{C}$ under taking direct summands of objects.
+See for instance [BS01] for triangulated categories.
+Given an open $U\subseteq\operatorname{Spc}(\mathscr{K})$ with closed complement $Z=\operatorname{Spc}(\mathscr{K})\setminus U$, we set
+$\mathscr{K}_{Z}:=\{x\in\mathscr{K}\mid\operatorname{supp}(x)\subseteq Z\},$
+and define the localization
+$\mathscr{K}|_{U}:=\left(\mathscr{K}/\mathscr{K}_{Z}\right)^{\natural}.$
+Then $\mathscr{K}_{Z}$ is a thick $\otimes$-ideal of $\mathscr{K}$, $\mathscr{K}|_{U}$ is again a tensor-triangulated category, and we have a homeomorphism
+$\operatorname{Spc}(\mathscr{K}|_{U})\cong U,$
+see [BF07, Proposition 1.11]. This assignment endows $\operatorname{Spc}(\mathscr{K})$ with the structure of a locally ringed space [BAL05, Remark 6.4].
+More generally, given any thick $\otimes$-ideal $\mathscr{I}\subseteq\mathscr{K}$ with canonical projection $\mathscr{K}\mapsto\mathscr{K}/\mathscr{I}$, the induced map on Balmer spectra $\operatorname{Spc}(\mathscr{K}/\mathscr{I})\to\operatorname{Spc}(\mathscr{K})$ induces a homeomorphism between $\operatorname{Spc}(\mathscr{K}/\mathscr{I})$ and the subspace $\{\mathscr{P}\in\operatorname{Spc}(\mathscr{K})\mid\mathscr{I}\subseteq\mathscr{P}\}$ of $\operatorname{Spc}(\mathscr{K})$ [BAL05, Proposition 3.11].
+Example 1.3.
+Let $G$ be a finite group (scheme), $k$ a field of characteristic $p$ prime, and $\operatorname{D}_{b}(kG)$ denote the bounded derived category of $kG$-modules. We have a homeomorphism
+$\operatorname{Spc}(\operatorname{D}_{b}(kG))\cong\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}(G;k)).$
+This homeomorphism is realized by Balmer’s comparison map [BAL10a]
+$\mathscr{P}\mapsto\langle\{\zeta\in\operatorname{H}^{\scriptscriptstyle\bullet}(G;k)\mid\operatorname{cone}(\zeta)\not\in\mathscr{P}\}\rangle.$
+We discuss an extended version of this comparison map in greater detail in the sequel.
+This theorem is due to Benson–Carlson–Rickard [BCR97] for finite groups (in very different language than we present) and was extended to finite group schemes by Friedlander–Pevtsova [FP07].
+That the comparison map is a homeomorphism follows as a consequence of the classification of thick $\otimes$-ideals in both settings and Balmer’s classification theorem [BAL05, Theorem 4.10].
+Surjectivity follows from Noetherianity of the cohomology ring [EVE61, FS97], and a general tt-geometric fact [BAL10a, Theorem 7.3], but as far as we are aware, there is no known direct proof showing the comparison map is even injective.
+1.2. Permutation modules
+Next, we review the basics of $p$-permutation modules.
+For a purely representation-theoretic overview of $p$-permutation modules and associate constructions, we refer the reader to [LIN18, Chapter 5] or [LAS23].
 Definition 1.4.
@@
     In particular, if $G$ is a finite $p$-group, every $p$-permutation module is a permutation module.
-We set $\mathscr{K}(G):=\operatorname{K}_{b}(p\operatorname{-perm})$, the bounded homotopy category of $p$-permutation modules.
+In this case, the indecomposable permutation modules are precisely those isomorphic to $kX$ for $X$ a transitive $G$-set.
+If $G$ is not a $p$-group, these modules may decompose further.
+We set $\mathscr{K}(G):=\operatorname{K}_{b}(p\operatorname{-perm}(kG))$, the bounded homotopy category of $p$-permutation modules.
+The category $\mathscr{K}(G)$ is a rigid tensor-triangulated category, i.e.
+there is a given by the internal hom $C^{*}:=\operatorname{Hom}_{k}(C,k)$.
+It is an essentially small and is the compact part of a rigidly-compactly-generated (i.e., “big”) tensor-triangulated category $\mathscr{D}(G)$, the derived category of permutation $kG$-modules.
+It
+We note this category can be defined in the context of cohomological Mackey functors, Artin motives, or equivariant stable homotopy theory, and refer the reader to [BG23b, FUH25] for details on these equivalences.
+There are two standard homomorphisms between permutation modules we’ll frequently consider.
 Definition 1.5.
 Let $K\leq H\leq G$ be subgroups of $G$.
-The augmentation homomorphism for $G$-sets $\operatorname{aug}\colon G/K\to G/H$ is the homomorphism induced by the assignment $1K\mapsto 1H$.
-The augmentation homomorphism between the permutation modules $\operatorname{aug}\colon k[G/K]\to k[G/H]$ is the induced map on vector spaces.
-We note the standard example of the augmentation homomorphism is when $H=G$ and $K=1$.
+The augmentation homomorphism between the permutation modules $\operatorname{aug}\colon k[G/K]\to k[G/H]$ is the homomorphism induced by the $G$-set morphism given by $1K\mapsto 1H$.
+The standard example of the augmentation homomorphism is when $H=G$ and $K=1$.
 We have an analogous coaugmentation homomorphism $\operatorname{coaug}\colon k[G/H]\to k[G/K]$ induced from the assignment $1H\mapsto\sum_{g\in[H/K]}gK$.
 This homomorphism of permutation modules is in no way induced from any $G$-set homomorphism.
-Remark 1.3.
-By [BG23a], we have a surjective tt-functor $\mathscr{K}(G)\twoheadrightarrow\operatorname{D}_{b}(kG)$ with kernel $\mathscr{K}_{ac}(G)$, the tt-ideal consisting of all acyclic complexes.
-Set $V_{G}:=\operatorname{Spc}(\operatorname{D}_{b}(kG))$, then [BG25, Proposition 3.22] asserts that this functor induces an open inclusion $V_{G}\hookrightarrow\operatorname{Spc}(\mathscr{K}(G))$.
-We say $V_{G}\subseteq\operatorname{Spc}(\mathscr{K}(G))$ is the cohomological open.
-As noted in the introduction of [BG25], the crux of the matter is its closed complement, $\operatorname{supp}(\mathscr{K}_{ac})$.
-The Brauer quotient is an explicit construction on the level of modules which mimics taking fixed points of $G$-sets.
-Balmer-Gallauer adapt this construction to work on the level of big categories, dubbing them modular fixed points functors.
+Definition 1.6.
+By [BG23a], we have a localization $\mathscr{K}(G)\twoheadrightarrow\mathscr{K}(G)/\mathscr{K}_{ac}(G)\cong\operatorname{D}_{b}(kG)$, where $\mathscr{K}_{ac}(G)$ denotes the thick $\otimes$-ideal of $\mathscr{K}(G)$ consisting of all acyclic complexes.
+We set $V_{G}:=\operatorname{Spc}(\operatorname{D}_{b}(kG))$, then [BG25, Proposition 3.22] asserts that this functor induces an open inclusion $V_{G}\hookrightarrow\operatorname{Spc}(\mathscr{K}(G))$.
+The subset $V_{G}\subseteq\operatorname{Spc}(\mathscr{K}(G))$ is called the cohomological open.
+As noted in the introduction of [BG25], the “interesting” part of $\mathscr{K}(G)$ is its closed complement, $\operatorname{supp}(\mathscr{K}_{ac})$.
+Brauer quotients are explicit constructions on $kG$-modules which mimic taking $H$-fixed points of $G$-sets when $H$ is a $p$-subgroup of $G$.
+Balmer–Gallauer adapt this construction to work on the level of big categories, dubbing them modular fixed points functors.
 Theorem 1.7.
@@
 such that $\Psi^{H}(kX)\cong k[X^{H}]$ for every $G$-set $X$. In particular, this functor preserves compacts and restricts to a tt-functor $\Psi^{H}\colon\mathscr{K}(G)\to\mathscr{K}({G}/\!\!/{H})$.
-Notation 1.5.
+This conservativity theorem is an example of “local-to-global” phenomena in modular representation theory.
+Theorem 1.8.
+[BG25, Theorem 2.11] The family of functors
+$\{\mathscr{D}(G)\xrightarrow{\Psi^{H}}\mathscr{D}({G}/\!\!/{H})\twoheadrightarrow\operatorname{K}(\operatorname{Inj}(k[{G}/\!\!/{H}]))\}_{H\in s_{p}(G)/G}$
+indexed by conjugacy classes of $p$-subgroups $H\leq G$ is conservative, i.e., detects vanishing of objects. This restricts to a conservative family of functors
+$\{\mathscr{K}(G)\xrightarrow{\Psi^{H}}\mathscr{K}({G}/\!\!/{H})\twoheadrightarrow\operatorname{D}_{b}(k[{G}/\!\!/{H}])\}_{H\in s_{p}(G)/G}$
+on the compact parts.
+Notation 1.9.
 Each tt-functor $\Psi^{H}$ induces a continuous map on spectra
@@
 $\hat{\psi}^{H}\colon V_{{G}/\!\!/{H}}\to\operatorname{Spc}(\mathscr{K}(G)).$
-In fact, the Balmer spectrum, as a set, is built entirely from these “$p$-local” cohomological opens.
+These modular fixed point functors and the localization to the bounded derived category are the keys to deducing $\operatorname{Spc}(\mathscr{K}(G))$ as a set.
+The Balmer spectrum, as a set, is built entirely from these $p$-local cohomological opens.
 Theorem 1.10.
-[BG25, Theorem 2.10] Every point of $\operatorname{Spc}(\mathscr{K}(G))$ is the image $\hat{\psi}^{H}(\mathfrak{p})$ of a point $\mathfrak{p}\in V_{{G}/\!\!/{H}}$ for some subgroup $H\leq G$ unique up to $G$-conjugation, i.e.
-we have $\hat{\psi}^{H}(\mathfrak{p})=\hat{\psi}^{H^{\prime}}(\mathfrak{p}^{\prime})$ if and only if there exists a $g\in G$ such that $^{g}H=H^{\prime}$ and $^{g}\mathfrak{p}=\mathfrak{p}^{\prime}$.
-Notation 1.7.
-We denote by $\mathscr{P}_{G}(H,\mathfrak{p})\in\operatorname{Spc}(\mathscr{K})$ the image of the prime $\mathfrak{p}\in V_{{G}/\!\!/{H}}$ under $\hat{\psi}^{H}:V_{{G}/\!\!/{H}}\to\operatorname{Spc}(\mathscr{K}(G))$.
+[BG25, Theorem 2.10] Every point of $\operatorname{Spc}(\mathscr{K}(G))$ is the image $\hat{\psi}^{H}(\mathfrak{p})$ of a point $\mathfrak{p}\in V_{{G}/\!\!/{H}}$ for some subgroup $H\leq G$ unique up to $G$-conjugation, i.e., we have $\hat{\psi}^{H}(\mathfrak{p})=\hat{\psi}^{H^{\prime}}(\mathfrak{p}^{\prime})$ if and only if there exists a $g\in G$ such that $^{g}H=H^{\prime}$ and $^{g}\mathfrak{p}=\mathfrak{p}^{\prime}$.
+In other words,
+$\operatorname{Spc}(\mathscr{K}(G))=\bigsqcup_{H\in s_{p}(G)/G}\operatorname{Spc}(\operatorname{D}_{b}(k[{G}/\!\!/{H}]))\cong\bigsqcup_{H\in s_{p}(G)/G}\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H};k)).$
+Notation 1.11.
+We denote by $\mathscr{P}_{G}(H,\mathfrak{p})\in\operatorname{Spc}(\mathscr{K}(G))$ the image of the prime $\mathfrak{p}\in V_{{G}/\!\!/{H}}$ under $\hat{\psi}^{H}:V_{{G}/\!\!/{H}}\to\operatorname{Spc}(\mathscr{K}(G))$.
+We omit the $G$ subscript if the ambient group is clear.
 By [BG25, Theorem 2.10], every prime is, up to $G$-conjugation, uniquely expressible in this way.
-Theorem 1.8.
-[BG25, Theorem 2.11] The family of functors
-$\{\mathscr{D}(G)\xrightarrow{\Psi^{H}}\mathscr{D}({G}/\!\!/{H})\twoheadrightarrow\operatorname{K}(\operatorname{Inj}(k[{G}/\!\!/{H}]))\}_{H\in s_{p}(G)/G}$
-indexed by conjugacy classes of $p$-subgroups $H\leq G$ is conservative.
-This restricts to a conservative family of functors $\{\mathscr{K}(G)\xrightarrow{\Psi^{H}}\mathscr{K}({G}/\!\!/{H})\twoheadrightarrow\operatorname{D}_{b}(k[{G}/\!\!/{H}])\}_{H\in s_{p}(G)/G}$ on compacts.
-Next, we briefly review the classification of endotrivial complexes.
+The subset $\{\mathscr{P}(H,\mathfrak{p})\mid\mathfrak{p}\in V_{{G}/\!\!/{H}}\}\subseteq\operatorname{Spc}(\mathscr{K}(G))$ is generally not open nor closed when $H\neq 1$, but contains the closed point $\mathscr{P}(H,0)\in\operatorname{Spc}(\mathscr{K}(G))$.
+Explicitly, one can have an inclusion of primes $\mathscr{P}(H,\mathfrak{p})\subseteq\mathscr{P}(H^{\prime},\mathfrak{q})$ where $H>_{G}H^{\prime}$, see [BG25, Corollary 7.13].
+2. Preliminaries: Endotrivial complexes and Borel-Smith functions
+We next review the classification of endotrivial complexes, following [MIL25b].
 Definition 2.1.
-A chain complex $C\in\mathscr{K}(G)$ is endotrivial if it is an invertible object, i.e. $C^{*}\otimes_{k}C\simeq k[0]$.
-Remark 1.10.
-By the Künneth formula for $kG$-modules, for any $p$-subgroup $H\leq G$, $\Psi^{H}(C)$ has nonzero homology in exactly one homological degree, with that homology having $k$-dimension one.
+A chain complex $C\in\mathscr{K}(G)$ is endotrivial if it is an invertible object in $\mathscr{K}(G)$, i.e., we have $C^{*}\otimes C\cong k$.
+The set $\operatorname{Pic}(\mathscr{K}(G))$ of all endotrivials up to isomorphism forms a group under $\otimes$, and is called the Picard group of $\mathscr{K}(G)$.
+By the Künneth formula, for any $p$-subgroup $H\leq G$, $\Psi^{H}(C)$ has nonzero homology in exactly one homological degree, with that homology having $k$-dimension one.
 Let $h_{C}(H)\in\mathbb{Z}$ denote the unique integer for which $\operatorname{H}_{h_{C}(H)}(\Psi^{H}(C))\neq 0$.
+The next proposition essentially reduces the study of endotrivials to considering superclass functions.
 Proposition 2.2.
-[Mil24, Definition 3.5]
+[MIL24, Definition 3.5, Theorem 3.6]The following hold.
   1. (a)
-    The function $h_{C}\colon s_{p}(G)\to\mathbb{Z}$ is a well-defined superclass function, i.e. a function constant on conjugacy classes of subgroups of $G$.
+    The function $h_{C}\colon s_{p}(G)\to\mathbb{Z}$ is a well-defined superclass function.
   2. (b)
-    The assignment $h\colon\operatorname{Pic}(\mathscr{K}(G))\to\operatorname{CF}(G,p)$ given by $C\mapsto h_{C}$ is a well-defined group homomorphism.
+    The assignment $h\colon\operatorname{Pic}(\mathscr{K}(G))\to\operatorname{CF}(G,p)$ given by $C\mapsto h_{C}$ is a well-defined group homomorphism, with kernel isomorphic to $\operatorname{Hom}(G,k^{\times})$.
+Definition 2.3.
 Given an endotrivial $C$, we call $h_{C}$ the h-mark function of $C$, and call the homomorphism $h$ the h-mark homomorphism.
@@
 Since the invertible objects of $\operatorname{D}_{b}(kG)$ consists of shifts of $k$-dimension one $kG$-modules, $h_{C}$ simply tracks the image of $C$ in $\operatorname{D}_{b}(k[{G}/\!\!/{H}])$ where $H$ runs through all $p$-subgroups of $G$.
+By [MIL24, Theorem 3.6], to determine $\operatorname{Pic}(\mathscr{K}(G))$, it effectively suffices to compute $\operatorname{im}(h)$.
 Definition 2.4.
@@
 The collection of Borel-Smith functions forms an additive subgroup $\operatorname{CF}_{b}(G,p)$ of $\operatorname{CF}(G,p)$.
-In fact, if $G$ is a finite $p$-group, then under the identification $\operatorname{CF}(G)\cong B^{*}(G):=\operatorname{Hom}(B(G),\mathbb{Z})$, $\operatorname{CF}_{b}(G)$ forms a rational $p$-biset subfunctor of $\operatorname{CF}(G)$.
+In fact, if $G$ is a finite $p$-group, then under the identification $\operatorname{CF}(G)\cong B^{*}(G):=\operatorname{Hom}(B(G),\mathbb{Z})$, where $B(G)$ denotes the Burnside ring of $G$, $\operatorname{CF}_{b}(G)$ is a rational $p$-biset subfunctor of $\operatorname{CF}(G)$.
 See [BY07, Proposition 3.7] for details.
@@
   1. (a)
-    [Mil25a, Theorem 4.6] Let $G$ be a finite $p$-group. The h-mark homomorphism is an isomorphism onto the subgroup of Borel-Smith functions
+    [MIL25b, Theorem 4.6] Let $G$ be a finite $p$-group. The h-mark homomorphism is an isomorphism onto the subgroup of Borel-Smith functions
     $h\colon\operatorname{Pic}(\mathscr{K}(G))\cong\operatorname{CF}_{b}(G).$
   2. (b)
-    [Mil25a, Corollary 6.4] Let $G$ be a finite group. The h-mark homomorphism has image $\operatorname{CF}_{b}(G,p)$ and induces a split exact sequence
+    [MIL25b, Corollary 6.4] Let $G$ be a finite group. The h-mark homomorphism has image $\operatorname{CF}_{b}(G,p)$ and induces a split exact sequence
     $0\to\operatorname{Hom}(G,k^{\times})\to\operatorname{Pic}(\mathscr{K}(G))\to\operatorname{CF}_{b}(G,p)\to 0.$
+Remark 2.6.
+The classification of endotrivials for a $p$-group can be deduced in multiple ways; one way is by identifying them with representation spheres, which we discuss in the following section.
+This method was considered by Bachmann in his dissertation [BAC16].
+Another way, the method we used in [MIL25b], is by recovering a short exact sequence constructed by Bouc–Yalçin [BY07] relating Borel-Smith functions to the Dade group of a finite group.
+This group parametrizes endopermutation $kG$-modules, modules $M$ for which $M^{*}\otimes M$ is permutation.
+The classification of these modules for $p$-groups was completed by Bouc [BOU06].
+The classification of endotrivials for a non-$p$-group follows from some technical results concerning the homomorphism induced by restriction $\operatorname{Res}^{G}_{H}\colon\operatorname{Pic}(\mathscr{K}(G))\to\operatorname{Pic}(\mathscr{K}(H))$ for subgroups $H$ containing a Sylow $p$-subgroup of $G$ deduced in [MIL25a].
 Example 2.7.
@@
 If $p=2$, one has an endotrivial for $C_{2}$ given by $kC_{2}\to k,$ with $k$ in homological degree 0 and the nonzero differential given by the augmentation homomorphism.
-Otherwise if $p$ is odd, one has an endotrivial for $C_{p}$ by truncating a periodic resolution of the trivial $kC_{p}$-module $k$, $kC_{p}\to kC_{p}\to k$ with $k$ in homological degree 0.
+If $p$ is odd, one has an endotrivial for $C_{p}$ by truncating a periodic resolution of the trivial $kC_{p}$-module $k$, $kC_{p}\to kC_{p}\to k$ with $k$ in homological degree 0.
 For any subgroup $N\triangleleft G$ of index $p$, inflation yields the following complex
 $u_{N}:=k[G/N]\to k[G/N]\to k$
-Here $p$ is assumed odd, if $p=2$ then the highest degree term can be deleted.
-A less easy example of an endotrivial for $p=2$ and $G=D_{2^{n}}$ is as follows.
-Let $H_{1},H_{2}$ be nonconjugate, noncentral subgroups of order 2 (this choice is unique up to conjugacy and reordering).
-Then the following complex is endotrivial, and a small example of a nontrivial faithful, irreducible endotrivial, i.e.
-one arising from a faithful irreducible real representation (see Definition 2.5),
-$kD_{2^{n}}\to k[D_{2^{n}}/H_{1}]\oplus k[D_{2^{n}}/H_{1}]\to k,$
-where all maps between indecomposable permutation modules are augmentation homomorphisms.
-2. Borel-Smith functions and representation spheres
-We begin by considering the topological properties of the main characters of our story, the endotrivials arising from genuine real representations, i.e.
-representations spheres.
-Such endotrivials come from objects of higher structure; they arise from chain complexes of free modules over the orbit category $\Gamma_{G}$, and in particular, their differentials are augmentation homomorphisms which are well-behaved with respect to local-global considerations.
-We discuss this now.
+Here $p$ is assumed odd, if $p=2$ then the degree 2 term is deleted. In the notation of Balmer–Gallauer, we set
+$2^{\prime}:=\begin{cases}1&\text{if }p=2;\\ 2&\text{if }p>2;\end{cases}$
+hence $u_{N}$ has length $2^{\prime}$. We keep the $u_{N}$ notation throughout the paper.
+A more secluded example for $p=2$ and $G=D_{2^{n}}$ (here, we mean the dihedral group with $2^{n}$ elements, rather than the dihedral group acting on the regular polygon with $2^{n}$ edges) is as follows.
+Let $H_{1},H_{2}$ be nonconjugate, non-central subgroups of order 2.
+This choice is unique up to conjugacy and reordering.
+Then the following complex is endotrivial:
+$kD_{2^{n}}\to k[D_{2^{n}}/H_{1}]\oplus k[D_{2^{n}}/H_{2}]\to k,$
+with all maps between indecomposable permutation modules being augmentation homomorphisms, $k$ in homological degree 0.
+Finally, we note a structural fact about endotrivial complexes; the trivial $kG$-module $k$ only appears up to homotopy once.
+Proposition 2.8.
+Let $G$ be a $p$-group and $C\in\operatorname{Pic}(\mathscr{K}(G))$ be an indecomposable endotrivial complex.
+Then $C_{h_{C}(G)}$ contains the trivial $kG$-module $k$ as a direct summand exactly once, and for all $j\neq h_{C}(G),$ $C_{j}$ does not contain $k$ as a direct summand.
+Proof.
+Since $\Psi^{G}(C)\cong k[h_{C}(G)]$, $C_{h_{C}(G)}$ must contain a direct summand isomorphic to $k$, and [MIL24, Lemma 5.8] with $\mathcal{Y}:=s_{p}(G)\setminus\operatorname{Syl}_{p}(G)$ proves the rest.
+∎
+2.1. Borel-Smith functions and representation spheres
+Borel-Smith functions classically arise not in the context of representation theory, but in equivariant topology.
+To close out our preliminaries, we consider their origin: representation spheres, following [TOM87].
 Definition 2.9.
@@
 Similarly, we say that an endotrivial $C$ is effective if its corresponding h-mark function $h_{C}$ is effective.
-This choice of terminology will become clear in the sequel.
-We first recall some important facts about Borel-Smith functions.
-Let $G$ be a nilpotent group and $F$ a field.
-Given a $FG$-module $V$, the dimension function associated to $V$ is the superclass function
+We first recall some important facts about Borel-Smith functions. Let $G$ be a group and $F$ a field. Given a $FG$-module $V$, the dimension function associated to $V$ is the superclass function
 $\dim\colon H\mapsto\dim_{F}V^{H}.$
 If $F$ has characteristic 0, this induces a group homomorphism $R_{F}(G)\to\operatorname{CF}(G)$.
-When $F=\mathbb{R}$, topologists may write $RO(G)=R_{\mathbb{R}}(G)$, as a representation sphere is nothing more than a one-point compactification of a real representation.
+When $F=\mathbb{R}$, it is common to write $\operatorname{RO}(G)=R_{\mathbb{R}}(G)$, where $\operatorname{RO}(G)$ denotes the Grothendieck ring of representation spheres, i.e., 1-point compactifications of real representations (see [TOM87, Page 12]).
+In this case, the dimension function is equivalently computing the dimension of the representation sphere $S^{V^{H}}=(S^{V})^{H}$ of $V^{H}$.
 Theorem 2.10.
-[tD87, Theorem 5.4 and Theorem 5.13, pages 211 and 216] Let $G$ be a nilpotent group.
+[TOM87, Theorem 5.4 and Theorem 5.13, pages 211 and 216] Let $G$ be a nilpotent group.
 The image of $\dim\colon RO(G)\to\operatorname{CF}(G)$ is the group of Borel-Smith functions $\operatorname{CF}_{b}(G)$.
-Moreover, if $f$ is a effective Borel-Smith function, there exists a real representation $V$ for which $\dim(V)=f$.
+Moreover, if $f$ is a effective Borel-Smith function returning non-negative values, there exists a real representation $V$ for which $\dim(V)=f$.
+The following is immediate.
 Corollary 2.11.
@@
 In particular, every endotrivial is isomorphic in $\mathscr{K}(G)$ to the product of an effective endotrivial and the dual of an effective endotrivial.
-Proof.
-This follows from [tD87, Theorem 5.4, page 211] since the dimension function of any representation is effective, and [Mil25a, Theorem 4.6], since the image of the h-mark homomorphism is precisely the group of Borel-Smith functions of $G$.
-∎
-The previous corollary tells us that it suffices to work with effective endotrivials, which are particularly well-behaved.
-In fact, when $G$ is a nilpotent group, the dimension function gives a canonical effective basis of $\operatorname{CF}_{b}(G)$, which arises from the real irreps of $G$.
+Therefore, it suffices to work with effective endotrivials, which we will see are particularly well-behaved.
+In fact, when $G$ is a nilpotent group, the dimension function gives a canonical effective basis of $\operatorname{CF}_{b}(G)$, which arises from the real irreducible representations of $G$.
 Therefore, if $G$ is a finite $p$-group, we have a corresponding canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K})$ consisting of effective endotrivials.
@@
 Since the image of the dimension homomorphism is precisely $\operatorname{CF}_{b}(G)$, it suffices to show that the set $\mathcal{B}:=\{f_{1},\dots,f_{n}\}$ is linearly independent after removing duplicates.
-[tD87, Proposition 5.9, page 213] asserts that $\ker(\dim)$ is generated by elements of the form $V-\psi^{k}(V)$, where $V$ is an irreducible real representation, $\psi^{k}$ is the $k$-th Adams operation, and $k$ is coprime to $|G|$.
-Therefore, the duplicates in $\mathcal{B}$ arise from Adams operation conjugates, and it follows that any set of real irreducible representations which are not Adams operation conjugates will correspond to a linearly independent set of Borel-Smith functions.
+[TOM87, Proposition 5.9, page 213] asserts that $\ker(\dim)$ is generated by elements of the form $V-\psi^{k}(V)$, where $V$ is an irreducible real representation, $\psi^{k}$ is the $k$-th Adams operation, and $k$ is coprime to $|G|$.
+Therefore, the duplicates in $\mathcal{B}$ arise from Adams-conjugates, hence any set of real irreducible representations with no Adams-conjugates will correspond to a linearly independent set of Borel-Smith functions.
 ∎
 Definition 2.13.
-We call the Borel-Smith functions (resp. endotrivials) associated to the real irreps of $G$ the irreducible Borel-Smith functions (resp. endotrivials). These objects are necessarily effective.
+We call the Borel-Smith functions (resp.
+endotrivials) associated to the real irreducible representations of $G$ the irreducible Borel-Smith functions (resp.
+endotrivials).
+These objects are necessarily effective.
+On the other hand, an indecomposable endotrivial means what one would expect.
+It is an easy exercise to verify that in $\mathscr{K}(G)$, for every element $C\in\operatorname{Pic}(\mathscr{K}(G))$, there is a unique indecomposable representative of the isomorphism class of $C$.
 Remark 2.14.
 Given a real representation $V$ of $G$ with corresponding character $\chi$, we have an equality
-$\dim_{\mathbb{R}}V^{H}=\frac{1}{|H|}\sum_{h\in H}\chi(h),$
-a practical method of computing the basis of $\operatorname{CF}_{b}(G)$.
+$\dim_{\mathbb{R}}V^{H}=\frac{1}{|H|}\sum_{h\in H}\chi(h)$
+providing a computational method of computing the basis of $\operatorname{CF}_{b}(G)$.
 Checking character tables of $G$ a finite $p$-group of normal $p$-rank one shows the basis of $\operatorname{Pic}(\mathscr{K}(G))$ obtained in [MIL24, Section 6] coincides with the canonical $\mathbb{Z}$-basis of $\operatorname{CF}_{b}(G)$.
-(The author remarks that he finds this is rather surprising, as the computations of [Mil24] were performed entirely ad-hoc and were done prior to the classification of $\operatorname{Pic}(\mathscr{K}(G))$.)
-Notation 2.7.
+We remark that this is rather surprising, as the computations of [MIL24] were performed entirely ad-hoc, as they occurred well before the classification of $\operatorname{Pic}(\mathscr{K}(G))$.
+Notation 2.15.
 Let $G$ be a finite $p$-group.
@@
 That is, $f_{V}:=\dim V$ and $C_{V}$ is the unique indecomposable endotrivial for which $h_{C_{V}}=\dim V$.
-Per tom Dieck [tD87], Borel-Smith functions correspond to representation spheres, which is simply a manifestation of 2.2.
-If $X$ is a finite-dimensional $G$-complex which is a $\mathbb{F}_{p}$-cohomology sphere, the fixed-point set $X^{H}$ is a $\mathbb{F}_{p}$-cohomology-sphere of dimension $d(H)$.
-The assignment $H\mapsto d(H)+1$ is (unfortunately) also called the dimension function of $X$.
-The set of dimension functions of representation spheres is precisely the effective non-negative Borel-Smith functions.
-Given a real representation $V$, let $S^{V}$ denote the corresponding representation sphere.
+Definition 3.6 will explain how to obtain $C_{V}$ from $S^{V}$ directly.
 3. Chain complexes over the orbit category
-Since effective endotrivials identify with representation spheres, they have additional structural properties, as representation spheres (and more generally, any $G$-CW-complex) produce free chain complexes over the orbit category $\Gamma_{G}$.
-We first review this notion, following [Yal17], then prove some technical results which will be critical for the sequel.
+The connection between representation spheres, endotrivial complexes and Borel-Smith functions lets us discern that effective endotrivials have additional structure: they are induced from representation spheres via Bredon (alternatively called $G$-equivariant) homology.
+$G$-CW-complexes produce chain complexes of free modules over the orbit category $\Gamma_{G}$, which themselves induce particularly nice objects of $\mathscr{K}(G)$, well-behaved with respect to local-global constructions.
+We first review the orbit category, following [YAL17].
 Definition 3.1.
@@
 In particular, $\operatorname{Hom}_{\Gamma_{G}}(H,K)$ is empty unless a conjugate of $H$ is a subgroup of $K$.
+We may write a subgroup $H\leq G$ to denote the object $G/H\in\Gamma_{G}$ for shorthand.
 A $k\Gamma_{G}$-module $M$ is a contravariant functor from the category $\Gamma_{G}$ to the category of $k$-modules.
 By identifying $\operatorname{Aut}_{\Gamma_{G}}(G/H)$ with G//H{G}/\!\!/{H}, $M(H)$ has k[G//H]k[{G}/\!\!/{H}]-module structure.
-The category of finitely generated $k\Gamma_{G}$-modules, denoted $\operatorname{mod}(\Gamma_{G})$ is abelian, so the usual categorical concepts apply.
-Given a $G$-set $X$, we define the $k\Gamma_{G}$-module $k[X^{?}]$ as the module with value at $G/H$ given by $k[X^{H}]$, with the obvious induced maps.
+The category of finitely generated $k\Gamma_{G}$-modules, denoted $\operatorname{mod}(k\Gamma_{G})$ is abelian.
+Given a $G$-set $X$, we define the $k\Gamma_{G}$-module $k[X^{?}]$ as the module with value at $G/H$ given by the subspace $k[X^{H}]$, with the obvious induced maps.
 A module over $\Gamma_{G}$ is free if it is isomorphic to a direct sum of modules of the form $k[(G/K)^{?}]$.
-Let $\operatorname{proj}(\Gamma_{G})$ denote the full subcategory of $\operatorname{mod}(\Gamma_{G})$ consisting of free $\Gamma_{G}$-modules (by the Yoneda lemma, every projective $\Gamma_{G}$-module is free, see [Yal17, Definition 3.1]).
+Let $\operatorname{proj}(k\Gamma_{G})$ denote the full subcategory of $\operatorname{mod}(k\Gamma_{G})$ consisting of free $\Gamma_{G}$-modules (by the Yoneda lemma, every projective $\Gamma_{G}$-module is free, see [YAL17, Definition 3.1]).
 The next propositions follow immediately from definition of $k\Gamma_{G}$-modules.
@@
 Any $k\Gamma_{G}$-homomorphism $f\colon k[X^{?}]\to k[Y^{?}]$ satisfies the following property: given any $x\in X$ and subgroup $H\leq G$, if $x\in X^{H}$, then $f(x)\in k[Y^{H}]$.
-In particular, $\operatorname{Hom}_{\operatorname{mod}(\Gamma_{G})}(k[X^{?}],k[Y^{?}])$ has $k$-basis induced from the set $\operatorname{Hom}_{G-\operatorname{set}}(X,Y)$.
+In particular, $\operatorname{Hom}_{\operatorname{mod}(k\Gamma_{G})}(k[X^{?}],k[Y^{?}])$ has $k$-basis induced from $\operatorname{Hom}_{G-\operatorname{set}}(X,Y)$.
 Proposition 3.3.
@@
 Remark 3.4.
-As a result, each free $k\Gamma_{G}$-module has a stratification by fixed points.
-That is, given a free $k\Gamma_{G}$-module $kX$, we may write $kX=kG^{\oplus a_{1}}\oplus k[G/H_{2}]^{\oplus a_{2}}\cdots\oplus k[G/G]^{a_{n}}$ for some $a_{1},\dots,a_{n}\in\mathbb{N}$, where each direct summand $k[G/H_{i}]^{\oplus a_{i}}$ is a uniquely determined submodule of $kX$.
+As a result, each free $k\Gamma_{G}$-module has a filtration by fixed points.
 Therefore, we may fix a “canonical permutation basis” in accordance with this stratification.
 We will make considerable use of this fact in the sequel.
-Given a $G$-set $X$ and two subgroups $K\leq H$ of $G$, we have a canonical inclusion homomorphism (of $k[N_{G}(H)\cap N_{G}(K)]$-modules) $i\colon k[X^{H}]\to k[X^{K}]$ associated to the $G$-set homomorphism $G/K\to G/H,1K\mapsto 1H$.
+Given a $G$-set $X$ and two subgroups $K\leq H$ of $G$, we have a canonical inclusion homomorphism of $k[N_{G}(H)\cap N_{G}(K)]$-modules $i\colon k[X^{H}]\to k[X^{K}]$ associated to the $G$-set homomorphism $G/K\to G/H,1K\mapsto 1H$.
 This comes associated with a (not necessarily unique) projection map $p\colon k[X^{K}]\to k[X^{H}]$ satisfying $p\circ i=\operatorname{id}$ and $p(m)=0$ if $m\in\Psi^{H}(k[X^{K}])=0$.
-After choosing a canonical permutation basis, we obtain a corresponding projection map associated to the basis.
-For any subgroup $H\leq G$, we have an obvious functor $(H)\colon\operatorname{proj}(\Gamma_{G})\to\operatorname{perm}(k[{G}/\!\!/{H}])$ from the assignment $k[X^{?}]\to k[X^{H}]$.
+After choosing a canonical permutation basis, we obtain an corresponding projection map associated to the basis.
+For any subgroup $H\leq G$, we have an obvious functor $(H)\colon\operatorname{proj}(k\Gamma_{G})\to\operatorname{perm}(k[{G}/\!\!/{H}])$ from the assignment $k[X^{?}]\to k[X^{H}]$.
 Modular fixed points behave as one would hope.
@@
 Let $P$ be a $p$-subgroup of $G$. Then the following diagram commutes up to natural isomorphism.
-$\operatorname{proj}(\Gamma_{G})$$\operatorname{perm}(kG)$$\operatorname{perm}(k[{G}/\!\!/{P}])$$\scriptstyle{-(1)}$$\scriptstyle{-(P)}$$\scriptstyle{\Psi^{P}}$
+$\operatorname{proj}(k\Gamma_{G})$$\operatorname{perm}(kG)$$\operatorname{perm}(k[{G}/\!\!/{P}])$$\scriptstyle{-(1)}$$\scriptstyle{-(P)}$$\scriptstyle{\Psi^{P}}$
 Proof.
 This follows from the natural isomorphism $k[X^{P}]\cong\Psi^{P}(kX)$. ∎
+We now explain how to obtain an endotrivial complex from a representation sphere, via a construction topologists will be familiar with.
 Definition 3.6.
 Let $X$ be a $G$-CW-complex.
-The reduced chain complex $X$ over the orbit category is the functor $\tilde{C}_{\ast}(X^{?};k)$ from the orbit category $\Gamma_{G}$ to the category of chain complexes of $k$-modules.
-This gives rise to a chain complex of free (by [Yal17, Lemma 3.2]) $k\Gamma_{G}$-modules
-$\tilde{C}_{\ast}(X^{?};k):=\cdots\to C_{i}(X^{?};k)\xrightarrow{d_{i}}C_{i-1}(X^{?};k)\to\cdots\to C_{0}(X^{?};k)\xrightarrow{\epsilon}\underline{k}\to 0,$
-where $\underline{k}$ denotes the constant functor with values $k(H)=k$ for all subgroups $H\leq G$, and $\epsilon$ denotes the augmentation homomorphism (this is essentially the construction of reduced Bredon homology, see [Ill73]).
+The reduced cellular chain complex of $X$ over the orbit category is the functor $\tilde{C}(X^{?}):=\tilde{C}(X^{?};k)$ from the orbit category $\Gamma_{G}$ to the category of chain complexes of $k$-modules.
+This gives rise to a chain complex of free $k\Gamma_{G}$-modules
+$\tilde{C}(X^{?}):=\cdots\to C_{i}(X^{?})\xrightarrow{d_{i}}C_{i-1}(X^{?})\to\cdots\to C_{0}(X^{?})\xrightarrow{\epsilon}\underline{k}\to 0,$
+where $\underline{k}$ denotes the constant functor with values $k(H)=k$ for all subgroups $H\leq G$, and $\epsilon$ denotes the augmentation homomorphism.
+By convention $\underline{k}$ is in homological degree -1.
+This is essentially the construction of reduced Bredon homology with coefficients in $k$, see [ILL73].
 We denote chain complexes of free $k\Gamma_{G}$-modules by $C^{?}$, and their evaluations at the subgroup $H\leq G$ by $C^{H}$ to inspire an aura of fixed points.
-If $G$ is a finite $p$-group and $S$ is a representation sphere of $G$, it may be realized as a $G$-CW-complex, hence producing a chain complex of free $k\Gamma_{G}$-modules.
-Therefore for any subgroup $H\leq G$, $\tilde{C}_{\ast}(S^{H};k)$ is an effective endotrivial complex of permutation k[G//H]k[{G}/\!\!/{H}]-modules.
-Conversely, after shifting, every effective endotrivial complex $C$ of $kG$-modules is homotopy equivalent to $\tilde{C}_{\ast}(S^{1};k)$ for some representation sphere $S$ of $G$ with dimension homomorphism $\dim(S)=h_{C}$ (see [tD87, Page 217, Theorem 5.16]), and hence up to homotopy arises from a chain complex of free $k\Gamma_{G}$-modules.
-This proposition is critical for many of the constructions to follow: it allows us to perform lifts of $p$-local homomorphisms $k\to\Psi^{H}(C)$.
+Given any chain complex of free $k\Gamma_{G}$-modules, evaluation at $1$ produces a chain complex of $p$-permutation kGkG-modules. This assignment induces a tt-functor
+$(-)^{1}\colon\operatorname{K}_{b}(\operatorname{proj}(k\Gamma_{G}))\to\mathscr{K}(G)$, and the analogous commutative diagram to 3.5 commutes.
+Let $G$ be a $p$-group.
+If $S^{V}$ is a representation sphere of $G$, it may be realized as a $G$-CW-complex, hence producing a chain complex of free $k\Gamma_{G}$-modules.
+Therefore for any subgroup $H\leq G$, $\tilde{C}(S^{H})$ is the corresponding effective endotrivial complex of permutation k[G//H]k[{G}/\!\!/{H}]-modules.
+It is clear there is an equality
+$h_{C}(H)=\operatorname{H}_{i}(\tilde{C}(S^{H}))=\dim_{\mathbb{R}}V^{H}$
+for all $H\leq G$.
+Conversely, by identifying an endotrivial $C$ with its h-mark homomorphism $h_{C}$, Theorem 2.10 asserts $C$ is up to shift homotopy equivalent to $\tilde{C}(S^{1})$ for some representation sphere $S$ of $G$ with dimension homomorphism $\dim(S)=h_{C}$ by 3.5.
+Therefore, up to homotopy, every effective endotrivial arises from a chain complex of free $k\Gamma_{G}$-modules.
+We will show that one can remove “up to homotopy” in the sequel.
 Proposition 3.7.
@@
 Suppose $k[X^{H}]\neq 0$, and let $M^{\prime}$ denote the unique minimal nonzero submodule of $k[X^{H}]$.
-If $M^{\prime}\subseteq\ker(f^{H})$, then there exists a submodule $M^{?}\subseteq k[X^{?}]$, minimal with respect to the property that $M^{L}\neq 0$ when $k[X^{L}]\neq 0$, such that $M^{H}=M^{\prime}$ and for all subgroups $K\leq_{G}H$, $M^{K}\subseteq\ker(f^{K})$.
+If $M^{\prime}\subseteq\ker(f^{H})$, then there exists a unique submodule $M^{?}\subseteq k[X^{?}]$, minimal with respect to the property that $M^{L}\neq 0$ when $k[X^{L}]\neq 0$, such that $M^{H}=M^{\prime}$ and for all subgroups $K\leq_{G}H$, $M^{K}\subseteq\ker(f^{K})$.
 Proof.
-We define $M^{?}\subseteq k[X^{?}]$ as follows: for $L\leq G$, $M^{L}\subset k[X^{L}]$ is the unique minimal submodule if $X$ has any $K$-fixed points, and the zero module if not.
-It is straightforward that, by construction, $M^{?}$ is minimal with respect to the property that $M^{L}\neq 0$ when $k[X^{L}]\neq 0$.
-Now by minimality of $M^{?}$, it suffices to show $f^{K}$ is not an injective k[G//K]k[{G}/\!\!/{K}]-module homomorphism for all $K\leq_{G}H$, but this follows directly from Proposition 3.3.
+We define $M^{?}\subseteq k[X^{?}]$ as follows: for $L\leq G$, $M^{L}\subset k[X^{L}]$ is the unique minimal submodule of $k$-dimension 1 if $X$ has any $L$-fixed points, and 0 if not.
+By construction, $M^{?}$ is minimal with respect to the property that $M^{L}\neq 0$ when $k[X^{L}]\neq 0$.
+Now by minimality of $M^{?}$, it suffices to show $f^{K}$ is not an injective k[G//K]k[{G}/\!\!/{K}]-module homomorphism for all $K\leq_{G}H$, and this follows from 3.3.
 ∎
 Remark 3.8.
-Proposition 3.7 asserts that we can construct “local” inclusions $k\to C[s]$ for chain complexes arising from chain complexes over the orbit category.
+3.7 asserts that we can lift “$p$-local” morphisms $\hat{f}\colon k\to\Psi^{H}(C[s])$ to “global” morphisms $f\colon k\to C[s]$ satisfying $\Psi^{H}(f)=\hat{f}$ for chain complexes arising from chain complexes over the orbit category.
 This does not hold for arbitrary objects of $\mathscr{K}(G)$.
 For instance, the two-term complex $C=k\xrightarrow{\operatorname{coaug}}kG$ (with $k$ in homological degree 0) has no nonzero global homomorphism $k\to C$, but there exist nonzero local homomorphisms $k\to\Psi^{H}(C)$ (in fact isomorphisms) for every nontrivial subgroup $H\leq G$, since $\Psi^{H}(kG)=0$.
+These constructions make sense in $\mathscr{K}(G)$ rather than $\operatorname{K}_{b}(\operatorname{mod}(k\Gamma_{G}))$, as $\operatorname{Hom}_{\operatorname{mod}(k\Gamma_{G})}(\underline{k},k[X^{?}])=0$ for any nontrivial transitive $G$-set $X$.
 3.1. Removing contractible summands
 We prove a couple technical results for finite $p$-groups that show we can remove contractible summands from chain complexes of free $k\Gamma_{G}$-modules if and only if we can for the corresponding chain complexes of permutation $kG$-modules (in particular, for effective endotrivials).
-Though most of these results are not strictly necessary for future sections, they provide elucidation and may be of independent interest.
 Remark 3.9.
-A chain complex of free $k\Gamma_{G}$ modules has in each homological degree a canonical permutation basis, and each differential respects the stabilizers grow condition.
+A chain complex of free $k\Gamma_{G}$ modules has in each homological degree a canonical permutation basis induced from the natural inclusions $X^{H}\hookrightarrow X^{K}$ for $K\leq_{G}H\leq G$, and each differential respects the stabilizers grow condition.
 Moreover, it is easy to see that given a chain complex $C$ of permutation $kG$-modules, if one can choose a canonical permutation basis of $G$-sets in each degree such that the stabilizers grow condition holds with respect to this basis, then $C$ may be realized as the image of a chain complex of free $k\Gamma_{G}$-modules evaluated at the subgroup $1\leq G$.
-In this sense, the chain complexes of permutation $kG$-modules satisfying the stabilizers grow condition on differentials are exactly those which arise from chain complexes of free $k\Gamma_{G}$-modules.
-If $C$ is a chain complex of permutation modules, we say $C$ arises from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules if there exists a chain complex of free $k\Gamma_{G}$-modules $C^{?}$ such that $C^{1}\cong C$ as chain complexes.
+In this sense, the chain complexes of permutation $kG$-modules satisfying the stabilizers grow condition on differentials with respect to some chosen permutation basis are exactly those which arise from chain complexes of free $k\Gamma_{G}$-modules.
+If $C$ is a chain complex of permutation modules, we say $C$ arises from a chain complex of free $k\Gamma_{G}$-modules if $C$ is in the essential image of the “evaluation at G/1G/1” $t$-functor $(-)^{1}\colon\operatorname{K}_{b}(\operatorname{proj}(k\Gamma_{G}))\to\mathscr{K}(G)$.
 Proposition 3.10.
@@
 Set $b$ equal to the index $b_{j}$ for which $Y_{b_{j}}$ has minimal stabilizers (this is well-defined since $Y_{b_{j}}$ is transitive).
-It follows from Corollary 3.3 that there exists an index $a\in\{1,\dots,m\}$ such that the composition $p_{b}\circ d_{i}\circ i_{a}$ is an isomorphism, where $p_{b}$ denotes projection onto $kY_{b}$ and $i_{a}$ denotes inclusion into $kX_{a}$.
+It follows from 3.3 that there exists an index $a\in\{1,\dots,m\}$ such that the composition $p_{b}\circ d_{i}\circ i_{a}$ is an isomorphism, where $p_{b}$ denotes projection onto $kY_{b}$ and $i_{a}$ denotes inclusion into $kX_{a}$.
 In particular, $X_{a}$ and $Y_{b}$ are isomorphic $G$-sets.
@@
 Choose any generator $t\in T$ of $T$, then we may write $t=p_{a}(t)+t^{\prime}$, with $t^{\prime}\in T_{1}\oplus\cdots\oplus\widehat{T}_{a}\oplus\cdots\oplus T_{m}$.
-The $kG$-modules $\langle p_{a}(t)\rangle$ and $\langle t^{\prime}\rangle$ are both isomorphic to $k$, i.e.
-are $G$-stable.
+The $kG$-modules $\langle p_{a}(t)\rangle$ and $\langle t^{\prime}\rangle$ are both isomorphic to $k$, i.e., are $G$-stable.
 We replace the $C_{i}$ basis elements $\{x_{1},\dots,x_{e}\}=X_{a}\subset kX_{a}\subseteq C_{i}$ with $\{x_{1}+t^{\prime},\dots,x_{e}+t^{\prime}\}=X_{a}^{\prime}\subset C_{i}$.
@@
 Therefore $p_{b}\circ d_{i}(y_{j}+t^{\prime})$ has the same $G$-stabilizer as both $y_{j}$ and $y_{j}+t^{\prime}$.
-Hence $Y_{b}^{\prime}$ is isomorphic as a $G$-set to $Y_{b}$ and $X_{a}$, and it follows that $C_{i-1}=C^{\prime}_{i-1}\oplus kY^{\prime}_{b}$.
+Hence $Y_{b}^{\prime}$ is isomorphic as a $G$-set to $Y_{b}$ and $X_{a}$, and $C_{i-1}=C^{\prime}_{i-1}\oplus kY^{\prime}_{b}$.
 Moreover, $d_{i}$ restricts to an isomorphism $kX_{a}^{\prime}\xrightarrow{\cong}kY_{b}^{\prime}$, and the result follows.
@@
 The next proposition demonstrates that if a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules satisfies that $C^{1}$ has a contractible summand (as a chain complex of $kG$-modules), then $C^{?}$ itself has a contractible summand (as a chain complex of $k\Gamma_{G}$-modules).
+This fact will be rather useful for studying effective endotrivials.
+It is unclear if a non-technical proof of this fact exists, but we suspect that for chain complexes arising from $G$-CW-complexes, it may be so.
 Proposition 3.11.
 Let $G$ be a finite group and suppose $C$ is a bounded chain complex of permutation $kG$-modules arising from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules.
-If $C$ contains a contractible direct summand $K$, then there exists a direct sum decomposition $C=K\oplus D$ such that $D$ also arises from a chain complex $D^{?}$ of free $k\Gamma_{G}$-modules.
+If $C$ contains a contractible direct summand $K$ of permutation modules, then there exists a direct sum decomposition $C=K\oplus D$ such that $D$ also arises from a chain complex $D^{?}$ of free $k\Gamma_{G}$-modules.
 In particular, $C$ has a contractible direct summand of permutation modules if and only if $C^{?}$ has a contractible direct summand.
@@
 To prove this, it suffices to find a decomposition of $C$ into direct summands $K,D$ as stated and show that each homological degree of $D$ has a canonical permutation basis satisfying the stabilizers grow condition.
+If $C$ is indecomposable, there is nothing to show, so assume $C$ decomposes.
 First, since $C$ has a contractible direct summand, there exists an integer $i\in\mathbb{Z}$ and $k$-dimension one submodule $T\subseteq C_{i}$ such that $T\not\in\ker(d_{i})$.
@@
 After projecting $T$ onto each $kX_{j}$, it follows that there exists an index $l$ for which $d_{i}|_{X_{l}}$ is injective.
-Choose the index $l$ such that the stabilizer of $X_{l}$ is minimal with respect to subgroup inclusion.
+Choose an index $l$ such that the stabilizer of $X_{l}$ is minimal with respect to group order.
 We modify the canonical bases of $C_{i}$ and $C_{i-1}$ as follows.
-First, it follows from Corollary 3.3 that there exists a (non-unique) index $h\in\{1,\dots,n\}$ such that $p_{h}\circ d_{i}\circ i_{l}$ is an isomorphism, where $i_{l}$ denotes inclusion into $kX_{l}$ and $p_{h}$ denotes projection onto $kY_{h}$.
+First, 3.3 implies there exists a (non-unique) index $h\in\{1,\dots,n\}$ such that $p_{h}\circ d_{i}\circ i_{l}$ is an isomorphism, where $i_{l}$ denotes inclusion into $kX_{l}$ and $p_{h}$ denotes projection onto $kY_{h}$.
 We replace the basis elements of $Y_{h}$ with the basis elements $Y_{h}^{\prime}=\{d_{i}(x)\mid x\in X_{l}\}$.
@@
 Set $C_{i}^{\prime}:=kX_{1}\oplus\cdots\oplus\widehat{kX}_{l}\oplus\cdots\oplus kX_{m}$, so we have $C_{i}=C_{i}^{\prime}\oplus K_{i}$.
-For each canonical permutation basis element $x$ of $C_{i}$ (i.e.
-some $x\in X_{j}$ for $j\neq l$) replace $x$ with $x^{\prime}:=x-((d_{i}|_{kX_{l}})^{-1}\circ p_{h}^{\prime}\circ d_{i})(x)$, where $p_{h}^{\prime}$ denotes projection onto $kY_{h}^{\prime}$ and $(d_{i}|_{kX_{l}})^{-1}$ denotes the inverse of the isomorphic projection of $kX_{l}$ onto $kY_{h}^{\prime}$.
+For each canonical permutation basis element $x$ of $C_{i}$ (i.e., some $x\in X_{j}$ for $j\neq l$) replace $x$ with $x^{\prime}:=x-((d_{i}|_{kX_{l}})^{-1}\circ p_{h}^{\prime}\circ d_{i})(x)$, where $p_{h}^{\prime}$ denotes projection onto $kY_{h}^{\prime}$ and $(d_{i}|_{kX_{l}})^{-1}$ denotes the inverse of the isomorphic projection of $kX_{l}$ onto $kY_{h}^{\prime}$.
 A straightforward computation shows $d_{i}(x^{\prime})\in D_{i-1}$.
@@
 Finally, for all $j\neq i,i-1$, set $D_{j}=C_{j}$.
-It follows that we have a direct sum decomposition $C=K\oplus D$.
+We have a direct sum decomposition $C=K\oplus D$.
 Finally, observe that for each $x^{\prime}$, under the image of $d_{i}$, the canonical permutation basis representation of $x^{\prime}$ in $D_{i-1}$ is identical to the canonical permutation basis representation of $x$ in $D_{i-1}$, therefore the stabilizers grow property holds for the differential $d_{i}$.
@@
 Thus $d_{i+1}$ also satisfies the stabilizers grow condition, as desired.
+Corollary 3.12.
+Let $G$ be a finite group. The tt-functor
+$(-)^{1}:\operatorname{K}_{b}(\operatorname{proj}(k\Gamma_{G}))\to\mathscr{K}(G)$
+induced by evaluation at the trivial subgroup $1$ of $G$ is conservative, i.e., detects vanishing of objects.
+Proof.
+3.11 implies the functor $(-)^{1}:\operatorname{K}_{b}(\operatorname{proj}(k\Gamma_{G}))\to\operatorname{K}_{b}(\operatorname{perm}(kG))$ is conservative, and the canonical inclusion $\operatorname{K}_{b}(\operatorname{perm}(kG))\to\operatorname{K}_{b}(\operatorname{perm}(kG))^{\natural}\cong\mathscr{K}(G)$ is conservative as well.
+∎
 In particular, if $G$ is a finite $p$-group, then every indecomposable effective endotrivial arises from a free chain complex of $k\Gamma_{G}$-modules.
-Corollary 3.12.
-Let $G$ be a finite $p$-group and let $C$ be an indecomposable effective endotrivial of $kG$-modules. Then $C$ arises from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules.
+Theorem 3.13.
+Let $G$ be a finite $p$-group and let $C$ be an indecomposable effective endotrivial of $kG$-modules.
+Then $C$ arises from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules.
+In particular, any effective endotrivial arises from a chain complex of free $k\Gamma_{G}$-modules.
 Proof.
@@
 The representation sphere $S^{V}$ is a $G$-CW-complex, therefore produces a (reduced) chain complex $C^{?}$ for which $C^{1}$ is an endotrivial with the same h-marks as $C$.
-By Corollary 3.11, we can remove all contractible summands from $C^{?}$ until $C^{1}$ is indecomposable, and since two indecomposable endotrivials with the same h-marks are isomorphic, the result follows.
+By 3.11, we can remove all contractible summands from $C^{?}$ until $C^{1}$ is indecomposable, and since two indecomposable endotrivials with the same h-marks are isomorphic, the result follows.
+The last statement follows easily by considering a direct sum decomposition of an effective endotrivial into its indecomposable endotrivial summand and contractible summand.
 ∎
-Moreover, we can completely identify the homomorphisms $k\to C$ for an effective endotrivial $C$.
+We also prove a tightening of 2.8 for effective endotrivials. The examples in Example 2.7 demonstrate this, but it is not hard to find counterexamples of the proposition for $C$ non-effective.
+Corollary 3.14.
+Let $G$ be a $p$-group and $C\in\operatorname{Pic}(\mathscr{K}(G))$ be an indecomposable effective endotrivial complex. Then $C_{h_{C}(G)}=k$.
+Proof.
+Without loss of generality we assume $h_{C}(G)=0$.
+By 2.8 we have $C_{0}$ has a direct summand isomorphic to $k$ and for all $i\neq 0$, $k$ is not a direct summand of $C_{i}$.
+Furthermore, since $C$ is effective, by a standard $p$-permutation argument (see [LIN18, Proposition 5.8.11]), $C_{i}=0$ for negative $i$.
+If $h_{C}\equiv 0$ then $C=k[0]$ so there is nothing to show.
+Otherwise, there exists a subgroup $H$ such that $h_{C}(K)>0$, and in particular $h_{C}(1)>0$.
+Now assume for contradiction that $C_{0}$ has another indecomposable summand $M\cong k[G/K]$ with $K<G$.
+Since $h_{C}(1)>0$, $p_{M}\circ d_{1}$ is surjective, and because $C$ arises from a complex of free $k\Gamma_{G}$-modules, the fiber of $p_{M}\circ d_{1}$ in $C_{1}$ consists of indecomposable modules $k[G/K^{\prime}]$ with $K^{\prime}\leq K$.
+Here $p_{M}$ denotes projection onto $p_{M}$.
+If $h_{C}(K)=0$, then monotonicity implies $h_{\operatorname{Res}^{G}_{K}}(C)\equiv 0$ so $\Psi^{K}(C)\cong k$.
+Therefore, there must be a summand $M^{\prime}$ of $C_{1}$ also isomorphic to $k[G/K]$ in the fiber of $p_{M}\circ d_{1}$ to “cancel” $\Psi^{K}(M)$ via homotopy equivalence.
+But by [LIN18, Proposition 5.10.7] the composition $p_{M}\circ d_{1}\circ i_{M^{\prime}}$ is an isomorphism if and only if $\Psi^{K}(\iota_{M^{\prime}}\circ d_{1}\circ p_{M})$ is.
+Here $i_{M^{\prime}}$ denotes the inclusion from $M^{\prime}$ into $C_{1}$.
+Therefore a contractible summand splits off (c.f.
+[BM26, Lemma 9.2]), contradicting indecomposability of $C$.
+Therefore, $h_{C}(K)>0$ must hold.
+So $\Psi^{K}(d_{1})$ is surjective, and there must also be a summand $M^{\prime}$ of $C_{1}$ in the fiber of $p_{M}\circ d_{1}$ such that $\Psi^{K}(p_{M}\circ d_{1}\circ i_{M^{\prime}})$ is surjective.
+Now a similar argument as before yields a contractible summand of $C$, a contradiction.
+We’ve exhausted all cases, thus $C_{0}=k$.
+This final corollary will be critical in the sequel: it shows that nonzero homomorphisms $k\to C[s]$ for $C$ indecomposable and arising from a complex of free $k\Gamma_{G}$ modules, can be drawn without concern.
+If $C$ is endotrivial, [GAL25] refers to such morphisms as global sections.
 Corollary 3.15.
@@
 Proof.
-The existence of such a homomorphism $f$ is equivalent to the inclusion $T\subseteq\ker(d_{i}),$ and this inclusion holds since if not, Proposition 3.10 implies the existence of a contractible summand of $C^{1}$, hence a contractible summand of $C^{?}$, which cannot occur.
+The existence of such a homomorphism $f$ is equivalent to the inclusion $T\subseteq\ker(d_{i}),$ and this inclusion holds since if not, 3.10 implies the existence of a contractible summand of $C^{1}$, hence a contractible summand of $C^{?}$, which cannot occur.
 The last statement is straightforward.
@@
 4. An open cover of the Balmer spectrum via endotrivials
-For the rest of the paper, we assume $G$ is a finite $p$-group.
-4.1. $p$-local quasi-isomorphisms
+Notation: For the rest of the paper, we assume $G$ is a finite $p$-group.
+4.1. Constructing forerunners
+We construct our “$p$-local quasi-isomorphisms,” the forerunners, for effective endotrivials. This first proposition establishes the uniqueness of their targets.
 Proposition 4.1.
@@
 First, $V$ is equivalently a faithful $\mathbb{R}[G/N]$-module, so it suffices to assume $N=1$ after replacing $G$ with $G/N$, and it suffices to show $C_{n}$ is indecomposable and projective.
-Since $C$ is effective, it follows by an inductive argument that for all subgroups $H\leq G$, $\Psi^{H}(C)$ is isomorphic to an indecomposable complex in $\mathscr{K}({G}/\!\!/{H})$ whose highest nonzero homological degree is $h_{C}(H)$ (see [Lin18, Proposition 5.8.11]).
+Since $C$ is effective, it follows by an inductive argument (using [LIN18, Proposition 5.8.11]) that for all subgroups $H\leq G$, $\Psi^{H}(C)$ is isomorphic to an indecomposable complex in $\mathscr{K}({G}/\!\!/{H})$ whose highest nonzero homological degree is $h_{C}(H)$ .
 In particular, $n=h_{C}(1)$ is the maximal nonzero degree of $C$, since $C$ is indecomposable.
@@
 Now by a standard homological algebra argument (c.f.
-[BM23, Lemma 9.2]), $M_{1}\xrightarrow{d_{1}^{1}}N_{1}$ splits off as a contractible direct summand of $C$, a contradiction since $C$ was assumed to be indecomposable.
+[BM26, Lemma 9.2]), $M_{1}\xrightarrow{d_{1}^{1}}N_{1}$ splits off as a contractible direct summand of $C$, a contradiction since $C$ was assumed to be indecomposable.
 Therefore, $j=h_{C}(H)$.
@@
 Suppose for contradiction $C_{n}\cong P_{1}\oplus P_{2}$ for projective $kG$-modules $P_{1},P_{2}$.
-Since $\dim_{k}H_{n}(C)=1$, either $d_{n}|_{P_{1}}$ or $d_{n}|_{P_{2}}$ is injective, hence split injective, and it follows that a contractible chain complex containing $P_{1}$ or $P_{2}$ splits off from $C$, contradicting indecomposability of $C$.
+Since $\dim_{k}H_{n}(C)=1$, either $d_{n}|_{P_{1}}$ or $d_{n}|_{P_{2}}$ is injective, hence split injective, so a contractible chain complex containing $P_{1}$ or $P_{2}$ splits off from $C$, contradicting indecomposability of $C$.
 Thus, $C_{n}$ is indecomposable projective, as desired.
 ∎
-In the situation described in Proposition 4.1, $\ker(d_{n})$ is the unique submodule of $k[G/N]$ of $k$-dimension 1.
-We are now ready to generalize the $a_{N}$ and $b_{N}$ homomorphisms constructed in [BG25, Definition 12.3] for every effective endotrivial $C$ and ($p$-)subgroup $H\leq G$.
+In the situation described in 4.1, $\ker(d_{n})$ is the unique submodule of $k[G/N]$ of $k$-dimension 1.
+Remark 4.2.
+We recall the morphisms $a_{N},b_{N},c_{N}$ constructed by Balmer–Gallauer, [BG25, Definition 12.3].
+Recall that given $N\triangleleft G$ a subgroup of $G$ of index $p$, one has the endotrivial of length $2^{\prime}$ inflated from $C_{p}$
+$u_{N}=k[G/N]\to k[G/N]\to k,$
+inflated from a truncated periodic resolution of the trivial module $k$.
+The above example is in the setting of $p>2$.
+Then we have $2^{\prime}+1$ morphisms in $\mathscr{K}(G)$, $a_{N}\colon k\to u_{N}$, $b_{N}\colon k\to u_{N}[-2^{\prime}]$, and if $p>2$, $c_{N}\colon k\to u_{N}[-1]$.
+The morphism $a_{N}$ in degree 0 is the identity map $k\xrightarrow{\operatorname{id}}k$.
+The morphisms $b_{N}$ and $c_{N}$ (if it exists) are induced by the coaugmentation homomorphism $k\to k[G/N]$.
+For all $H\leq G$, exactly one of $\Psi^{H}(a_{N})$ or $\Psi^{H}(b_{N})$ is a quasi-isomorphism, c.f. [BG25, Proposition 12.9]. On the other hand, $c_{N}$ is $\otimes$-nilpotent.
+We now generalize these maps for all endotrivials, as announced by the introduction.
+The theorem statement does not capture all the properties of the maps, their construction in the proof will be relevant.
 Theorem 4.3.
@@
 $\iota^{H}_{C}\colon k[h_{C}(H)]\to C$
-unique up to scaling, such that $\Psi^{H}(\iota^{H}_{C})$ is an isomorphism $k[h_{C}(H)]\cong\Psi^{H}(C)$ in $\operatorname{D}_{b}(k[{G}/\!\!/{H}])$.
+such that $\Psi^{H}(\iota^{H}_{C})$ is an isomorphism $k[h_{C}(H)]\cong\Psi^{H}(C)$ in $\operatorname{D}_{b}(k[{G}/\!\!/{H}])$.
 We have $\iota^{H}_{C}=\iota^{{}^{g}H}_{C}$ for all $g\in G$.
@@
 Proof.
-Since $C$ is effective, we may assume (after possibly shifting $C$) by Corollary 3.12 that there exists an indecomposable chain complex of free $k\Gamma_{G}$-modules $C^{?}$ such that $C\cong C^{1}$.
+Since $C$ is effective, we may assume (after possibly shifting $C$) by Theorem 3.13 that there exists an indecomposable chain complex of free $k\Gamma_{G}$-modules $C^{?}$ such that $C\cong C^{1}$.
 Denote the differentials of $C^{?}$ by $d_{i}^{?}$.
-By Proposition 4.1, the chain complex of permutation k[G//H]k[{G}/\!\!/{H}]-modules $C^{H}$, which by Proposition 3.5 is isomorphic in $\mathscr{K}({G}/\!\!/{H})$ to $\Psi^{H}(C)$, contains up to homotopy a unique indecomposable k[G//H]k[{G}/\!\!/{H}]-module $M$ in top homological degree $h_{C}(H)$, and $\dim_{k}\ker(d_{h_{C}(H)}^{H})=1$.
-It follows by the stabilizers grow condition that in the canonical permutation basis of $C^{?}$, $M$ corresponds to a unique direct summand in homological degree $h_{C}(H)$.
-Therefore, we are in the situation of Proposition 3.7 after restricting $d_{i}^{?}$ to $M$.
-Choosing a nonzero $m\in\ker(d_{h_{C}(H)}^{H})$ (this choice is unique up to scaling), Proposition 3.7 asserts the existence of a $m^{\prime}\in\ker(d_{h_{C}(H)}^{1})$ generating a (unique) $k$-dimension one submodule of $C^{1}_{h_{C}(H)}$.
-We have a unique (up to scaling) nonzero homomorphism $\iota^{H}_{C}$ of chain complexes as follows.
+By 4.1, the chain complex of permutation k[G//H]k[{G}/\!\!/{H}]-modules $C^{H}$, which by 3.5 is isomorphic in $\mathscr{K}({G}/\!\!/{H})$ to $\Psi^{H}(C)$, contains up to homotopy (after removing contractible summands) a unique indecomposable k[G//H]k[{G}/\!\!/{H}]-module $M$ in top homological degree $h_{C}(H)$, and $\dim_{k}\ker(d_{h_{C}(H)}^{H})=1$.
+By the stabilizers grow condition, in the canonical permutation basis of $C^{?}$, $M$ corresponds to a unique direct summand in homological degree $h_{C}(H)$.
+Therefore, we are in the situation of 3.7 after restricting $d_{i}^{?}$ to $M$.
+Choosing a nonzero $m\in\ker(d_{h_{C}(H)}^{H})$, 3.7 asserts the existence of a $m^{\prime}\in\ker(d_{h_{C}(H)}^{1})$ generating a uniquely determined $k$-dimension one submodule of $C^{1}_{h_{C}(H)}$.
+We have a unique (up to scaling) nonzero homomorphism $\iota^{H}_{C}$ of chain complexes as follows:
 […]
@@
 The statement $\iota^{H}_{C}=\iota^{{}^{g}H}_{C}$ for all $g\in G$ follows since $c_{g}\circ\Psi^{H}=\Psi^{H^{\prime}}$ for any $g\in G$ satisfying $^{g}H=H^{\prime}$.
-The final statement follows from the final statement of Proposition 3.7.
+The final statement follows from the final statement of 3.7.
 ∎
 Remark 4.4.
-One has to take care in how the maps $\iota^{H}_{C}$ are constructed; such maps that become quasi-isomorphisms locally are non-unique (not even up to choice of identification as in [BG25, Remark 12.5]).
-For instance, let $G=D_{16}$, $C$ be the endotrivial
-$C:=kD_{16}\xrightarrow{d_{2}}k[D_{16}/H_{1}]\oplus k[D_{16}/H_{2}]\xrightarrow{d_{1}}k,$
-where $H_{1}$ and $H_{2}$ are non-conjugate non-central subgroups of order 2, $d_{2}$ is induced by the assignment $g\mapsto(gH_{1},gH_{2})$, and $d_{1}$ is induced by augmentation homomorphisms.
+We refer to the maps constructed in Theorem 4.3 as forerunners, and $\iota^{H}_{C}$ as the $H$-forerunner of $C$, as the maps herald the $p$-local isomorphisms $\hat{\Psi}^{H}(C)\cong k[h_{C}(H)]$.
+We have much to say about these maps.
+  1. (a)
+    One can also define these via the trace homomorphism
+    $t^{G}_{H}:\operatorname{Hom}_{\mathscr{K}(H)}(\operatorname{Res}^{G}_{H}C,\operatorname{Res}^{G}_{H}D)\to\operatorname{Hom}_{\mathscr{K}(G)}(C,D)$
+    defined by the assignment
+    $f\mapsto\sum_{g\in[G/H]}g\cdot f(g^{-1}\cdot-).$
+    See [BEN98a, Definition 3.6.2].
+Setting $C=k$ and $D$ a shifted effective endotrivial, we are in the situation of Theorem 4.3, and have a morphism $f\colon k\to C^{H}[-h_{C}(G)]\hookrightarrow C^{1}[-h_{C}(H)]$.
+This morphism is $K$-stable for some maximum $K\geq H$, hence can be regarded as a morphism $f\in\operatorname{Hom}_{\mathscr{K}(K)}(k,\operatorname{Res}^{G}_{K}C[-h_{C}(H)])$.
+Setting $\iota^{H}_{C}=t^{G}_{K}(f)$ obtains our desired morphism.
+  2. (b)
+    The forerunner construction is quite ad-hoc, and appears to not coincide with any obvious topological construction, as the map from the tensor unit is a coaugmentation homomorphism.
+Moreover, their construction depends on a certain choice of direct sum decomposition; although this will not matter for our purposes, it still is dissatisfying, and in general the uniqueness (even up to scalar) of the forerunners is not clear.
+Whether a more natural construction, or a topological interpretation of these maps exist, are pertinent questions; we expect answers in the affirmative for both.
+Such a construction would allow for a much better understanding of these morphisms.
+    For instance, with our current construction it is unclear whether the following natural property holds: given two effective endotrivials $C_{1},C_{2}$ and subgroup $H\leq G$, do we have an identification, up to scalars and contractibles, between the images of $\iota^{H}_{C_{1}\otimes C_{2}}$ and $\iota^{H}_{C_{1}}\iota^{H}_{C_{2}}$?
+  3. (c)
+    By 4.1 and Corollary 3.14, for any effective endotrivial $C$, $\iota^{1}_{C}$ and $\iota^{G}_{C}$ are unique up to scaling, since if $C$ is indecomposable, $C_{h_{C}(1)}$ is indecomposable and $C_{h_{C}(G)}=k$.
+It follows that $\iota^{H}_{C_{1}\otimes C_{2}}=\iota^{H}_{C_{1}}\iota^{H}_{C_{2}}$ up to scaling for $H=1$ and $H=G$.
+  4. (d)
+    One has to take care in how the forerunners are constructed, as such maps that become quasi-isomorphisms locally are non-unique (not even up to choice of identification as in [BG25, Remark 12.5]).
+For instance, let $G=D_{2^{n}}$ for $n\geq 3$, and $C$ be the endotrivial of Example 2.7.
+    $C:=kD_{16}\xrightarrow{}k[D_{16}/H_{1}]\oplus k[D_{16}/H_{2}]\xrightarrow{}k,$
+    where $H_{1}$ and $H_{2}$ are non-conjugate non-central subgroups of order 2 and the differentials are induced by augmentation homomorphisms.
 In this case, we could choose $\iota^{H_{1}}_{C}$ and $\iota^{H_{2}}_{C}$ to be the inclusion
@@
     In this case, both $\hat{\Psi}^{H_{1}}(\iota)$ and $\hat{\Psi}^{H_{2}}(\iota)$ are isomorphisms, a desired property.
-However, from the construction in Proposition 3.7, it follows that $\iota^{H_{1}}_{C}$ and $\iota^{H_{2}}_{C}$ satisfy $\Psi^{H_{1}}(\iota^{H_{2}}_{C})=0$ and vice versa, since neither $H_{1}\geq_{G}H_{2}$ or vice versa.
-Additionally, given two subgroups $K,H\leq G$, $\iota^{H}_{C}$ and $\iota^{K}_{C}$ may coincide.
-For instance, any shift of the tensor unit has $\iota^{H}_{C}=\iota^{G}_{C}$ for all subgroups $H\leq G$.
+However, from the construction in 3.7, one observes that $\iota^{H_{1}}_{C}$ and $\iota^{H_{2}}_{C}$ satisfy $\Psi^{H_{1}}(\iota^{H_{2}}_{C})=0$ and vice versa, since neither $H_{1}\geq_{G}H_{2}$ or vice versa.
+This property is established in full generality in Lemma 4.11, and are necessary for our open cover to stratify the closed points of $\operatorname{Spc}(\mathscr{K}(G))$.
+  5. (e)
+    Given two non-conjugate subgroups $K,H\leq G$, $\iota^{H}_{C}$ and $\iota^{K}_{C}$ may coincide.
+As an easy example, any shift of the tensor unit has $\iota^{H}_{C}=\iota^{G}_{C}$ for all subgroups $H\leq G$.
 Proposition 4.5.
 Let $H,K$ be subgroups of $G$ and let $C$ be an effective endotrivial of $kG$-modules.
-Then $\iota^{H}_{C}=\iota^{K}_{C}$ if and only if there exists a subgroup $B$ of $G$ such that $h_{C}(B)=h_{C}(H)=h_{C}(K)$ and $H,K\leq_{G}B$.
+Then we may construct $\iota^{H}_{C}=\iota^{K}_{C}$ if and only if there exists a subgroup $B\leq G$ such that $h_{C}(B)=h_{C}(H)=h_{C}(K)$ and $H,K\leq_{G}B$.
 Proof.
-The forward implication is straightforward by the construction in Proposition 4.2 - in particular, we may choose $B$ to be the vertex of the permutation $kG$-submodule containing the image of $\iota^{H}_{C}$.
+The forward implication follows from the construction in Theorem 4.3 - in particular, one may choose $B$ to be the vertex of the permutation $kG$-submodule containing the image of $\iota^{H}_{C}$.
 Conversely, assume $h_{C}(H)=h_{C}(K)$.
 Suppose there exists a subgroup $B$ satisfying $h_{C}(B)=h_{C}(H)=h_{C}(K)$ and $H,K\leq_{G}B$.
-From the construction in Proposition 4.2, it follows that the image of $\iota^{B}_{C}$ is the unique minimal submodule of a transitive permutation module isomorphic to $k[G/B^{\prime}]$ for some $B^{\prime}$ containing $B$.
-It follows that $\iota^{H}_{C}$ and $\iota^{K}_{C}$ also have image contained in this permutation module, and hence also have image the unique minimal submodule of $k[G/B^{\prime}]$, as desired.
+From the construction of Theorem 4.3, the image of $\iota^{B}_{C}$ is the unique minimal submodule of a transitive permutation module isomorphic to $k[G/B^{\prime}]$ for some $B^{\prime}$ containing $B$.
+Then $\iota^{H}_{C}$ and $\iota^{K}_{C}$ $p$-locally also have image contained in this permutation module, and hence also have image the unique minimal submodule of $k[G/B^{\prime}]$, as desired.
 ∎
-4.2. The open cover
-Construction 4.5.
-With the notation of [BG25, Definition 12.3], given any normal subgroup $N$ of $G$ of index $p$, the maps $a_{N}$ and $b_{N}$ are examples of such morphisms (however, $c_{N}$ is not!).
-For instance, the endotrivial
+4.2. An open cover of the spectrum
+Using the forerunners, we construct the promised open cover of $\operatorname{Spc}(\mathscr{K}(G))$.
+Recall that every prime of $\operatorname{Spc}(\mathscr{K}(G))$ is of the form $\mathscr{P}(H,\mathfrak{p}):=\hat{\psi}^{H}(\mathfrak{p})$ for some subgroup $H\leq G$ and $\mathfrak{p}\in V_{{G}/\!\!/{H}}:=\operatorname{Spc}(\operatorname{D}_{b}(kG))$, and $\mathscr{P}(H,\mathfrak{p})=\mathscr{P}(H^{\prime},\mathfrak{p}^{\prime})$ if and only if $H^{\prime}={}^{g}H$ and $\mathfrak{p}^{\prime}={}^{g}\mathfrak{p}$ for some $g\in G$.
+Construction 4.6.
+With the notation of [BG25, Definition 12.3], given any normal subgroup $N\leq G$ of index $p$, the maps $a_{N}$ and $b_{N}$ are examples of such morphisms (however, $c_{N}$ is not!). The endotrivial
 $u_{N}=k[G/N]\to k[G/N]\to k$
 (assuming $p$ odd) satisfies $a_{N}=\iota_{u_{N}}^{H}$ for any $H\not\leq N$ and $b_{N}=\iota_{u_{N}}^{H}$ for any $H\leq N$.
-Given this observation, we generalize the open cover of $\operatorname{Spc}(\mathscr{K})$ presented in [BG25, Proposition 13.11] as follows.
-Let $\mathcal{B}(G)$ denote the subset of the canonical $\mathbb{Z}$-basis of $\operatorname{CF}_{b}(G)$ not induced from the trivial $\mathbb{R}G$-module $\mathbb{R}$, i.e.
-the effective endotrivials arising from irreducible real representation spheres, excluding $k[1]$.
-Define an open of $\operatorname{Spc}(\mathscr{K})$ by
+Given this observation, we generalize the open cover of $\operatorname{Spc}(\mathscr{K}(G))$ presented in [BG25, Proposition 13.11] as follows.
+Let $\mathcal{B}(G)$ denote the subset of the canonical $\mathbb{Z}$-basis of $\operatorname{CF}_{b}(G)$ not induced from the trivial $\mathbb{R}G$-module $\mathbb{R}$, i.e., the effective endotrivials arising from irreducible real representation spheres, excluding $k[1]$.
+For any subgroup $H\leq G$, we define an open subset of $\operatorname{Spc}(\mathscr{K}(G))$ via
 $U(H):=\bigcap_{C\in\mathcal{B}(G)}\operatorname{open}(\iota_{C}^{H}).$
@@
 $\operatorname{open}(f):=\operatorname{open}(\operatorname{cone}(f))=\{\mathscr{P}\mid f\text{ is invertible in }\mathscr{K}(G)/\mathscr{P}\}.$
-Note that in defining the opens $U(H)$, the canonical $\mathbb{Z}$-basis element $k[1]\in\mathcal{B}(G)$ would play no significant role, so we may exclude $k[1]$ from $\mathcal{B}(G)$ without issue.
+Since $\iota^{H}_{C}=\iota^{{}^{g}H}_{C}$ for any $g\in G$, $U(H)=U({}^{g}H)$.
+Moreover, $\operatorname{open}(\iota_{k[1]}^{H})=\operatorname{Spc}(\mathscr{K}(G))$ for any subgroup $H\leq G$.
+Therefore, there is no harm in excluding $k[1]=C_{\mathbb{R}}$ from $\mathcal{B}(G)$; this exclusion is mostly for notation and indexing purposes.
 Remark 4.7.
-Let us recall the closed points of $\operatorname{Spc}(\mathscr{K})$.
-Recall that closed points are exactly the minimal primes for inclusion, and every prime contains a minimal prime.
-[BG25, Corollary 7.31] asserts that the minimal primes of $\operatorname{Spc}(\mathscr{K})$ are those explicitly of the form $\mathfrak{m}_{H}:=\mathscr{P}(H,0)$ for some subgroup $H$, where 0 denotes the unique closed point of $V_{{G}/\!\!/{H}}$.
-These closed points are precisely the kernel of the residue tt-functor $\mathbb{F}^{H}:=\operatorname{Res}_{1}^{{G}/\!\!/{H}}\circ\Psi^{H}=\Psi^{H}\circ\operatorname{Res}^{G}_{H}$ (see [BG25, Definition 7.26]).
-The collection $\{U(H)\}_{H\leq G}$ is an open cover of $\operatorname{Spc}(\mathscr{K})$, with the closed point $\mathfrak{m}_{H}$ belonging to $U(H)$.
+Let us recall the closed points of $\operatorname{Spc}(\mathscr{K}(G))$, i.e., the minimal primes with respect to inclusion.
+[BG25, Corollary 7.31] asserts that the minimal primes of $\operatorname{Spc}(\mathscr{K}(G))$ are those $\mathscr{M}(H):=\mathscr{P}(H,0)$ for some subgroup $H$, where 0 denotes the unique closed point of $V_{{G}/\!\!/{H}}$.
+The closed point $\mathscr{M}(H)$ is precisely the kernel of the residue tt-functor $\mathbb{F}^{H}:=\operatorname{Res}_{1}^{{G}/\!\!/{H}}\circ\Psi^{H}=\Psi^{H}\circ\operatorname{Res}^{G}_{H}$, see [BG25, Definition 7.26], and $\mathcal{M}(H)$ is also the kernel of $\hat{\Psi}^{H}$.
+The collection $\{U(H)\}_{H\leq G}$ is an open cover of $\operatorname{Spc}(\mathscr{K}(G))$, with the closed point $\mathscr{M}(H)$ belonging to $U(H)$.
 Proposition 4.8.
 Let $H$ be a subgroup of $G$ and $C$ be an effective endotrivial.
-Recall the residue tt-functor $\mathbb{F}^{H}\colon\operatorname{Res}^{{G}/\!\!/{H}}_{1}\circ\Psi^{H}\colon\mathscr{K}\to\operatorname{D}_{b}(k)$ of the closed point $\mathfrak{m}_{H}$.
+Recall the residue tt-functor $\mathbb{F}^{H}\colon\operatorname{Res}^{{G}/\!\!/{H}}_{1}\circ\Psi^{H}\colon\mathscr{K}(G)\to\operatorname{D}_{b}(k)$ of the closed point $\mathscr{M}(H)$.
 $\mathbb{F}^{H}(\iota^{H}_{C})$ is an isomorphism.
 Moreover, if $K$ is a subgroup of $G$ satisfying $h_{C}(K)\neq h_{C}(H)$, $\mathbb{F}^{K}(\iota^{H}_{C})$ is not an isomorphism.
-In particular, the open $U(H)$ contains $\mathfrak{m}_{H}$, therefore the set of opens $\{U(H)\}_{H\leq G}$ is an open cover of $\operatorname{Spc}(\mathscr{K})$.
+In particular, the open $U(H)$ contains $\mathscr{M}(H)$, therefore the set of opens $\{U(H)\}_{H\leq G}$ is an open cover of $\operatorname{Spc}(\mathscr{K}(G))$.
 Proof.
-The fact that $\mathbb{F}^{H}(\iota^{H}_{C})$ is an isomorphism follows immediately since $\Psi^{H}(\iota^{H}_{C})$ is an isomorphism.
+The fact that $\mathbb{F}^{H}(\iota^{H}_{C})$ is an isomorphism follows immediately since $\Psi^{H}(\iota^{H}_{C})$ is a quasi-isomorphism by construction.
 Moreover, if $h_{C}(K)\neq h_{C}(H)$, $\mathbb{F}^{K}(\iota^{H}_{C})$ cannot be an isomorphism since $k[h_{C}(H)]$ is not isomorphic to $\Psi^{K}(C)$ in $\operatorname{D}_{b}(k[{G}/\!\!/{K}])$.
-It follows by conservativity that for any effective endotrivial $C$, $\iota^{H}_{C}$ is an isomorphism in $\mathscr{K}/\mathfrak{m}_{H}$ since $\mathbb{F}^{H}$ is the residue functor of $\mathfrak{m}_{H}$, so $U(H)$ contains $\mathfrak{m}_{H}$.
-Therefore by general tt-geometry, $\{U(H)\}_{H\leq G}$ is an open cover of $\operatorname{Spc}(\mathscr{K})$, as every prime specializes to some $\mathfrak{m}_{H}$, which are precisely the closed points of $\operatorname{Spc}(\mathscr{K}(G)$.
+It follows by conservativity that for any effective endotrivial $C$, $\iota^{H}_{C}$ is an isomorphism in $\mathscr{K}/\mathscr{M}(H)$, as $\mathbb{F}^{H}$ is the residue functor of $\mathscr{M}(H)$, so $U(H)$ contains $\mathscr{M}(H)$.
+Therefore by general tt-geometry, $\{U(H)\}_{H\leq G}$ is an open cover of $\operatorname{Spc}(\mathscr{K}(G))$, as every prime specializes to some $\mathscr{M}(H)$, which are precisely the closed points of $\operatorname{Spc}(\mathscr{K}(G)$.
 ∎
-We next show that if $K,H$ are non-conjugate subgroups of $G$, then $\mathfrak{m}_{K}\not\in U(H)$.
+We next show that if $K,H$ are non-conjugate subgroups of $G$, then $\mathscr{M}(K)\not\in U(H)$.
 This takes a bit more work.
-It follows from the previous proposition that if there exists a Borel-Smith function $f$ for which $f(K)\neq f(H)$, then $\mathfrak{m}_{K}\not\in U(H)$.
+It follows from the previous proposition that if there exists a Borel-Smith function $f$ for which $f(K)\neq f(H)$, then $\mathscr{M}(K)\not\in U(H)$.
 However, this may not necessarily occur!
+We take a brief detour to explore this.
 Definition 4.9.
@@
 In this case, $C_{2}\times C_{2}$ acts on $D_{8}$ faithfully.
-$G$ has two nonconjugate subgroups isomorphic to $V_{4}$,
+$G$ has two nonconjugate subgroups isomorphic to $V_{4}=C_{2}\times C_{2}$,
 $H:=\{1,b,c,bc\},\quad K:=\{1,a^{2}b,a^{2}c,bc\},$
@@
 $\dim V^{H}=\frac{1}{|H|}\sum_{h\in H}\chi_{V}(h).$
-Since $\operatorname{im}(\dim)=\operatorname{CF}_{b}(G)$, every Borel-Smith function $f$ satisfies $f(H)=f(K)$. The author thanks Math.SE user testaccount for this example.
+Since $\operatorname{im}(\dim)=\operatorname{CF}_{b}(G)$, every Borel-Smith function $f$ satisfies $f(H)=f(K)$. We thank Math.SE user testaccount for finding and sharing this example [TES24].
 Lemma 4.11.
-Let $H,K$ be non-conjugate subgroups of $G$ with the same order, let $L$ be the smallest subgroup of $G$ containing both a conjugate of $H$ and a conjugate of $K$, and let $C$ be an effective endotrivial for which $h_{C}(L)<h_{C}(H)$.
-Then $\mathbb{F}^{K}(\iota^{H}_{C})$ and $\mathbb{F}^{H}(\iota^{K}_{C})$ are 0.
+Let $H,K$ be non-conjugate subgroups of $G$ with the same order, let $L$ be the smallest subgroup of $G$ for which $H,K\leq_{G}L$, and let $C$ be an effective endotrivial for which $h_{C}(L)<h_{C}(H)$.
+Then $\mathbb{F}^{K}(\iota^{H}_{C})=0$ and $\mathbb{F}^{H}(\iota^{K}_{C})=0$.
 Proof.
 It suffices to assume $h_{C}(H)=h_{C}(K)$.
-From Proposition 4.1 and the construction of $\iota^{H}_{C}$ in Theorem 4.2, the image of $\iota^{H}_{C}$ in homological degree $h_{C}(H)$ is contained in an indecomposable permutation $kG$-module $M$ with stabilizer containing $H$ but strictly contained in $L$.
-Therefore, $\Psi^{K}(M)=0$, so the image of $\mathbb{F}^{K}(\iota^{H}_{C})$ in degree $h_{C}(H)$ is the zero map, as desired.
-An analogous argument demonstrates $\mathbb{F}^{H}(\iota^{K}_{C})$ is also the zero map.
+From 4.1 and the construction of $\iota^{H}_{C}$ in Theorem 4.3, there exists a direct sum decomposition of $C_{h_{C}(H)}$ such that the image of $\iota^{H}_{C}$ in homological degree $h_{C}(H)$ is contained in an indecomposable permutation $kG$-module summand $M$, with stabilizer containing $H$ but strictly contained in $L$.
+Therefore, by minimality of $L$, $\Psi^{K}(M)=0$, so $\mathbb{F}^{K}(\iota^{H}_{C})=0$, as desired.
+An analogous argument demonstrates $\mathbb{F}^{H}(\iota^{K}_{C})=0$.
 ∎
-Corollary 4.11.
-The prime $\mathfrak{m}_{H}$ is the only closed point of $\operatorname{Spc}(\mathscr{K})$ contained in the open $U(H)$.
+Putting everything together, we show that our open cover stratifies the closed points.
+Theorem 4.12.
+The prime $\mathscr{M}(H)$ is the only closed point of $\operatorname{Spc}(\mathscr{K}(G))$ contained in the open $U(H)$.
 Proof.
-We previously showed in Proposition 4.7 that $\mathfrak{m}_{H}\in U(H)$.
-Since $U(H)$ is defined by iterating over the irreducible endotrivials, if $h_{C}(H)\neq h_{C}(K)$ for some irreducible endotrivial $C$, then $\mathfrak{m}_{K}\not\in U(H)$.
+We previously showed in 4.8 that $\mathscr{M}(H)\in U(H)$.
+Since $U(H)$ is defined by iterating over the irreducible endotrivials, if $h_{C}(H)\neq h_{C}(K)$ for some irreducible endotrivial $C$, then $\mathscr{M}(K)\not\in U(H)$.
 Therefore, it suffices to consider the case where $G$ has a pair of indistinguishable subgroups $H,K$.
@@
 Hence, there exists an element $C$ of the canonical $\mathbb{Z}$-basis $\mathcal{B}(G)\cup\{k[1]\}$ of $\operatorname{Pic}(\mathscr{K}(G))$ with $h_{C}(H)\neq h_{C}(L)$.
-The previous lemma implies $\mathbb{F}^{K}(\iota^{H}_{C})$ is not an isomorphism, and it follows by definition that $\mathfrak{m}_{K}\not\in U(H)$, as desired.
+The previous lemma implies $\mathbb{F}^{K}(\iota^{H}_{C})$ is not an isomorphism, so by definition, $\mathscr{M}(K)\not\in U(H)$, as desired.
 ∎
-Warning: Corollary 4.11 does not imply that the open $U(H)$ has a unique closed point viewed as a subspace of $\operatorname{Spc}(\mathscr{K})$.
-The open $U(H)$ has a minimal closed point if and only if $\mathscr{K}(G)|_{U(H)}\cong\mathscr{K}(G)/\mathfrak{m}_{H}$, i.e.
-$U(H)$ consists of all points specializing to $\mathfrak{m}_{H}$.
-In general this may not occur.
-Remark 4.12.
-Corollary 4.11 result does not hold for the open cover of [BG25, Section 13] for non elementary abelian $p$-groups.
-For instance, [BG25, Proposition 13.14] states the closed complement of the open $U^{\prime}(1)$ (we write $U^{\prime}(1)$ to denote the open $U_{G}(1)$ of [BG25, Proposition 13.14] in order to distinguish the covers apart) is the support of $\operatorname{kos}_{G}(F)$, where $F$ denotes the Frattini subgroup of $G$ (i.e.
-the intersection of all maximal subgroups of $G$).
-By [BG25, Corollary 7.17], we have $\mathfrak{m}_{H}\in\operatorname{supp}(\operatorname{kos}_{G}(F))$ if and only if $H\not\leq_{G}F$, hence $\mathfrak{m}_{H}\in U^{\prime}(1)$ if and only if $H\leq_{G}F$.
-Therefore, $U^{\prime}(1)$ contains a unique closed point of $\mathscr{K}(G)$ if and only if $F=1$ if and only if $G$ is an elementary abelian $p$-group.
-Corollary 4.13.
+Warning: Theorem 4.12 does not imply that the open $U(H)$ has a unique closed point, as a subspace of $\operatorname{Spc}(\mathscr{K}(G))$.
+The open $U(H)$ has a minimal closed point if and only if $U(H)$ consists of all points specializing to $\mathscr{M}(H)$.
+In general this may not occur; this can already be observed for elementary abelian $p$-groups.
+As a result, we conclude that all endotrivials for $p$-groups are line bundles.
+Theorem 4.13.
 Every element $C\in\operatorname{Pic}(\mathscr{K}(G))$ is a line bundle under the open cover $\{U(H)\}_{H\leq G}$.
@@
 Proof.
-By construction of $U(H)$, every irreducible endotrivial corresponding to an element of the canonical $\mathbb{Z}$-basis of $\operatorname{CF}_{b}(G)$ is isomorphic to $k[h_{C}(H)]$ in $\mathscr{K}(G)|_{U(H)}$.
-As the set of irreducible endotrivials is a basis of $\operatorname{Pic}(\mathscr{K}(G))$, it follows that every endotrivial $C$ is isomorphic to $k[h_{C}(H)].$
+By construction of $U(H)$ and the forerunners, every irreducible endotrivial corresponding to an element of the canonical $\mathbb{Z}$-basis of $\operatorname{CF}_{b}(G)$ is isomorphic to $k[h_{C}(H)]$ in $\mathscr{K}(G)|_{U(H)}$.
+As the set of irreducible endotrivials is a $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K}(G))$, it follows that every endotrivial $C\in\operatorname{Pic}(\mathscr{K}(G))$ satisfies $C\cong k[h_{C}(H)]$ in $\mathscr{K}(G)|_{U(H)}$.
+∎
 Remark 4.14.
-Corollary 4.13 does not imply the existence of a homomorphism $f:k[h_{C}(H)]\to C$ which is an isomorphism in $\mathscr{K}(G)|_{U(H)}$.
+Given an arbitrary endotrivial $C$, Theorem 4.13 does not imply the existence of an $H$-forerunner, i.e., a homomorphism $f:k[h_{C}(H)]\to C$ which descends to an isomorphism in $\mathscr{K}(G)|_{U(H)}$.
 Such homomorphisms may not exist in general for non-effective endotrivials.
 For example, let $p=2$ and $G=C_{2}$, then the two-term endotrivial $C:=k\xrightarrow{\operatorname{coaug}}kC_{2}$ with $k$ in homological degree 0 satisfies $\operatorname{Hom}_{\mathscr{K}(G)}(k,C)=0$, but we have a local isomorphism $k[0]\cong C$ in the localization $\mathscr{K}(G)|_{U(C_{2})}$.
-We turn to our open $U(1).$ First, we need a lemma about faithful endotrivials, i.e.
-endotrivials that arise from a faithful real representation $V$.
+Definition 4.15.
+We recall Balmer–Gallauer’s Koszul complexes: given a subgroup $H\leq G$, set
+$\operatorname{kos}_{H}(G):=\operatorname{Ten}^{G}_{H}(0\to k\xrightarrow{\operatorname{id}}k\to 0),$
+where $\operatorname{Ten}^{G}_{H}$ denotes tensor induction of chain complexes, see [BEN98b, Section 4.1].
+Note that while tensor induction is monoidal, it does not preserve contractibility, and hence is not functorial for $\mathscr{K}(G)$.
+One has that the tt-ideal $\ker(\operatorname{Res}^{G}_{H})$ is generated by $\operatorname{kos}_{G}(H)$ by [BG25, Proposition 3.21].
+In particular, $\operatorname{supp}(\operatorname{kos}_{G}(H))=\operatorname{supp}(\ker(\operatorname{Res}^{G}_{H}))$.
+Remark 4.16.
+Theorem 4.12 result does not hold for the open cover of [BG25, Section 13] for non elementary abelian $p$-groups, demonstrating that to stratify $\operatorname{Spc}(\mathscr{K}(G))$, one needs to consider more endotrivials than just the collection of $u_{N}$s.
+For instance, [BG25, Proposition 13.14] states the closed complement of the open $U^{\prime}(1)$ (writing $U^{\prime}(1)$ to denote the open $U_{G}(1)$ of [BG25, Proposition 13.14], in order to distinguish the open covers apart) is the support of $\operatorname{kos}_{G}(F)$, where $F$ denotes the Frattini subgroup of $G$ (i.e., the intersection of all maximal subgroups of $G$).
+By [BG25, Corollary 7.17], we have $\mathscr{M}(H)\in\operatorname{supp}(\operatorname{kos}_{G}(F))$ if and only if $H\not\leq_{G}F$, hence $\mathscr{M}(H)\in U^{\prime}(1)$ if and only if $H\leq_{G}F$.
+Therefore, $U^{\prime}(1)$ contains a unique closed point of $\mathscr{K}(G)$ if and only if $F=1$ if and only if $G$ is an elementary abelian $p$-group.
+We turn to our open $U(1).$ Note that for any effective endotrivial $C$, $\iota^{1}_{C}$ is unique up to scalar since $C_{h_{C}(1)}$ is indecomposable by 4.1.
+First, we need a lemma about faithful endotrivials, i.e., endotrivials that arise from a faithful real representation $V$.
 Equivalently, these are endotrivials whose h-marks are all 0 for any subgroup $H$ containing a nonzero normal subgroup of $G$, see [MIL24].
@@
 Proof.
-We have that $\ker(V)=1$, so by Proposition 4.1 if $n$ is the highest nonzero homological degree of $C$, $C_{n}\cong kG$.
-Now consider the complex $\operatorname{cone}(\iota^{1}_{C})^{*}$.
+We have that $\ker(V)=1$, so by 4.1 if $n$ is the highest nonzero homological degree of $C$, $C_{n}\cong kG$.
+Now consider the dual complex $\operatorname{cone}(\iota^{1}_{C})^{*}$.
 After shifting, this complex satisfies $C_{1}\cong kG$ and $C_{0}\cong k$.
-Therefore, we are in the situation of [BG25, Corollary 3.20], and it follows that $\operatorname{cone}(\iota^{1}_{C})^{*}$, hence $\operatorname{cone}(\iota^{1}_{C})$, generates $\mathscr{K}_{ac}(G)$ as a tt-ideal, and $\operatorname{supp}(\mathscr{K}_{ac}(G))=\operatorname{supp}(\operatorname{cone}(\iota^{1}_{C}))$, as desired.
+Therefore, we are in the situation of [BG25, Corollary 3.20], so $\operatorname{cone}(\iota^{1}_{C})^{*}$, hence $\operatorname{cone}(\iota^{1}_{C})$, generates $\mathscr{K}_{ac}(G)$ as a tt-ideal, and $\operatorname{supp}(\mathscr{K}_{ac}(G))=\operatorname{supp}(\operatorname{cone}(\iota^{1}_{C}))$, as desired.
 ∎
 Remark 4.18.
-Analogously, if $V$ is a real representation with kernel $N\trianglelefteq G$, then it is a faithful real representation of $G/N$, and the associated endotrivial $C$ is a faithful endotrivial complex of $k[G/N]$-modules.
-It follows that in this case, $\operatorname{cone}(\iota^{1}_{C})=\operatorname{cone}(\iota^{N}_{C})$ (as they are the same map) and $\operatorname{supp}(\ker(\operatorname{Res}^{G}_{N}))=\operatorname{supp}(\operatorname{kos}_{G}(N))$, recovering [BG25, Lemma 13.2].
+Analogously, if $V$ is a real representation with kernel $N\trianglelefteq G$, then it is a faithful real representation of $G/N$, and the associated endotrivial $C_{V}$ is a faithful endotrivial complex of $k[G/N]$-modules.
+In this case, $\operatorname{cone}(\iota^{1}_{C_{V}})=\operatorname{cone}(\iota^{N}_{C_{V}})$ (as they are the same map) and
+$\operatorname{supp}(\ker(\operatorname{Res}^{G}_{N}))=\operatorname{supp}(\operatorname{kos}_{G}(N))=\operatorname{supp}(\operatorname{Inf}^{G}_{G/N}(\operatorname{kos}_{G/N}(N/N))),$
+with the rightmost equality implied by [BG25, Lemma 3.17]. Therefore we have
+$\operatorname{supp}(\iota_{C_{V}}^{N})=\operatorname{supp}(\ker(\operatorname{Res}^{G}_{N})),$
+generalizing [BG25, Lemma 13.2(b)] to arbitrary endotrivials.
 Theorem 4.19.
 Let $G$ be a finite $p$-group.
-The closed complement of the open $U(1)$ is the support of $\operatorname{kos}_{G}(1)$, i.e.
-the closed support of the tt-ideal $\mathscr{K}_{ac}(G)=\ker(\operatorname{Res}^{G}_{1})$.
+The closed complement of the open $U(1)$ is the support of $\operatorname{kos}_{G}(1)$, i.e., the closed support of the tt-ideal $\mathscr{K}_{ac}(G)=\ker(\operatorname{Res}^{G}_{1})$.
 In particular, $U(1)$ is equal to the cohomological open $V_{G}=\operatorname{Spc}(\operatorname{D}_{b}(kG)).$
@@
 Since $\mathcal{B}(G)\cup\{k[1]\}$ is a basis of $\operatorname{Pic}(\mathscr{K}(G))$, for any $\mathscr{P}\in U(1)$, all endotrivials are isomorphic to a shift of the tensor unit in $\mathscr{K}(G)/\mathscr{P}$.
-In particular, there exists a faithful real representation $V$ (for instance the regular representation $\mathbb{R}G$) and a corresponding endotrivial $C_{V}$, for which $\iota_{C_{V}}^{1}\colon k[h_{C_{V}}(1)]\to C_{V}$ is an isomorphism in $\mathscr{K}(G)/\mathscr{P}$ (in fact, an isomorphism in $\mathscr{K}|_{U(1)}$ by Corollary 4.13).
+In particular, there exists a faithful real representation $V$ (for instance the regular representation $\mathbb{R}G$) and a corresponding endotrivial $C_{V}$, for which $\iota_{C_{V}}^{1}\colon k[h_{C_{V}}(1)]\to C_{V}$ is an isomorphism in $\mathscr{K}(G)/\mathscr{P}$ (in fact, an isomorphism in $\mathscr{K}(G)|_{U(1)}$ by Theorem 4.13).
 Therefore, $\operatorname{open}(\operatorname{cone}(\iota^{1}_{C_{V}}))\supseteq U(1)$.
-By Proposition 4.15, we have $\operatorname{supp}(\operatorname{kos}_{G}(1))=\operatorname{supp}(\operatorname{cone}(\iota_{C_{V}}^{1}))\subseteq U(1)^{c}$.
+By 4.17, we have $\operatorname{supp}(\operatorname{kos}_{G}(1))=\operatorname{supp}(\operatorname{cone}(\iota_{C_{V}}^{1}))\subseteq U(1)^{c}$.
 Conversely, by definition we have
@@
 ∎
+The following corollary is immediate by general tt-geometry, since $\mathscr{K}(G)$ is rigid.
 Corollary 4.20.
-We have an tt-equivalence $\mathscr{K}(G)|_{U(1)}\cong\operatorname{D}_{b}(kG)$.
-Proof.
-This follows immediately from Proposition 4.17 and general tt-geometry. ∎
-In particular, the open $U(1)$ has a unique closed point as a subspace of $\mathscr{K}(G)$. However, we stress that in general the other opens $U(H)$ need not have unique closed points.
-5. The (re-)twisted cohomology ring
-We are now ready to define the twisted cohomology ring for finite $p$-groups.
-When $G$ is a finite elementary abelian $p$-group, we recover the previously constructed twisted cohomology ring, since in this case the set of endotrivials $\{u_{N}\}_{N\in\mathcal{N}}$ as defined in [BG25, Example 12.1] and the shift of the tensor unit $k[1]$ form precisely the canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K}(G))$.
+We have a tt-equivalence $\mathscr{K}(G)|_{U(1)}\cong\operatorname{D}_{b}(kG)$.
+In particular, the open $U(1)$ has a unique closed point as a subspace of $\mathscr{K}(G)$.
+However, we stress that the other opens $U(H)$ need not have unique closed points.
+In particular, the spaces $\operatorname{Spc}(\mathscr{K}(G)/\mathscr{M}(H))\subseteq\operatorname{Spc}(\mathscr{K}(G)|_{U(H)})$ need not be homeomorphic, as the former has a unique closed point, but the latter may not.
+5. The remixed twisted cohomology ring
+We introduce the twisted cohomology ring for finite $p$-groups.
 Definition 5.1.
-Let $\mathbb{N}^{\mathcal{B}(G)}=\{q\colon\mathcal{B}(G)\to\mathbb{N}\}$ be the monoid of twists, i.e.
-tuples of non-negative integers indexed by the $\mathcal{B}(G)$.
+Let $\mathbb{N}^{\mathcal{B}(G)}=\{q\colon\mathcal{B}(G)\to\mathbb{N}\}$ be the monoid of twists, i.e., tuples of non-negative integers indexed by elements of $\mathcal{B}(G)$.
 Equivalently, the monoid of twists identifies with the submonoid of $\operatorname{Pic}(\mathscr{K}(G))$ generated by $\mathcal{B}(G)$.
@@
 $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)=\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G;k):=\bigoplus_{s\in\mathbb{Z}}\bigoplus_{q\in\mathbb{N}^{\mathcal{B}(G)}}\operatorname{Hom}_{\mathscr{K}(G)}\left(k,\bigotimes_{C\in\mathcal{B}(G)}C^{\otimes q(C)}[s]\right).$
-Its multiplication is induced by the tensor product in $\mathscr{K}(G)$.
+Multiplication is induced by the tensor product in $\mathscr{K}(G)$.
 Note that only non-positive shifts $s\leq 0$ produce non-zero homomorphisms.
-We call $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ the twisted cohomology ring of $G$.
+We call $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ the (permutation) twisted cohomology ring of $G$, as in [BG25].
 It is convenient to write
@@
 Remark 5.2.
-This construction may not immediately adapt if $G$ is not a $p$-group, since in this setting it is not known if one has a canonical $\mathbb{Z}$-basis for $\operatorname{Pic}(\mathscr{K}(G))$.
-A possible replacement could be the subgroup of $\mathscr{K}(G)$ of endotrivials arising from representation spheres for $G$.
-We propose that this subgroup has finite index in $\mathscr{K}(G),$ hence the analogous construction of an open cover should also hold.
-To examine the $k$-vector space $\operatorname{H}^{s,q}(G)$, it suffices to choose an indecomposable representative due to the stabilizers grow property.
+  1. (a)
+    When $G$ is a finite elementary abelian $p$-group, we recover the twisted cohomology ring constructed in [BG25], since in this case the set of endotrivials $\{u_{N}\}_{N\in\mathcal{N}}$ as defined in [BG25, Example 12.1] and the shift of the tensor unit $k[1]$ form precisely the canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K}(G))$, as deduced in [MIL24, Section 6].
+  2. (b)
+    This construction may not immediately adapt for non-$p$-groups, since in this setting it is not known if one has a canonical $\mathbb{Z}$-basis for $\operatorname{Pic}(\mathscr{K}(G))$.
+A possible replacement we propose could be the subgroup of $\mathscr{K}(G)$ of endotrivials arising from representation spheres for $G$.
+It is unknown if the subgroup of $\operatorname{Pic}(\mathscr{K}(G))$ spanned by such endotrivials even has maximal rank; we expect these questions to be nontrivial.
+To examine the $k$-vector space $\operatorname{H}^{s,q}(G)$, it suffices to choose an indecomposable representative of $k(q)$.
+This proposition also classifies all the global sections of an effective endotrivial, in the sense of [GAL25].
 Proposition 5.3.
@@
 Let $f\colon k\to C[s]$ be a nonzero chain complex homomorphism.
-We have that $f\in\operatorname{Hom}_{\mathscr{K}(G)}(k,C[s])=0$ if and only if $\operatorname{im}(f)\subset C_{s}$ is contained in a contractible direct summand of $C$.
-In particular, if $C$ is indecomposable, then
+We have that $f\in\operatorname{Hom}_{\mathscr{K}(G)}(k,C[s])=0$ if and only if $\operatorname{im}(f_{0})\subset C_{s}$ is contained in a contractible direct summand of $C$.
+In particular, if $C$ is an indecomposable effective endotrivial, then
 $\operatorname{H}^{s,q}(G):=\operatorname{Hom}_{\mathscr{K}(G)}(k,C[s])=\operatorname{Hom}_{\operatorname{Ch}(p-\operatorname{perm}(kG))}(k,C[s]).$
@@
 Suppose $f\in\operatorname{Hom}_{\mathscr{K}(G)}(k,C[s])=0$, then there exists a homotopy $h\colon k\to C_{s+1}$ such that $f=d_{s+1}\circ h$.
-Choose a canonical $\mathbb{Z}$-basis of $C$, and consider the projection of the image of $f$ onto each indecomposable summand.
+Choose a canonical basis of $C$, and consider the projection of the image of $f$ onto each indecomposable summand.
 Necessarily on each summand $k[G/H]$ whose projection is nonzero, the image must be the unique minimal nonzero submodule $k\subseteq k[G/H]$.
@@
 An induction argument demonstrates $\operatorname{im}(f)$ is contained in a contractible direct summand of $C$ as desired.
+The last statement follows from Corollary 3.15.
 ∎
-Just like in [Bal10a, BG25], we have a canonical comparison map.
+As in [BAL10a, BG25], we have a canonical comparison map, the heart of the matter.
+Its existence follows identically to [BG25, Proposition 13.4].
+See also [GAL25, Appendix A] for a much more general treatment of these twisted cohomology rings.
 Proposition 5.4.
@@
 where $Z(f)=\{\mathfrak{p}\mid f\in\mathfrak{p}\}$ is the closed subset of $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ defined by $f$.
-Proof.
-Same as [BG25, Proposition 13.4]. ∎
-Recall the conservative functor of Theorem 1.8, the collection of modular fixed points functors $\{\hat{\Psi}^{H}\}_{H\leq G}$.
-For shorthand, we denote this functor by $\prod\hat{\Psi}:\mathscr{K}(G)\to\prod_{H\leq G}\operatorname{D}_{b}({G}/\!\!/{H})$.
-We’ll describe the “twisted” analogue of the induced homomorphism on the respective cohomology rings.
-Definition 5.5.
-Let $C\in\mathcal{B}(G)$, $s\in\mathbb{Z}$, and $f\colon k\to C[s]$ be a homogeneous element of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$.
-For every subgroup $H\leq G$, $\hat{\Psi}^{H}(C)\in\operatorname{Hom}_{\operatorname{D}_{b}({G}/\!\!/{H})}(k,\Psi^{H}(C)[s])$, and since we have an isomorphism $\Psi^{H}(C)[s]\cong k[h_{C}(H)+s]$ (determined by $\iota^{H}_{C}$), this determines an element in group cohomology $f_{H}\in\operatorname{H}^{h_{C}(H)+s}({G}/\!\!/{H},k)=\operatorname{Ext}^{h_{C}(H)+s}_{k[{G}/\!\!/{H}]}(k,k)$ determined by the fraction
-$k\xrightarrow{\Psi^{H}(f)}C[s]\xleftarrow{\Psi^{H}(\iota^{H}_{C})}k[h_{C}(H)+s].$
-We consider $\hat{\Psi}^{H}(f)$ an element of $\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H})$, i.e.
-as a morphism in $\operatorname{D}_{b}(k[{G}/\!\!/{H}])$.
-Note that for the endotrivial $u_{N}$ defined in [BG25, Definition 12.3] and $H=1$, the fraction $\hat{\Psi}^{1}(\iota^{G}_{u_{N}})$ is precisely the fraction $\zeta_{N}^{+}$ of [BG25, Remark 12.7].
-If $H=G$, the fraction $\hat{\Psi}^{G}(\iota^{1}_{u_{N}})$ is precisely the fraction $\zeta_{N}^{-}$.
-Performing this construction over all subgroups $H\leq G$ up to conjugacy determines an element $\left(\prod\hat{\Psi}\right)(f)$ in the $\operatorname{CF}(G)$-graded ring
-$\operatorname{dH}^{{\scriptscriptstyle\bullet}}(G):=\prod_{H\in s_{p}(G)/G}\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H},k).$
-For $\prod\hat{\Psi}$ to be a well-defined ring homomorphism, the denominator associated to an arbitrary effective endotrivial $C$ is defined as the product of $\iota^{H}_{C}$s corresponding to the unique tensor product factorization of $C$ into a product of irreducible endotrivials.
-This product of $\iota^{H}$s remains a quasi-isomorphism.
-Note that this product is well-defined up to reordering, since $h_{C}$ returns only even values for $p$ odd.
-The assignment described describes the ring homomorphism $\prod\hat{\Psi}\colon\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{dH}^{\scriptscriptstyle\bullet}(G)$ induced by the conservative functor of [BG25, Theorem 7.2].
-It is clear that this homomorphism is linear with respect to addition.
-Linearity over multiplication follows from the following roof; let $C$ and $D$ be endotrivials, $H$ a subgroup of $G$, $i$ and $j$ integers, and $f\colon k\to C[i]$ and $g\colon k\to D[j]$ homogeneous elements of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$.
-For shorthand, we write $\iota^{H}_{C}$ and $\iota^{H}_{D}$ for the products of $\iota^{H}_{C^{\prime}}$ and $\iota^{H}_{D^{\prime}}$ running over all irreducible endotrivial $C^{\prime}$ and $D^{\prime}$ occurring in the tensor decompositions of $C$ and $D$ respectively.
-Then, the following roof commutes.
-[…]
-Note that $\prod\hat{\Psi}$ sends tensor products of homomorphisms to Yoneda products of cohomology classes, sums of homomorphisms to Baer sums, and sends a homogeneous element in homological degree $(s,k(q))$ to a homogeneous element in degree $\underline{s}+h_{k(q)}$, where $\underline{s}$ denotes the constant superclass function returning only $s$.
-Finally, we of course have a homeomorphism
-$\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))\cong\bigsqcup_{H\in s_{p}(G)/G}V_{{G}/\!\!/{H}}.$
-We denote the induced map on spectrum by $\prod\hat{\psi}\colon\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$, similar to the categorical case.
-Proposition 5.6.
-The kernel of $\prod\hat{\Psi}$ is $\otimes$-nilpotent, i.e. elements $f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ for which $f^{\otimes n}=0$ for some $n\geq 1$.
-Proof.
-This follows directly from [BG25, Theorems 7.1, 7.2], since $\prod\hat{\Psi}$ is induced from the conservative family of functors $\{\hat{\Psi}^{H}\}$. ∎
-Proposition 5.7.
-The following diagram in $\operatorname{Top}$ commutes.
-[…]
-In particular, the image of the comparison map $\operatorname{comp}_{G}$ is as a set, precisely the image of $\prod\hat{\psi}$.
-Moreover, the image of $\prod\hat{\psi}$ is dense in $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$.
-Proof.
-Every prime $\mathscr{P}(H,\mathfrak{p})\in\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))$ (in the notation of [BG25]) is explicitly the product of the prime $\mathfrak{p}$ of $\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H})$ and the full rings $\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{K})$ with $K$ different from $H$.
-The map $\prod\hat{\psi}$ sends the prime ideal $\mathscr{P}(H,\mathfrak{p})\in V_{{G}/\!\!/{H}}\subseteq\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))$ to the ideal $\{f\in\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))\mid\left(\prod\hat{\psi}\right)(f)\in\mathscr{P}(H,\mathfrak{p})\}$, and we have $\left(\prod\hat{\psi}\right)(f)\in\mathscr{P}(H,\mathfrak{p})$ if and only if $\hat{\Psi}^{H}(f)\in\mathfrak{p}$.
-Verification of commutativity now follows from the fact that given any $f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$, $H\leq G$, and $\mathfrak{p}\in V_{{G}/\!\!/{H}}$, we have $\operatorname{cone}(f)\not\in\mathscr{P}(H,\mathfrak{p})$ if and only if $\hat{\Psi}^{H}(f)\in\mathfrak{p}$.
-Indeed, it is well-known that the comparison map on the tt-category $\operatorname{D}_{b}({G}/\!\!/{H})$ is a homeomorphism $\operatorname{Spc}(\operatorname{D}_{b}({G}/\!\!/{H}))\xrightarrow{\cong}\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H}))$, sending a prime $\mathscr{P}(H,\mathfrak{p})$ to the prime $\mathfrak{p}=\{\zeta\mid\operatorname{cone}(\zeta)\not\in\mathfrak{p}\}\subset\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H},k)$.
-Thus, $\zeta\in\mathfrak{p}$ if and only if $\operatorname{cone}(\zeta)\not\in\mathscr{P}(H,\mathfrak{p})$, and setting $\zeta=\hat{\Psi}^{H}(f)$ shows the result.
-For the final statement, Proposition 5.6 implies the kernel of $d$ is nilpotent, so the result follows by a standard algebraic geometry result. ∎
 Remark 5.5.
-In particular, the image of the prime $\mathscr{P}(H,\mathfrak{p})$ under the comparison map $\operatorname{comp}_{G}$ is equivalently
-$\left(\prod\hat{\psi}\right)(\mathscr{P}(H,\mathfrak{p}))=\{f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\mid\hat{\Psi}^{H}(f)\in\mathfrak{p}\}.$
+Recall for a morphism $f$ in $\mathscr{K}(G)$ we write
+$\operatorname{open}(f):=\operatorname{open}(\operatorname{cone}(f))=\{\mathscr{P}\mid f\text{ is invertible in }\mathscr{K}(G)/\mathscr{P}\}.$
+The open $\operatorname{open}(f)$ is the preimage of $\operatorname{comp}_{G}$ of the principal quasi-compact open $Z(f)^{c}=\{\mathfrak{p}\mid f\not\in\mathfrak{p}\}$, and is equivalently the open locus of $\operatorname{Spc}(\mathscr{K}(G))$ where $f$ is invertible.
 5.1. Twisted cohomology under localization and tt-functors
-Next, we’ll extend the results of [BG25, Section 14] regarding localization.
+Next, we’ll extend the results of [BG25, Section 14] regarding localization and functoriality. The following constructions rely on the forerunners constructed in Theorem 4.3.
 Definition 5.6.
 Let $H$ be a subgroup of $G$.
-Let $S_{H}\subset\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ be the multiplicative subset generated by all $\iota^{H}_{C}$, where $C$ runs over the irreducible endotrivials $C\in\mathcal{B}(G)$.
+Let $S_{H}\subset\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ be the multiplicative subset generated by all forerunners $\iota^{H}_{C}$, where $C$ runs over the irreducible endotrivials $C\in\mathcal{B}(G)$.
 We define a $\mathbb{Z}$-graded ring
@@
 The homological degree of $\frac{f}{g}$ is the difference sts-t between the shifts of $f$ and $g$.
-In particular, $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(1)\cong\operatorname{H}^{\scriptscriptstyle\bullet}(G)$, following directly from Definition 5.5.
-Remark 5.10.
-Recall for a morphism $f$ in $\mathscr{K}(G)$ we write
-$\operatorname{open}(f):=\operatorname{open}(\operatorname{cone}(f))=\{\mathscr{P}\mid f\text{ is invertible in }\mathscr{K}(G)/\mathscr{P}\}.$
-We have that $\operatorname{open}(f)$ is the preimage of $\operatorname{comp}_{G}$ of the principle open $Z(f)^{c}=\{\mathfrak{p}\mid f\not\in\mathfrak{p}\}$, and is the open locus of $\operatorname{Spc}(\mathscr{K}(G))$ where $f$ is invertible.
-Construction 5.11.
-We can perform the central localization (see [Bal10a]) of the whole category $\mathscr{K}(G)$
+Construction 5.7.
+  1. (a)
+    We can perform the central localization (c.f. [BAL10a]) of the whole category $\mathscr{K}(G)$
     $\mathcal{L}_{G}(H):=\mathscr{K}(G)[S_{H}^{-1}].$
@@
 Moreover, the $\mathbb{Z}$-graded endomorphism ring $\operatorname{End}^{\scriptscriptstyle\bullet}_{\mathcal{L}_{G}(H)}(k)$ of the unit in $\mathcal{L}_{G}(H)$ is the $\mathbb{Z}$-graded ring $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$.
-Construction 5.12.
-Twisted cohomology $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is graded over a monoid over the form $\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)}$.
-The ring homomorphisms induced by tt-functors or localization will be homogeneous with respect to a certain homomorphism $\gamma$ on the corresponding grading monoids.
-Let $H\leq G$ be a subgroup and consider the central localization $(-)_{U(H)}\colon\mathscr{K}(G)\twoheadrightarrow\mathcal{L}_{G}(H)$.
-Here, the morphisms $\iota^{H}_{C}$ become isomorphisms, yielding a homomorphism on the grading
-$\gamma=\gamma_{U(H)}\colon\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)}\to\mathbb{Z}$
-defined by $\gamma(s,q)=s+h_{k(q)}(H)$ and we obtain a ring homomorphism
+In particular, $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(1)\cong\operatorname{H}^{\scriptscriptstyle\bullet}(G)$ by Theorem 4.19.
+We refer the reader to [BAL10a, Section 3] for this construction in abstract.
+  2. (b)
+    Twisted cohomology $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is graded over the monoid $\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)}$.
+Ring homomorphisms induced by tt-functors or localization will be homogeneous with respect to a certain homomorphism $\gamma$ on the corresponding grading monoids.
+    Let $H\leq G$ be a subgroup and consider the central localization
+    $(-)_{U(H)}\colon\mathscr{K}(G)\twoheadrightarrow\mathcal{L}_{G}(H).$
+    Here, the $H$-forerunners all become isomorphisms, yielding a homomorphism on the grading
+    $\gamma_{U(H)}\colon\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)}\to\mathbb{Z}$
+    defined by $\gamma_{U(H)}(s,q)=s+h_{k(q)}(H)$ and one obtains a ring homomorphism
     $(-)_{U(H)}\colon\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{End}^{{\scriptscriptstyle\bullet}}_{\mathcal{L}_{G}(H)}(k)=\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H),$
-homogeneous with respect to the homomorphism $\gamma_{U(H)}$.
-We may also consider the effects of restriction and modular fixed points on twisted cohomology.
-Note $\mathcal{B}(G)$ is more complicated than $\mathcal{N}\subseteq\mathcal{B}(G)$, where $\mathcal{N}$ consists of endotrivials inflated from normal subgroups of $G$ of index $p$, so in general, the shifts $\gamma$ may be more difficult to compute in the case of restriction.
-Construction 5.13.
+    homogeneous with respect to the monoid homomorphism $\gamma_{U(H)}$.
+  3. (c)
     Let $H\trianglelefteq G$ be a normal subgroup.
 The tt-functor $\Psi^{H}\colon\mathscr{K}(G)\to\mathscr{K}(G/H)$ maps effective endotrivials with h-mark at $G$ equal to 0 to effective endotrivials at $G/H$ equal to 0.
-In fact, since deflation preserves irreducibility of a representation, $\Psi^{H}$ sends elements of $\mathcal{B}(G)$ to $\mathcal{B}(G/H)$.
+In fact, since deflation preserves irreducibility of a representation, $\Psi^{H}$ induces a surjective map $\mathcal{B}(G)\to\mathcal{B}(G/H)\sqcup\{k[1]\}$.
     This defines a homomorphism of graded monoids
@@
     given by $\gamma(s,q)=(s,\overline{q})$ where $\overline{q}$ is given by the surjection $\mathbb{N}^{\mathcal{B}(G)}\twoheadrightarrow\mathbb{N}^{\mathcal{B}(G/H)}$ along the inclusion $\mathcal{B}(G/H)\hookrightarrow\mathcal{B}(G)$ induced from inflation.
-Therefore, modular fixed points defines a ring homomorphism $\Psi^{H}:\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G/H)$ homogeneous with respect to $\gamma_{\Psi^{H}}$.
-Given a group homomorphism $\alpha:G^{\prime}\to G$, restriction along $\alpha$ also defines a tt-functor $\alpha^{*}:\mathscr{K}(G)\to\mathscr{K}(G^{\prime})$.
-This again defines a corresponding ring homomorphism $\alpha^{*}:\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G^{\prime})$ homogeneous with respect to $\gamma_{\alpha^{*}}$.
-Note in this case if $\alpha$ is not surjective, then $\gamma_{\alpha^{*}}$ may not be constant on the shift $s$ and in general $\overline{q}$ will be difficult to compute.
-Conversely, if $\alpha$ is surjective, then $\alpha^{*}$ is simply inflation, which may be easily computed, as it is section of modular fixed points.
-It follows that the homomorphism $\Psi^{H}$ on twisted cohomology is split surjective.
+Therefore, modular fixed points defines a ring homomorphism $\Psi^{H}\colon\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G/H)$ homogeneous with respect to $\gamma_{\Psi^{H}}$.
+Because $\operatorname{id}_{\mathscr{K}(G/H)}=\Psi^{H}\circ\operatorname{Inf}^{G}_{G/H}$ general algebraic geometry implies $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G/H))$ is homeomorphically embedded into $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ via the map induced by $\Psi^{H}$ on spectra.
+  4. (d)
+    Given a group homomorphism $\alpha\colon G^{\prime}\to G$, restriction along $\alpha$ also defines a tt-functor $\alpha^{*}:\mathscr{K}(G)\to\mathscr{K}(G^{\prime})$.
+This again defines a corresponding ring homomorphism $\alpha^{*}\colon\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G^{\prime})$ homogeneous with respect to $\gamma_{\alpha^{*}}$.
+If $\alpha$ is surjective, then $\alpha^{*}$ is simply inflation $\operatorname{Inf}^{G}_{G/N}$ for $N=\ker(\alpha)$.
+If $\alpha$ is injective, then $\alpha^{*}$ is restriction.
+We write
+    $\rho^{G}_{H}:=\left(\operatorname{Res}^{G}_{H}\right)^{*}\colon\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)).$
+    These are essentially the only two cases one need consider, as every group homomorphism factors as a surjection composed with an injection.
+Remark 5.8.
+For any subgroup $H\leq G$, the tt-functor $\operatorname{Res}^{G}_{H}\colon\mathscr{K}(G)\to\mathscr{K}(H)$ induces a closed map on Balmer spectra $\operatorname{Spc}(\mathscr{K}(H))\to\operatorname{Spc}(\mathscr{K}(G))$ by [BG25, Proposition 4.7]; this follows by monadicity of the restriction/induction adjunction, see [BAL16].
+The same does not hold for the induced map on twisted cohomology $\rho^{G}_{H}\colon\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$.
+Indeed, let $G=C_{2}$, then $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(C_{2})\cong k[a,b]$ with $a,b$ the usual $a_{1},b_{1}$ morphisms in $\mathscr{K}(C_{2})$:
+[…]
+Now consider the restriction $\rho^{C_{2}}_{1}\colon\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(C_{2})\to\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(1)$.
+The image of $\rho^{C_{2}}_{1}$ is the prime $\langle b\rangle\subset\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(C_{2})$, which is neither an open nor closed point of $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(C_{2}))$.
 Remark 5.9.
@@
 $\operatorname{Spc}(\mathscr{K}(G^{\prime}))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G^{\prime}))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{F^{*}}$$\scriptstyle{\operatorname{comp}_{G^{\prime}}}$$\scriptstyle{\operatorname{comp}_{G}}$$\scriptstyle{F^{*}}$
-In particular, setting the functor $F$ equal to the localization functor $(-)|_{U(H)}$ obtains the following commutative square.
-$\operatorname{Spc}(\mathcal{L}_{G}(H))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\operatorname{comp}_{\mathcal{L}_{G}(H)}}$$\scriptstyle{\operatorname{comp}_{G}}$
-Here, the left hand vertical map is the usual twist-free comparison map from [Bal10a] for the tt-category $\mathcal{L}_{G}(H)$ and the $\otimes$-invertible $k[1]$.
-We combine the functors as before to strengthen [BG25, Proposition 14.21].
+Similarly, the localization functor $(-)|_{U(H)}$ obtains the following commutative square.
+$U(H)=\operatorname{Spc}(\mathcal{L}_{G}(H))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\operatorname{comp}_{\mathcal{L}_{G}(H)}}$$\scriptstyle{\operatorname{comp}_{G}}$
+Here, the left hand vertical map is the usual graded comparison map from [BAL10a] for the tt-category $\mathcal{L}_{G}(H)$.
+We have an analogous version of [BG25, Proposition 14.21].
 Proposition 5.10.
-Let $H\trianglelefteq G$ be a normal subgroup. Then we have a commutative square
-$V_{G/H}=\operatorname{Spc}(\operatorname{D}_{b}(k[G/H]))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}(G/H))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\hat{\psi^{H}}}$$\scriptstyle{\operatorname{comp}_{\operatorname{D}_{b}(k[G/H])}}$$\scriptstyle{\operatorname{comp}_{G}}$
-In particular, the diagonal is injective.
+Let $H\leq G$ be a subgroup. Then we have a commutative square
+$V_{{G}/\!\!/{H}}=\operatorname{Spc}(\operatorname{D}_{b}(k[{G}/\!\!/{H}]))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H}))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\hat{\psi^{H}}}$$\scriptstyle{\operatorname{comp}_{\operatorname{D}_{b}(k[{G}/\!\!/{H}])}}$$\scriptstyle{\operatorname{comp}_{G}}$
+If $H\trianglelefteq G$ is normal, then the diagonal is injective.
 Proof.
-The proof follows essentially the same as the proof of [BG25, Proposition 14.21], except without the added assumption that $G/H$ is elementary abelian.
-For this, we use Proposition 4.17 which asserts the quotient $\mathscr{K}(G/H)\to\operatorname{D}_{b}(k[G/H])$ is the central localization $(-)|_{U(1)}$ for any group $G/H$.
+The above diagram factors as follows:
+$\operatorname{Spc}(\operatorname{D}_{b}(k[{G}/\!\!/{H}]))$$\operatorname{Spc}(\mathscr{K}({G}/\!\!/{H}))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H}))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}({G}/\!\!/{H}))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\operatorname{comp}_{\operatorname{D}_{b}(k[{G}/\!\!/{H}])}}$$\scriptstyle{\psi^{H}}$$\scriptstyle{\operatorname{comp}_{{G}/\!\!/{H}}}$$\scriptstyle{\operatorname{comp}_{G}}$$\scriptstyle{\psi^{H}}$
+Here, the left-hand square arises from the central localization $(-)|_{U(H/H)}$ which by Theorem 4.19 is identified with the localization $\mathscr{K}({G}/\!\!/{H})\to\operatorname{D}_{b}(k[{G}/\!\!/{H}])$, and the right-hand square is induced from the tt-functor $\Psi^{H}\colon\mathscr{K}(G)\to\mathscr{K}({G}/\!\!/{H})$.
+Remark 5.9 asserts both diagrams commute, thus the square commutes as well.
+If $H\trianglelefteq G$, then $\Psi^{H}\colon\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G/H)$ is surjective, as noted in 5.7(c), hence $\psi^{H}:\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G/H))\hookrightarrow\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ is injective.
+Additionally, the comparison map $\operatorname{comp}_{\operatorname{D}_{b}(k[{G}/\!\!/{H}])}$ is always a homeomorphism, thus the bottom left composition is injective, and therefore so is the diagram.
+5.2. An identification for the image of the comparison map
+We may consider the commutative diagram of 5.10 for all conjugacy classes of subgroups of $G$ at once. Recall the conservative functor of Theorem 1.8, the collection of modular fixed points functors
+$\{\hat{\Psi}^{H}\}_{H\in s_{p}(G)/G}\colon\mathscr{K}(G)\to\prod_{H\in s_{p}(G)/G}\operatorname{D}_{b}(k[{G}/\!\!/{H}]).$
+For shorthand, we denote this functor by $\hat{\Psi}$. Denote the cohomology ring of $\prod_{H\in s_{p}(G)/G}\operatorname{D}_{b}(k[{G}/\!\!/{H}])$ by
+$\operatorname{dH}^{{\scriptscriptstyle\bullet}}(G)=\operatorname{dH}^{{\scriptscriptstyle\bullet}}(G,k)=\prod_{H\in s_{p}(G)/G}\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H},k).$
+We have an induced map on cohomology from 5.7, also denoted by $\hat{\Psi}$.
+$\hat{\Psi}\colon\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\xrightarrow{\prod\Psi^{H}}\prod_{H\in s_{p}(G)/G}\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}({G}/\!\!/{H})\to\prod_{H\in s_{p}(G)/G}\mathcal{O}_{{G}/\!\!/{H}}^{\scriptscriptstyle\bullet}(H/H)=\operatorname{dH}^{\scriptscriptstyle\bullet}(G).$
+We denote the associated map on homogeneous ring spectra by $\hat{\psi}^{h}\colon\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$.
+The following observation is easy but crucial, as it allows us to understand the comparison map $\operatorname{comp}_{G}$ purely in terms of induced maps on (Zariski) spectra.
+Injectivity of the diagonal is the crux of Section 6.
+Proposition 5.11.
+The following diagram in $\operatorname{Top}$ commutes.
+[…]
+In particular, the image of the comparison map $\operatorname{comp}_{G}$ is as a set, precisely the image of $\hat{\psi}^{h}$.
+Proof.
+Commutativity follows directly commutativity of the diagrams of 5.10 for all subgroups $H\leq G$. The latter statement follows since the top and left arrows are bijections of sets. ∎
+Remark 5.12.
+Unwinding things, the image of the prime $\mathscr{P}(H,\mathfrak{p})\in\operatorname{Spc}(\mathscr{K}(G))$, with $\mathfrak{p}$ regarded as an element of $\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H}))\cong V_{{G}/\!\!/{H}}$, under the comparison map $\operatorname{comp}_{G}$ is precisely
+$\hat{\psi}^{h}(\mathscr{P}(H,\mathfrak{p}))=\langle f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\mid\hat{\Psi}^{H}(f)\in\mathfrak{p}\rangle.$
+While the comparison map is in general not surjective, we deduce that it has dense image.
+The proof of this mirrors a standard commutative algebra fact, but with a slight modification due to the grading.
+Corollary 5.13.
+The image of the comparison map, $\operatorname{im}(\operatorname{comp}_{G})$, has dense image in $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$.
+Proof.
+Let $f\in\operatorname{H}^{s,q}(G)$ be a homogeneous element of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$.
+By [BG25, Theorems 7.1, 7.2], if $\hat{\Psi}(f)=0$, then $f$ is $\otimes$-nilpotent.
+We claim this fact is sufficient to prove that $\operatorname{im}(\operatorname{comp}_{G})=\operatorname{im}(\hat{\psi}^{h})$ is dense in $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$.
+Indeed, this follows by a standard argument.
+Recall that $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ has an open basis given by supports $\operatorname{supp}(f)$ of homogeneous $f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$.
+Therefore, for all homogeneous $f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$, if $\operatorname{supp}(f)\neq\emptyset$, then $f\not\in\mathfrak{p}$ for some homogeneous $\mathfrak{p}\in\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$.
+Hence $f$ is not $\otimes$-nilpotent, and it follows by our fact that neither is $\hat{\Psi}(f)$, so there exists a prime $\mathfrak{q}\in\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))$ with $\hat{\Psi}(f)\not\in\mathfrak{q}$.
+Therefore $f\not\in\hat{\psi}^{h}(\mathfrak{q})$, so $\hat{\psi}^{h}(\mathfrak{q})\in\operatorname{supp}(f)$.
+Thus $\operatorname{im}(\hat{\psi}^{h})$ intersects every nonempty basic open set $\operatorname{supp}(f)$, as desired.
 ∎
+Remark 5.14.
+[…]
 6. Injectivity of the comparison map
-The crux of injectivity of the comparison map is showing that, the diagonal of Proposition 5.15 is injective when $H$ is not necessarily normal in $G$ as well.
+The crux of injectivity of the comparison map is showing the diagonal of 5.10 is injective when a subgroup $H\leq G$ is not necessarily normal.
+To show this, we exploit the fact that $p$-groups are nilpotent and induct up a subnormal chain of subgroups of $G$.
+We first make a seemingly straightforward but subtle observation.
+Proposition 6.1.
+Let $H\leq G$, $C^{?}$ be a chain complex of free $k\Gamma_{G}$-modules, and set $C:=C^{1}$.
+Then $f\in\operatorname{Hom}_{\mathscr{K}(H)}(k,\operatorname{Res}^{G}_{H}C[s])$ is $G$-stable if and only if there exists a $\hat{f}\in\operatorname{Hom}_{\mathscr{K}(G)}(k,C[s])$ satisfying $\operatorname{Res}^{G}_{H}\hat{f}=f$.
+Proof.
+If one knows that $f$ is $G$-stable on the nose, the result is clear, but $f$ may only be $G$-stable only up to homotopy.
+By choosing a subnormal sequence $H:=H_{0}\triangleleft H_{1}\triangleleft\cdots\triangleleft H_{n}=G$ with $H_{i+1}:H_{i}]=p$, it suffices to assume $[G:H]=p$, as the result can be inductively applied to the subnormal sequence.
+We show $f$ is homotopic to a $G$-stable chain complex homomorphism.
+We have a direct sum decomposition $\operatorname{Res}^{G}_{H}C=D\oplus K$, where $D$ has no contractible direct summands.
+Let $p_{D}$ and $p_{K}$ denote projection onto $D$ and $K$ respectively followed by inclusion into $C$.
+If $f$ is $G$-stable in $\mathscr{K}(G)$, then by 5.3, for all $g\in G$, $f$ and $^{g}f$ agree on $D$, that is, $p_{D}\circ{}^{g}f=p_{D}\circ f$ as chain complex homomorphisms for all $g\in G$, where $p_{D}$ denotes projection onto $D$ followed by inclusion into $\operatorname{Res}^{G}_{H}C$.
+Of course, $p_{D}+p_{K}=\operatorname{id}$.
+[…]
+[…]
+[…]
+is homotopic to $f$.
+We note the exclusion of the term $1H\in[G/H]$ is purely for making computation easier, since $p_{K}\circ p_{D}=0$.
+We claim $\hat{f}$ is also a $G$-stable chain complex homomorphism.
+Indeed, for all $g\in G$, we compute for $p_{K}$ and $p_{D}$:
+$p_{D}\circ{}^{g}\hat{f}=p_{D}\circ{}^{g}(p_{D}\circ f)+p_{D}\circ{}^{g}p_{K}\circ\left(\sum_{g^{\prime}\in[G/H]-gH}{}^{g^{\prime}}(p_{D}\circ f)\right)=p_{D}\circ{}^{g}(p_{D}\circ f)=p_{D}\circ\hat{f}$,
+and
+[…]
+[…]
+[…]
+$=p_{K}\circ\hat{f}$.
 Theorem 6.2.
@@
 $\operatorname{comp}_{G}\colon\operatorname{Spc}(\mathscr{K}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$
-of Proposition 5.4 is injective.
+of 5.4 is injective.
 Proof.
@@
 Suppose for contradiction that $H_{2}\not\leq_{G}H_{1}$.
-By Corollary 4.11, there exists an effective endotrivial $C\in\mathcal{B}(G)$ for which $\Psi^{H_{2}}(\iota_{C}^{H_{1}})$ is the zero map, and vice versa.
+By Theorem 4.12, there exists an effective endotrivial $C\in\mathcal{B}(G)$ for which $\Psi^{H_{2}}(\iota_{C}^{H_{1}})$ is the zero map, and vice versa.
 Now, consider the map $\hat{\psi}^{H_{2}}\colon V_{{G}/\!\!/{H_{2}}}\hookrightarrow\operatorname{Spc}(\mathscr{K}(G))$.
@@
 Therefore, $V_{{G}/\!\!/{H_{2}}}\cap\operatorname{open}(\iota^{H_{1}}_{C})=\emptyset$ in $\operatorname{Spc}(\mathscr{K}(G))$.
-On the other hand, $V_{{G}/\!\!/{H_{1}}}\subseteq\operatorname{open}(\iota^{H_{1}}_{C})$, since $\iota^{H_{1}}_{C}$ is invertible in $\mathscr{K}(G)/\mathfrak{m}_{H_{1}}$.
+On the other hand, $V_{{G}/\!\!/{H_{1}}}\subseteq\operatorname{open}(\iota^{H_{1}}_{C})$, since $\iota^{H_{1}}_{C}$ is invertible in $\mathscr{K}(G)/\mathscr{M}(H_{1})$.
 Therefore, $\mathscr{P}\in\operatorname{open}(\iota^{H_{1}}_{C})$ but $\mathscr{Q}\not\in\operatorname{open}(\iota^{H_{1}}_{C})$, a contradiction.
@@
 $V_{{G}/\!\!/{H}}\xrightarrow{\hat{\psi}^{H}}\operatorname{Spc}(\mathscr{K}(G))\xrightarrow{\operatorname{comp}_{G}}\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$
-and this map is injective from Proposition 5.15. Otherwise, we have the following commutative diagram.
-$V_{N_{G}(H)/H}$$\operatorname{Spc}(\mathscr{K}(N_{G}(H))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(N_{G}(H)))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\operatorname{Res}^{*}}$$\scriptstyle{\operatorname{comp}_{N_{G}(H)}}$$\scriptstyle{\operatorname{comp}_{G}}$$\scriptstyle{\operatorname{Res}^{*}}$
+and this map is injective from 5.10. Otherwise, we have the following commutative diagram.
+$V_{N_{G}(H)/H}$$\operatorname{Spc}(\mathscr{K}(N_{G}(H))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(N_{G}(H)))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\rho^{G}_{N_{G}(H)}}$$\scriptstyle{\operatorname{comp}_{N_{G}(H)}}$$\scriptstyle{\operatorname{comp}_{G}}$$\scriptstyle{\rho^{G}_{N_{G}(H)}}$
 Here, the diagonal arrow is injective since $H\trianglelefteq N_{G}(H)$, and the top row is injective since $\operatorname{Spc}(\mathscr{K}(G))=\bigsqcup_{H\in s_{p}(G)/G}V_{{G}/\!\!/{H}}$, and the composition of the above maps is precisely $\hat{\psi}^{H}$.
@@
 $V_{N_{G}(H)/H}\hookrightarrow\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(N_{G}(H)))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$
-is injective. We establish this as a separate lemma.
+is injective. We establish this as a separate lemma below. ∎
 Lemma 6.3.
@@
 We have the following setup depicted by the commutative diagram below.
-$V_{N_{G}(H)/H}$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{0}))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{1}))$$\cdots$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{n}))$}
-Here, the horizontal arrows are induced by restriction and the downwards arrows are $\hat{\psi}^{H}$.
+$V_{N_{G}(H)/H}$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{0}))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{1}))$$\cdots$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{n}))$$\scriptstyle{\rho_{H_{0}}^{H_{1}}}$$\scriptstyle{\rho_{H_{1}}^{H_{2}}}$$\scriptstyle{\rho_{H_{n-1}}^{H_{n}}}$
+Here we identify $V_{N_{G}(H)/H}=\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}(N_{G}(H)/H,k))$ via 5.11, the horizontal arrows are induced by restriction, and the downwards arrows are each the map $\hat{\psi}^{H}$.
 We will prove this statement inductively by assuming that the map $V_{N_{G}(H)}\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))$ is injective and showing that the composition $V_{N_{G}(H)}\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1}))$ is injective, which then of course implies the next map $V_{N_{G}(H)}\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1}))$ is injective as well.
-Injectivity of the base case $H_{0}=N_{G}(H)$ is established by Proposition 5.15.
+Injectivity of the base case $H_{0}=N_{G}(H)$ is established by 5.10.
 Let $\mathfrak{p},\mathfrak{q}\in\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))$.
-We claim that $\operatorname{Spec}(\operatorname{Res}^{H_{i+1}}_{H_{i}})(\mathfrak{p})=\operatorname{Spec}(\operatorname{Res}^{H_{i+1}}_{H_{i}})(\mathfrak{q})$ if and only if the following property $(\ast)$ is satisfied:
+We claim that $\rho_{H_{i}}^{H_{i+1}}(\mathfrak{p})=\rho_{H_{i}}^{H_{i+1}}(\mathfrak{q})$ if and only if the following property $(\ast)$ is satisfied for the pair $(\mathfrak{p},\mathfrak{q})$:
 $(\ast)$  For every $f\in\mathfrak{p}$, there exists a $g\in H_{i+1}$ such that $^{g}f\in\mathfrak{q}$.
-The converse is straightforward. Suppose there exists an $f\colon k\to C[s]\in\mathfrak{p}$ for which $^{g}f\not\in\mathfrak{q}$ for all $g\in H_{i+1}$. Then, the trace product of $H_{i+1}$-conjugates
+The converse is straightforward.
+We show the contrapositive: suppose there exists an $f\colon k\to C[s]\in\mathfrak{p}$ for which $^{g}f\not\in\mathfrak{q}$ for all $g\in H_{i+1}$.
+Then, the ‘norm’ product of $H_{i+1}$-conjugates
 $f^{\prime}:=\prod_{g\in H_{i+1}/H_{i+1}}{}^{g}f$
-is a morphism $f^{\prime}\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1})$. Indeed, the tensor product of chain complexes
+may be lifted to a morphism $f^{\prime}\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1})$. Indeed, the tensor product of chain complexes
 […]
-has $kH_{i+1}$-module structure (one may see for instance that $C^{\prime}$ is the restriction of the tensor induced complex $\operatorname{Ten}^{H_{i+1}}_{H_{i}}C$ via the Mackey formula - while tensor induction is not defined in general up to homotopy, in this case the restriction is well-behaved up to homotopy), and the image of $k$ in $C^{\prime}$ is $H_{i+1}$-stable.
-By primality of $\mathfrak{p}$ and $\mathfrak{q}$, we have $f^{\prime}\in\mathfrak{p}$ but $f^{\prime}\not\in\mathfrak{q}$, as desired.
+is up to homotopy in the image of restriction from $kH_{i+1}$.
+This follows from the Mackey formula for Borel-Smith functions, as $\operatorname{CF}_{b}(-)$ is a rational $p$-biset functor (see [BY07] - in particular it is a Mackey functor) so we have,
+$\operatorname{Res}^{H_{i+1}}_{H_{i}}\operatorname{Ind}^{H_{i+1}}_{H_{i}}h_{C}=\sum_{g\in H_{i+1}/H_{i}}{}^{g}h_{C},$
+and we have an identification of rational $p$-biset functors $\operatorname{CF}_{b}(-)\cong\operatorname{Pic}(\mathscr{K}(-))$ from [MIL25b, Section 5].
+Since the image of $f^{\prime}$ in $C^{\prime}$ is $H_{i+1}$-stable, the ‘norm’ product $f^{\prime}\colon k\mapsto C^{\prime}$ is in the image of restriction by 6.1.
+Now by primality of $\mathfrak{p}$ and $\mathfrak{q}$, we have $f^{\prime}\in\mathfrak{p}$ but $f^{\prime}\not\in\mathfrak{q}$, which is what we wanted.
 Now, we show that if $\mathfrak{p},\mathfrak{q}\in\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))$ satisfy $(\ast)$ and belong to the image of $V_{N_{G}(H)/H}$, then $\mathfrak{p}=\mathfrak{q}$.
-The statement is trivial if $H_{i+1}/H_{i}$ acts on $H_{i}$ trivially, so assume the action is nontrivial.
-Recall from Proposition 5.7 that in this case, $\mathfrak{p}=\rho(H,\overline{\mathfrak{p}})$ with $\overline{\mathfrak{p}}\in V_{N_{G}(H)/H}$, and we have $\mathfrak{p}=\{f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i})\mid\hat{\Psi}^{H}(f)\in\overline{\mathfrak{p}}\}$.
+The statement is straightforward if $H_{i+1}/H_{i}$ acts trivially on $H_{i}$, so let us consider the case where the action is nontrivial.
+From 5.11, we have $\mathfrak{p}=\mathscr{P}(H,\overline{\mathfrak{p}})$ with $\overline{\mathfrak{p}}\in V_{N_{G}(H)/H}$, and we have $\mathfrak{p}=\{f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i})\mid\hat{\Psi}^{H}(f)\in\overline{\mathfrak{p}}\}$.
 Let $f\in\mathfrak{p}$.
-By Lemma 4.10, there exists an effective endotrivial chain complex $C$ of $kH_{i}$-modules and morphism $\iota^{H}_{C}$ such that $\hat{\Psi}^{H}(\iota^{H}_{C})=\operatorname{id}_{k}$ and $\hat{\Psi}^{H}({}^{g}(\iota^{H}_{C}))=0$ for all $g\in H_{i+1}\setminus H_{i}$.
-Therefore, $\iota^{H}_{C}\not\in\mathfrak{p}$ (since $\overline{\mathfrak{p}}\neq V_{N_{G}(H)/H}$), but for all $g\in H_{i+1}\setminus H_{i}$, we have $^{g}(\iota^{H}_{C})\in\varphi(H,0)\subseteq\mathfrak{p}$.
+By Lemma 4.11, there exists an effective endotrivial chain complex $C$ of $kH_{i}$-modules and morphism $\iota^{H}_{C}$ such that $\hat{\Psi}^{H}(\iota^{H}_{C})=\operatorname{id}_{k}$ and $\hat{\Psi}^{{}^{g}H}(\iota^{H}_{C})=0$ for all $g\in H_{i+1}\setminus H_{i}$, noting that $g\not\in N_{G}(H)$ so $^{g}H\neq H$.
+For instance, one may take the endotrivial $C$ arising from the regular representation $kH_{i}$.
+Therefore, $\iota^{H}_{C}\not\in\mathfrak{p}$, but for all $g\in H_{i+1}\setminus H_{i}$, we have $^{g}(\iota^{H}_{C})\in\mathscr{P}(H,0)\subseteq\mathfrak{p}$.
 We construct a morphism $f^{\prime\prime}\in\mathfrak{p}$ as follows,
@@
 By construction $\hat{\Psi}^{({}^{g}H)}(f^{\prime\prime})$ is an isomorphism for all $g\in H_{i+1}\setminus H_{i}$, therefore $^{g}(f^{\prime\prime})\not\in\mathfrak{q}$.
-But since $\mathfrak{p},\mathfrak{q}$ satisfy $(\ast)$, it follows that $f^{\prime\prime}\in\mathfrak{q}$ as well.
-Since each $^{g}(\iota^{H}_{C})\in\varphi(H,0)\subseteq\mathfrak{q}$, cancellation implies $f\in\mathfrak{q}$ as well.
-Thus $\mathfrak{p}\subseteq\mathfrak{q}$, and symmetry implies $\mathfrak{p}=\mathfrak{q}$, and the composition $V_{N_{G}(H)}\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1}))$ is injective, as desired.
+But since $(\mathfrak{p},\mathfrak{q})$ satisfies $(\ast)$, it must be the case that $f^{\prime\prime}\in\mathfrak{q}$.
+Since each $^{g}(\iota^{H}_{C})\in\mathscr{P}(H,0)\subseteq\mathfrak{q}$, cancellation or homogeneity of $\mathfrak{p}$ implies $f\in\mathfrak{q}$ as well.
+Thus $\mathfrak{p}\subseteq\mathfrak{q}$, and symmetry implies $\mathfrak{p}=\mathfrak{q}$.
+Putting everything together, we conclude the composition $V_{N_{G}(H)}\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1}))$ is injective.
 ∎
-As an immediate corollary, we obtain a strengthening of Proposition 5.15.
+Note that the restriction $\rho_{N_{G}(H)}^{G}\colon\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(N_{G}(H)))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ need not be injective, as fusion phenomena will occur.
+As an immediate corollary, we strengthen 5.10 and 5.11.
 Corollary 6.4.
-Let $H$ be any subgroup of $G$. Then we have a commutative square
+Let $H\leq G$ be any subgroup. Then the commutative square
 $V_{{G}/\!\!/{H}}=\operatorname{Spc}(\operatorname{D}_{b}(k[{G}/\!\!/{H}]))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H}))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\hat{\psi^{H}}}$$\scriptstyle{\operatorname{comp}_{\operatorname{D}_{b}(k[{G}/\!\!/{H}])}}$$\scriptstyle{\operatorname{comp}_{G}}$
-In particular, the diagonal is injective, and the product
-$\prod\hat{\psi}\colon\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$
-is injective.
+satisfies that the diagonal is injective. Moreover,
+$\hat{\psi}^{h}\colon\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$
+is injective with dense image.
+Remark 6.5.
+The comparison map is an injective map of spectral spaces, therefore Stone duality implies the pullback
+$\Omega(\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)))\to\Omega(\operatorname{Spc}(\mathscr{K}(G))),\,U\mapsto\operatorname{comp}_{G}^{-1}(U)$
+is an epimorphism in CFrm, where $\Omega(X)$ denotes the coherent frame of open subsets of a spectral topological space $X$.
+However, epimorphisms in the category CFrm of bounded distributive lattices need not be surjective.
+This is in contrast to the category of posets, where epimorphisms are precisely the surjective maps.
+In fact, if the induced map on opens is surjective, then by definition, the submonoid of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ of effective endotrivials is by definition an ample family [GAL25, Definition 5.4], and by [GAL25, Proposition 5.8], the comparison map would be an open immersion.
+Therefore, in order to completely deduce the topology of $\operatorname{Spc}(\mathscr{K}(G))$, more work is required.
+On the other hand, by [DST19, Theorem 5.2.2], we have the comparison map induces a surjective map
+$\mathcal{K}(\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)))\to\mathcal{K}(\operatorname{Spc}(\mathscr{K}(G))),$
+where $\mathcal{K}(X)$ denotes the poset of clopen subsets of $X$.
+If we know twisted cohomology is Noetherian, we further can show that the comparison map restricts to a homeomorphism upon localization to the opens $U(H)$, and therefore is an open immersion, extending Balmer–Gallauer’s [BG25, Theorem 15.3].
+The proof of this is quite short, thanks to machinery recently developed by Sanders [SAN25] regarding fully faithful tt-functors.
 Corollary 6.6.
@@
 $\operatorname{comp}_{G}\colon U(H)\cong\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)).$
-In particular, $\operatorname{comp}_{G}\colon\operatorname{Spc}(\mathscr{K}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ is an open immersion, with image the open subspace of $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ with closed points (maximal primes)
-$\operatorname{comp}_{G}(\mathfrak{m}_{H})=\{f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\mid\Psi^{H}(f)\text{ is not a quasi-isomorphism}\}.$
+In particular, in this case $\operatorname{comp}_{G}\colon\operatorname{Spc}(\mathscr{K}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ is an open immersion, with image the open subspace of $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ with closed points
+$\operatorname{comp}_{G}(\mathscr{M}(H))=\{f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\mid\Psi^{H}(f)\text{ is not a quasi-isomorphism}\}.$
 Proof.
 We have already an injective map by Remark 5.9 and Theorem 6.2
-$\operatorname{comp}_{\mathscr{L}_{G}(H)}:\operatorname{Spc}(\mathscr{L}_{G}(H))=U(H)\to\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H))).$
-Noetherianity of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ implies $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$ is noetherian as well, hence [Bal10a, Theorem 7.3] implies $\operatorname{comp}_{\mathscr{L}_{G}(H)}$ is a continuous bijection.
-It remains to show this map is a homeomorphism; we show it is closed.
-Let $\mathscr{L}_{G}(H)_{\langle k\rangle}$ denote the unitization of $\mathscr{L}_{G}(H)$, i.e.
-the tt-subcategory compactly generated by the tensor unit.
-Of course, the unitization has the same cohomology ring $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$.
-Then [San25, Corollary 10.2] asserts that the inclusion $\mathscr{L}_{G}(H)_{\langle k\rangle}\to\mathscr{L}_{G}(H)$ induces a homeomorphism $\operatorname{Spc}(\mathscr{L}_{G}(H)_{\langle k\rangle})\cong\operatorname{Spc}(\mathscr{L}_{G}(H))$.
-The comparison map factors through $\operatorname{Spc}(\mathscr{L}_{G}(H))_{\langle k\rangle}$ as follows:
-$\operatorname{Spc}(\mathscr{L}_{G}(H))\to\operatorname{Spc}(\mathscr{L}_{G}(H))_{\langle k\rangle}\to\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)).$
-Now, $\mathscr{L}_{G}(H)_{\langle k\rangle}$ is (by definition) generated by its tensor unit.
-By noetherianity of $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$, $\mathscr{L}_{G}(H)_{\langle k\rangle}$ is therefore End-finite in the sense of [Lau23, Definition 2.6], hence [Lau23, Proposition 2.7] asserts the induced map on spectra $\operatorname{Spc}(\mathscr{L}_{G}(H))_{\langle k\rangle}\to\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H))$ is a homeomorphism, thus $\operatorname{comp}_{\mathscr{L}_{G}(H)}:U(H)\to\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H))$ is a homeomorphism as well, as desired.
-Therefore, $\operatorname{comp}_{G}$ is a homeomorphism onto its image in $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$.
-It is easy to check the closed points $\mathfrak{m}_{H}$ are mapped to the corresponding maximal primes in the image as described in the theorem statement.
+$\operatorname{comp}_{\mathscr{L}_{G}(H)}\colon\operatorname{Spc}(\mathscr{L}_{G}(H))=U(H)\to\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H))).$
+Noetherianity of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ implies $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$ is Noetherian as well.
+Now, a theorem of Sanders [SAN25, Theorem 10.8] immediately $\operatorname{comp}_{\mathscr{L}_{G}(H)}$ is a homeomorphism, as desired.
+Thus, globally $\operatorname{comp}_{G}$ is a homeomorphism onto its image in $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$.
+It is easy to check the closed points $\mathscr{M}(H)$ are mapped to the corresponding maximal primes in the image as described in the theorem statement.
 ∎
@@
 Let $\mathcal{O}_{G}^{\scriptscriptstyle\bullet}$ denote the sheaf of $\mathbb{Z}$-graded rings on $\operatorname{Spc}(\mathscr{K}(G))$ obtained by sheafifying $U\mapsto\operatorname{End}^{\scriptscriptstyle\bullet}_{\mathscr{K}(G)|_{U}}(k)$.
-If $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian, $(\operatorname{Spc}(\mathscr{K}(G)),\mathcal{O}_{G}^{\scriptscriptstyle\bullet})$ is a Dirac scheme.
+If $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is Noetherian, $(\operatorname{Spc}(\mathscr{K}(G)),\mathcal{O}_{G}^{\scriptscriptstyle\bullet})$ is a Dirac scheme in the sense of [HP23].
+Of course, the pertinent question now is to deduce when Corollary 6.6 and Corollary 6.7 hold.
+7. Towards Noetherianity of twisted cohomology
+To conclude, we show the twisted cohomology ring is Noetherian in some cases, and propose a general strategy.
+Definition 7.1.
+We introduce the following terminology: a homogeneous element $f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ of the form $f\colon k\to C[s]$ for $C\in\mathcal{B}(G)$ is a minimally graded morphism.
+An effective endotrivial complex $C$ is res-onto for $H\leq G$ if it satisfies the following property: for all $s\in\mathbb{Z}$, the homomorphism induced by restriction
+$\operatorname{Res}^{G}_{H}\colon\operatorname{Hom}_{\mathscr{K}(G)}(k,C[s])\to\operatorname{Hom}_{\mathscr{K}(H)}(k,\operatorname{Res}^{G}_{H}C[s])$
+is surjective. We say $C$ is res-onto if it is res-onto for all subgroups of $G$, and a $p$-group $G$ is res-onto if all $C\in\mathcal{B}(G)$ are res-onto.
+Example 7.2.
+Being res-onto is fairly restrictive. Let’s explore some examples and non-examples. Some of these cases will be referred to later for proofs of Noetherianity as well.
+  1. (a)
+    Balmer–Gallauer compute that all elementary abelian $p$-groups are res-onto; for $u_{N}$ (c.f.
+Example 2.7), [BG25, Proposition 12.10] proves so.
+Using the classification of endotrivials for all abelian $p$-groups, a similar computation shows all abelian $p$-groups are res-onto.
+Indeed, for an abelian $p$-group AA, $\mathcal{B}(A)$ consists of endotrivials inflated from nontrivial cyclic quotients of $A$.
+Given $N\triangleleft A$, let
+    $u_{N}:=k[A/N]\to k[A/N]\to k[A/N]$
+    (assuming $[A:N]>2$) denote the associated irreducible endotrivial.
+Then one has analogous maps $a_{N},b_{N},c_{N}$ (c.f.
+Remark 4.2) in $\mathscr{K}(A)$ and an analogous computation shows $A$ is res-onto.
+More generally, for any group $G$, the inflated endotrivial $u_{N}$ is res-onto, by the same computation.
+  2. (b)
+    The quaternion group $Q_{8}$ is res-onto.
+Indeed, $\mathcal{B}(Q_{8})$ has three two-term endotrivials inflated from cyclic quotients which are res-onto, and one faithful endotrivial $C$, a truncated periodic resolution of the trivial $kQ_{8}$-module $k$ (recalling $Q_{8}$ has periodicity 4).
+Explicitly,
+    $h_{C}(H)=\begin{cases}4&\text{if }H=1;\\ 0&\text{if }H\neq 1.\end{cases}$
+    One has that $\operatorname{Res}^{G}_{H}C$ is also a truncated periodic resolution for all $H<G$.
+In particular, $\operatorname{Res}^{G}_{H}C$ after removing contractible summands has in each degree at most one indecomposable term, and it follows that $C$ is res-onto.
+Note that in this case, restriction does not induce a surjective homomorphism $\operatorname{Res}^{G}_{H}\colon\operatorname{Pic}(\mathscr{K}(G))\to\operatorname{Pic}(\mathscr{K}(H))$ for $H$ any cyclic subgroup of order 4.
+  3. (c)
+    More generally, all $p$-groups which are Dedekind, i.e., groups with every subgroup normal, are res-onto.
+All finite Dedekind groups were classified by Baer [BAE66].
+For $p$-groups, if $p>2$, then the only Dedekind groups are abelian groups.
+If $p=2$, then if $G$ is a nonabelian Dedekind $2$-group, then $G=Q_{8}\times E$ for $E$ elementary abelian.
+Every complex irreducible representation of $G$ is tensor product of an (inflated) irreducible representation of $Q_{8}$ and one of $E$.
+The complex irreducible representations of $Q_{8}$ are linear of real type or degree 2 faithful of quaternionic type, and the complex irreducible representations of $E$ are all linear of real type; therefore the irreducible representations of $G$ are either linear of real type or quaternionic of degree 2.
+Hence the real irreducible representations of $Q_{8}$ are either linear or degree 4, and explicit computation determines that the degree 4 representations are inflated from quotients $G/N\cong Q_{8}$.
+    Therefore, the elements of $\mathcal{B}(G)$ are inflated from the unique element of $\mathcal{B}(C_{2})$ and the unique faithful endotrivial of $\mathcal{B}(Q_{8})$, and it follows by the same computations as before that $G$ is res-onto.
+  4. (d)
+    On the other hand, the dihedral group $D_{8}$ has a unique faithful irreducible endotrivial (c.f. Example 2.7) which is not res-onto. We use the presentation
+    $D_{8}=\langle a,b\mid a^{4}=b^{2}=1,ab=ba^{-1}\rangle.$
+    Set $H:=\langle a^{2},b\rangle$, $K_{0}:=\langle b\rangle$, $K_{1}=\langle ab\rangle$, and $K_{2}=\langle a^{2}b\rangle$; then $K_{0},K_{2}\leq H$ and are $D_{8}$-conjugate but $K_{1}$ is nonconjugate to $K_{0},K_{2}$.
+Then we have the faithful irreducible
+    $C=kD_{8}\to k[D_{8}/K_{0}]\oplus k[D_{8}/K_{1}]\to k$
+    and it follows either by direct computation or by checking h-marks that
+    $\operatorname{Res}^{G}_{H}C\cong kH\to k[H/K_{0}]\oplus k[H/K_{2}]\to k.$
+    Moreover, one may compute that
+    $\operatorname{Res}^{G}_{H}\iota_{C}^{K_{0}}=\iota_{\operatorname{Res}^{G}_{H}C}^{K_{0}}+\iota_{\operatorname{Res}^{G}_{H}C}^{K_{2}}\text{ and }\operatorname{Res}^{G}_{H}\iota_{C}^{K_{1}}=0.$
+    In particular, $\iota_{\operatorname{Res}^{G}_{H}C}^{K_{0}}$ and $\iota_{\operatorname{Res}^{G}_{H}}^{K_{2}}$ are not in the image of restriction.
+This can be viewed as a fusion phenomenon: $K_{0}$ and $K_{2}$ are $D_{8}$-conjugate but not $H$-conjugate, so $\iota_{\operatorname{Res}^{G}_{H}C}^{K_{0}}$ and $\iota_{\operatorname{Res}^{G}_{H}}^{K_{2}}$ are not $G$-stable, and cannot be in the image of restriction.
+    Since the inflation of an irreducible endotrivial or representation is again irreducible, any $p$-group $G$ with any quotient isomorphic to $D_{8}$ is not res-onto either.
+  5. (e)
+    By an analogous conjugacy argument, one may see the same holds for any extraspecial group $G$ of order $p^{3}$ for $p$ odd.
+These are the minimal nonabelian $p$-groups, and there are two types, ++ type and - type.
+The ++ type is isomorphic to the Heisenberg group $\operatorname{He}_{p}$, the group of upper triangular matrices over $\mathbb{F}_{p}$ with 1s across the diagonal.
+For $p=2$ these are $D_{8}$ and $Q_{8}$ respectively.
+For odd primes $p$, they have the following presentations:
+    $p^{2+1}_{+}:=\langle a,b,c\mid a^{p}=b^{p}=c^{p}=1,ba=abc,ca=ac,cb=bc\rangle\cong\operatorname{He}_{p},$
+    $p^{2+1}_{-}:=\langle a,b,c\mid a^{p}=b^{p}=c,c^{p}=1,ba=abc,ca=ac,cb=bc\rangle.$
+    We sketch the details: every extraspecial group $G$ of order $p^{3}$ has a faithful complex (hence faithful real) representation $V$ by [FIT06, Part II Theorem I], as $|Z(G)|=p$.
+In fact, it via character theory that $G$ has complex-type irreducible representations of order $p$, all of which are Adams-conjugate, hence the faithful endotrivial $C_{\partial}$ has length $2p$.
+Because the dimension homomorphism
+    $\dim\colon\operatorname{RO}(G)\to\operatorname{CF}_{b}(G)\cong\operatorname{Pic}(\mathscr{K}(G))$
+    is a morphism of rational $p$-biset functors, it restricts to a morphism between faithful components, see [BOU10, Section 6.3].
+Therefore $C_{\partial}$ is faithful in the sense of [MIL24]: one has that $h_{C_{\partial}}(K)=0$ if $K$ contains a proper normal subgroup of $G$.
+On the other hand, there must exist a non-normal subgroup $H$ (necessarily of order $p$) of $G$ such that $h_{C_{\partial}}(H)>0$, otherwise $C_{\partial}$ would be a truncated projective resolution of the trivial $kG$-module $k$, which would be absurd since $G$ is not periodic.
+    There exists an elementary abelian subgroup $E\leq G$ of order $p^{2}$ containing $H$ and its $G$-conjugates.
+Indeed, if $G$ is of ++ type, then all subgroup of order $p^{2}$ suffice, as every subgroup of this order is the join of $Z(G)$ and an $H$.
+If $G$ is of - type, then there is a unique index $p$ subgroup which is elementary abelian, which is also the join of $E$ and $Z(G)$.
+We claim
+    $\operatorname{Res}^{G}_{E}\colon\operatorname{Hom}_{\mathscr{K}(G)}(k,C_{\partial})\to\operatorname{Hom}_{\mathscr{K}(E)}(k,\operatorname{Res}^{G}_{E}C_{\partial})$
+    is not surjective.
+Indeed, since $E$ is abelian, no two subgroups are $E$-conjugate.
+We have that $h_{C_{\partial}}(E)>h_{C_{\partial}}(H)$ by the classification of endotrivials for elementary abelian groups, as $h_{C_{\partial}}(G)=0$ and there are multiple conjugates of $H$ contained in $E$.
+In this case we have
+    $\operatorname{Res}^{G}_{E}\iota_{C_{\partial}}^{H}=\sum_{H^{\prime}\sim_{G}H}\iota_{\operatorname{Res}^{G}_{E}{C_{\partial}}}^{H^{\prime}},$
+    possibly up to some choice of scaling.
+Each individual $\iota_{\operatorname{Res}^{G}_{E}C_{\partial}}^{H^{\prime}}$ is not $G$-stable (even up to homotopy), hence cannot be in the image of restriction.
+Similarly, any $p$-group $G$ with extraspecial quotient is not res-onto.
+We expect that a finite $p$-group is res-onto if and only if it is Dedekind. However, as this is not our focus, we will leave working out the finer details of this expectation to the reader.
+Proposition 7.3.
+Let $C$ be an effective endotrivial complex of $kG$-modules and $H\leq G$. The following diagram commutes.
+$\operatorname{Hom}_{\mathscr{K}(G)}(k,C[s])$$\operatorname{Hom}_{\mathscr{K}(H)}(k,\operatorname{Res}^{G}_{H}C[s])$$\operatorname{Hom}_{\mathscr{K}(G)}(k,C[s]\otimes k[G/H])$$\scriptstyle{\operatorname{Res}^{G}_{H}}$$\scriptstyle{\cdot\operatorname{coaug}}\scriptstyle{\cong}$$\scriptstyle{t^{G}_{H}}$$\scriptstyle{\operatorname{id}\otimes\operatorname{aug}}$
+Here, the bottom isomorphism arises from the induction/restriction adjunction and Frobenius reciprocity (projection formula), $t^{G}_{H}$ is the transfer or trace map
+$f\mapsto\sum_{g\in[G/H]}{}^{g}f,$
+see [BEN98a, Definition 3.6.2].
+In particular, $C$ is res-onto if and only if $\operatorname{coaug}$ is surjective, that is, every element of $\operatorname{Hom}_{\mathscr{K}(G)}(k,C[s]\otimes k[G/H])$ is homotopic to a simple tensor $f\otimes\operatorname{coaug}$.
+Proof.
+Routine verification. ∎
+We now prove Noetherianity for res-onto groups. The proof involves three separate inductions.
+Theorem 7.4.
+Let $G$ be a res-onto $p$-group.
+Then $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is generated as graded $k$-algebra by the minimally graded elements $f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$.
+In particular, $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is Noetherian.
+Proof.
+Let $q\in\mathbb{N}^{\mathcal{B}(G)}$ be an element in the monoid of twists and $k(q)$ be the associated effective endotrivial.
+Then $k(q)=C_{1}\otimes\cdots\otimes C_{n}$ for some unique up to ordering collection of irreducible endotrivials $C_{1},\dots,C_{n}\in\mathcal{B}(G)$.
+We induct on $n$, $|G|$, and the number of nonzero terms in $C_{n}$ (where we assume $C_{n}$ is indecomposable); we assume the elements in the homogeneous component $\operatorname{Hom}_{\mathscr{K}(G)}(k,(C_{1}\otimes\cdots\otimes C_{n-1})[s])$ are expressible as polynomials of minimally graded elements, and that the theorem statement holds for subgroups of $G$.
+If $n=0$ or $n=1$, there is nothing to show, and similarly for $|G|=1$ for any $n$.
+The case $|G|=p$ is shown by [BG25, Lemma 12.12].
+We establish notation.
+Set $D:=C_{1}\otimes\cdots\otimes C_{n-1}[s]$ for some $s\in\mathbb{Z}$ and $C:=C_{n}$ for short.
+We may assume without loss of generality that each $C_{i}$ is indecomposable as a chain complex.
+Let $C_{i}:=M_{i}^{1}\oplus\cdots\oplus M_{i}^{a_{i}}$ be a decomposition of $C_{i}$ into indecomposable permutation $kG$-modules, with $M_{i}^{j}\cong k[G/H_{i}^{j}]$.
+Theorem 3.13 implies $C$ arises from a chain complex of free $k\Gamma_{G}$-modules, and since $C$ is also indecomposable, Corollary 3.15 implies there exists a chain complex homomorphism $\gamma^{j}_{i}\colon k\to C[i]$, unique up to scaling, with image in $C_{i}$ the unique minimal $kG$-submodule of $M_{j}$.
+Such morphisms are coaugmentation maps.
+By Corollary 3.14 and because every nontrivial real irreducible representation $V$ satisfies $V^{G}={0}$, we have $M_{0}^{1}=C_{0}=k$.
+In particular, $\gamma^{1}_{0}=\iota^{G}_{C}$ (up to scalar).
+This gives an exact triangle
+[…]
+Tensoring the above triangle with $D$ and applying the cohomological functor $\operatorname{Hom}_{\mathscr{K}(G)}(k,-)$ gives an exact sequence
+$\operatorname{Hom}_{\mathscr{K}(G)}(k,D)\xrightarrow{\cdot\iota^{G}_{C}}\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes C)\to\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes{}^{1}\hat{C}),$
+noting that the map induced by $\iota^{G}_{C}$ is identified with the map $\cdot\iota^{G}_{C}\colon\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$.
+More generally, we define the truncated complex $^{i}\hat{C}$ in the obvious way,
+$^{i}\hat{C}_{j}:=\begin{cases}C_{j}&\text{if }j\geq i;\\ 0&\text{ if}j<i,\end{cases}$
+with the obvious differentials. Each truncated complex has an associated exact triangle
+[…]
+Tensoring by $D$ and applying the cohomological functor $\operatorname{Hom}_{\mathscr{K}(G)}(k,-)$ again yields an exact sequence
+$\operatorname{Hom}_{\mathscr{K}(G)}\left(k,D\otimes\bigoplus_{j=1}^{a_{i-1}}M_{i-1}^{j}[i-1]\right)\to\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes{}^{i-1}\hat{C})\to\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes{}^{i}\hat{C}).$
+We claim that every $^{i}\hat{C}$ (including $C={}^{0}\hat{C}$) satisfies the following property: all elements of $\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes{}^{i}\hat{C})$ can be expressed as polynomials in minimally graded morphisms $k\to C_{i}[s^{\prime}]$ and the $\gamma$ morphisms previously described.
+Note each morphism $\gamma_{l}^{j}$ is also well-defined on $^{i}\hat{C}$ when $l\geq i$ because $C$ arises from a complex of free $k\Gamma_{G}$-modules, and therefore so do its truncations.
+We proceed with the base case of $^{n}\hat{C}$ and $^{n-1}\hat{C}$.
+Let $N\trianglelefteq G$ be the kernel of the irreducible representation associated to $D$, then 4.1 asserts $C_{n}=k[G/N]$, hence $^{n}\hat{C}=k[G/N][n]$.
+In particular, the $n$th and final exact sequence is of the form
+$\operatorname{Hom}_{\mathscr{K}(G)}\left(k,D\otimes\bigoplus_{i=1}^{a_{n-1}}M^{i}_{n-1}[n-1]\right)\to\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes{}^{n-1}\hat{C})\to\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes k[G/N][n]).$
+Let $f\in\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes{}^{n-1}\hat{C})$ and set $f^{\prime}$ equal to its image in $\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes k[G/N][n])$.
+By adjunction, $\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes k[G/N][n])\cong\operatorname{Hom}_{\mathscr{K}(N)}(k,\operatorname{Res}^{G}_{N}D[n])$, therefore the image of $f^{\prime}$ under adjunction is polynomial in minimally graded morphisms.
+By 7.3, $f^{\prime}$ is also polynomial in minimally graded morphisms, as $D$ is res-onto by induction, so $f^{\prime}=f^{\prime\prime}\otimes\operatorname{coaug}=f^{\prime\prime}\otimes\gamma_{n}^{1}$ (up to a scalar) for some $f^{\prime\prime}\in\operatorname{Hom}_{\mathscr{K}(G)}(k,D[s])$.
+Now $f^{\prime\prime}\gamma_{n}^{1}\in\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes{}^{n-1}\hat{C})$, and by computation, $f$ and $f^{\prime\prime}\gamma_{n}^{1}$ have same image in $\operatorname{Hom}_{\mathscr{K}(G)}(k,D[s]\otimes k[G/N])$.
+Hence $f-f^{\prime\prime}\cdot\gamma_{n}^{1}$ is in the image of the left-side map.
+Again the following isomorphism holds by adjunction
+$\operatorname{Hom}_{\mathscr{K}(G)}\left(k,D\otimes\bigoplus_{j=1}^{a_{n-1}}M^{j}_{n-1}[n-1]\right)\cong\bigoplus_{j=1}^{a_{n-1}}\operatorname{Hom}_{\mathscr{K}(H_{n-1}^{j})}\left(k,\operatorname{Res}^{G}_{H_{n-1}^{j}}D[n-1]\right),$
+so again we may use induction and 7.3 as before to find a collection of morphisms $(f_{1}^{\prime\prime\prime},\dots,f_{a_{n-1}}^{\prime\prime\prime})$, with $f_{i}^{\prime\prime\prime}\in\operatorname{Hom}_{\mathscr{K}(G)}(k,D[n-1])$, such that $f_{1}^{\prime\prime\prime}\gamma_{n-1}^{1}+\cdots+f_{a_{n-1}}^{\prime\prime\prime}\gamma_{n-1}^{a_{n-1}}=f-f^{\prime\prime}\cdot\gamma_{n}^{1}$ in $\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes{}^{n-1}\hat{C})$, as desired.
+[…]
+$\operatorname{Hom}_{\mathscr{K}(G)}\left(k,D\otimes\bigoplus_{i=1}^{a_{i-1}}M^{j}_{i-1}[i-1]\right)\cong\bigoplus_{j=1}^{a_{i-1}}\operatorname{Hom}_{\mathscr{K}(H_{i-1}^{j})}\left(k,\operatorname{Res}^{G}_{H_{i-1}^{j}}D[i-1]\right).$
+Then induction and 7.3 allow us to find a collection of morphisms $(f_{1}^{\prime\prime},\dots,f_{a_{i-1}}^{\prime\prime})$ such that $f_{1}^{\prime\prime}\gamma_{i-1}^{1}+\cdots+f_{a_{i-1}}^{\prime\prime}\gamma_{i-1}^{a_{i-1}}=f-f^{\prime}$.
+Since all terms in this equality besides $f$ in this equality are polynomial in minimal elements by induction, we have shown that $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is generated by its minimally graded elements.
+Since the finite collection of $k$-vector spaces associated to each minimal grading is finite-dimensional, selecting a $k$-basis for each vector space obtains a finite list of generators of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$, as desired.
+∎
+7.1. Noetherianity for groups of order $p^{3}$
+Res-onto-ness is not a necessary condition for Noetherianity.
+We prove $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is Noetherian for all remaining groups of order $p^{3}$.
+This consists of all extraspecial groups of order $p^{3}$ and the minimal nonabelian groups (except for $Q_{8}$, which is res-onto).
+A difference here is that for extraspecial groups $G$, $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is no longer generated by minimally graded elements.
+We split the proofs for $D_{8}$ and for odd order groups, since the odd order groups have an additional step.
+Definition 7.5.
+Recall that $|\mathcal{B}(D_{8})|=4$, with three endotrivials inflated from the 3 maximal proper subgroups of $D_{8}$, and one faithful endotrivial $C_{\partial}$, which fails to be res-onto for the subgroups of $D_{8}$ isomorphic to $V_{4}=C_{2}\times C_{2}$, by Example 7.2.
+We say a homogeneous element $f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(D_{8})$ is $D_{8}$-minimal if either $f$ is minimally graded or if it is of the form $f\colon k\to C_{\partial}^{\otimes 2}[-2]$.
+Theorem 7.6.
+Set $G=D_{8}$.
+Then $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is generated as graded $k$-algebra by the $D_{8}$-minimal morphisms.
+In particular, $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is Noetherian.
+Proof.
+Explicit computation shows that the only obstruction to $G$ being res-onto comes from the faithful irreducible endotrivial (see Example 7.2), which we label by $C_{f}$, when restricted to the two subgroups of $D_{8}$ isomorphic to $V_{4}$.
+Therefore, proceeding by induction on $n$ and $|G|$ as before, the only new case to consider is if, in the decomposition $C_{1}\otimes\cdots\otimes C_{n}$, $C_{n}=k[G/H]\to k$ with the subgroup $H\leq G$ isomorphic to $V_{4}$.
+We set the same notation as before: let $s\in\mathbb{Z}$, set $D:=C_{1}\otimes\cdots\otimes C_{n-1}[s]$ and $C:=k[G/H]\to k$ for $H\cong V_{4}$.
+We again have an exact triangle
+[…]
+Tensoring with $D$ and applying $\operatorname{Hom}_{\mathscr{K}(G)}(k,-)$ gives the exact sequence
+$\operatorname{Hom}_{\mathscr{K}(G)}(k,D)$ $\xrightarrow{\cdot\iota_{C}^{G}}\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes C)$
+$\xrightarrow{\,}\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes k[G/H][1])$
+$\xrightarrow{\operatorname{id}\otimes\operatorname{aug}}\operatorname{Hom}_{\mathscr{K}(G)}(k,D[1]),$
+and the final three terms identify under adjunction and 7.3 as
+$\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes C)\xrightarrow{\phi}\operatorname{Hom}_{\mathscr{K}(H)}(k,\operatorname{Res}^{G}_{H}D[1])\xrightarrow{t^{G}_{H}}\operatorname{Hom}_{\mathscr{K}(G)}(k,D[1]).$
+Let $f\in\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes C)$ and let $f^{\prime}$ denote its image in $\operatorname{Hom}_{\mathscr{K}(H)}(k,\operatorname{Res}^{G}_{H}D[1])$.
+Again by induction, we have that $f^{\prime}$ can be expressed as a polynomial in morphisms $k\to\operatorname{Res}^{G}_{H}C_{i}[s^{\prime}]$.
+However, because $D_{8}$ is not res-onto, we cannot automatically lift the same polynomial to one in $\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes C)$ as before.
+Since the sequence is exact, we have $\operatorname{im}(\phi)=\ker(\operatorname{id}\otimes\operatorname{aug})$.
+Since $[G:H]=p=2$, 6.1 and 7.3 demonstrate that in fact, $\ker(t^{G}_{H})=\operatorname{im}(\operatorname{Res}^{G}_{H})$ with $\operatorname{im}(\operatorname{Res}^{G}_{H})$ equivalently the $G/H$-stable morphisms.
+Equivalently, $\ker(\operatorname{id}\otimes\operatorname{aug})=\operatorname{im}(\cdot\operatorname{coaug})$.
+In particular, the image of $\phi$ consists of $G$-stable morphisms, i.e., the ring of invariants.
+By induction, $f^{\prime}$ is a $G$-stable polynomial in minimally graded elements.
+Note that while $\operatorname{Res}^{G}_{H}C_{\partial}$ is not irreducible, and therefore the morphisms into it are not minimally graded, any minimally graded morphism (in $H$) in the polynomial representation of $f^{\prime}$ is necessarily multiplied with another to produce a morphism $k\to\operatorname{Res}^{G}_{H}C_{\partial}[s^{\prime}]$.
+In this way, $p$ is also polynomial with respect to homogeneous morphisms with grading in $\operatorname{Res}^{G}_{H}\mathcal{B}(G)\subset\operatorname{Pic}(\mathscr{K}(H))$.
+We refer to this as being generated by res-minimally graded elements.
+In order to lift the $f^{\prime}$ from $H$ to $G$, we moreover need to check that in fact, this polynomial can be written such that each homogeneous term is $G$-stable.
+In other words, we require that the ring of invariants $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H)^{G}$ admits a $G$-stable generating set.
+The key insight here is that $f^{\prime}$ is $G$-stable if and only if it can be written as a polynomial in $G$-stable res-minimal elements and symmetric polynomials of non-$G$-stable elements.
+There are only 2 (up to scaling and homotopy) res-minimally graded elements which do not belong, as described in Example 7.2.
+We denote these by $\iota_{1},\iota_{2}\colon k\to\operatorname{Res}^{G}_{H}C_{\partial}[-1]$.
+The symmetric polynomials $\iota_{1}+\iota_{2}$ and $\iota_{1}\iota_{2}\colon k\to C_{\partial}^{\otimes 2}[-2]$ are of course $G$-stable (as the action of $G/H$ permutes the two), hence both belong to $\operatorname{im}(\operatorname{Res}^{G}_{H})$ by 6.1.
+Note that $\iota_{1}\iota_{2}$ has a lift which is $D_{8}$-minimal but is not minimally graded.
+Thus, $f^{\prime}$ can be “copied” into a polynomial $f^{\prime\prime}\in\operatorname{Hom}_{\mathscr{K}(G)}(k,D)$ in $D_{8}$-minimal morphisms satisfying that $f-\iota^{1}_{C}f^{\prime\prime}\in\ker(\phi)$, hence $f-\iota^{1}_{C}f^{\prime\prime}\in\operatorname{im}(\cdot\iota^{G}_{C})$ as before.
+It follows as before that we obtain that $f-\iota^{1}_{C}f^{\prime\prime}=\iota^{G}_{C}f^{\prime\prime\prime}$ for some $f^{\prime\prime\prime}\in\operatorname{Hom}_{\mathscr{K}(G)}(k,D)$, and by induction, $f^{\prime\prime\prime}$ is also polynomial in $D_{8}$-minimal elements.
+Thus we have $f=\iota^{1}_{C}f^{\prime\prime}+\iota^{G}_{C}f^{\prime\prime\prime}$, and we conclude $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is generated by the $D_{8}$-minimally graded morphisms.
+Since the (finite) collection of $k$-vector spaces associated to $D_{8}$-minimality is finite-dimensional, selecting a $k$-basis for each vector space obtains a finite list of generators of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$, as desired.
+∎
+Finally, we extend this proof for all extraspecial groups of order $p^{3}$ with $p$ odd.
+The proof follows a similar logic, but has one additional step.
+The subgroup lattices of $3^{1+2}_{+}$ and $3^{1+2}_{-}$ are given for reference.
+$C_{1}$$C_{3}$$C_{3}$$C_{3}$$C_{3}$$C_{3}$$C_{3}^{2}$$C_{3}^{2}$$C_{3}^{2}$$C_{3}^{2}$$^{1+2}_{+}33$33
+Why hello there!
+$C_{1}$$C_{3}$$C_{3}$$C_{9}$$C_{9}$$C_{9}$$C_{3}^{2}$$3_{-}^{1+2}$3
+Figure 1.
+Subgroup lattice of $3^{1+2}_{+}$ and $3^{1+2}_{-}$ respectively, with conjugate subgroups merged.
+Purple indicates normal, blue indicates characteristic, subscript counts number of conjugates if non-normal.
+Diagram obtained from [25].
+Definition 7.7.
+Let $G$ be an extraspecial group of order $p^{3}$.
+Denote the unique faithful irreducible endotrivial for $kG$ by $C_{\partial}$ (see Example 7.2 for justification of their existence).
+We say a homogeneous element $f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is $G$-minimal (for the specified extraspecial group $G$) if either $f$ is minimally graded or if it is of the form $f\colon k\to C_{\partial}^{\otimes n}[s]$ for $n\in\{2,\dots,p\}$ and $s\in\mathbb{Z}$.
+Theorem 7.8.
+Let $p>2$ and $G$ be an extraspecial group of order $p^{3}$.
+Then $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is generated as graded $k$-algebra by its $G$-minimal morphisms.
+In particular, $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is Noetherian.
+Proof.
+The proof follows similarly to Theorem 7.6, therefore for brevity, we omit a few repetitive details.
+Since $p>2$, the irreducible endotrivials inflated from $C_{p}$ have length 3.
+We have again that $\mathcal{B}(G)$ consists of those inflated endotrivials and the faithful endotrivial $C_{\partial}$ of length $2p+1$ described in Example 7.2.
+The inflated endotrivials are res-onto, and $C_{\partial}$ is res-onto for all groups of order $p$.
+With the same induction setup as before, setting $D=C_{1}\otimes\cdots\otimes C_{n-1}[s]$ and $C$ an irreducible effective endotrivial, the only case to evaluate is $C=k[G/H]\to k[G/H]\to k$ for $H\leq G$ satisfying $[G:H]=p$.
+We again have an exact triangle, but this time, we truncate from the left.
+[…]
+Like the previous case, we have an exact sequence associated to $\hat{C}$,
+$\operatorname{Hom}_{\mathscr{K}(G)}(k,D)\to\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes\hat{C})\xrightarrow{\phi}\operatorname{Hom}_{\mathscr{K}(H)}(k,\operatorname{Res}^{G}_{H}D[1])\xrightarrow{\cdot t^{G}_{H}}\operatorname{Hom}_{\mathscr{K}(G)}(k,D[1]).$
+[…]
+Now we consider the first exact sequence associated to $C$,
+[…]
+$\xrightarrow{\phi}\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes k[G/H][2])$
+$\xrightarrow{\operatorname{id}\otimes(d_{2}\circ-)}\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes\hat{C}[2]).$
+Once again, we must show $\operatorname{im}(\phi)$ is copyable.
+In this case, it is easier to show without the induction/restriction adjunction.
+We consider $d_{2}\colon k[G/H]\to k[G/H]$, which is induced by the assignment $1H\mapsto 1H-gH$ for some $g\not\in H$, and satisfies
+$\ker(d_{2})=\left\langle\sum_{g\in[G/H]}gH\right\rangle\cong k.$
+[…]
+Corollary 7.9.
+Let $G$ be a $p$-group of order at most $p^{3}$. Then $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is Noetherian.
+Proof.
+All $p$-groups of order $p^{3}$ are either abelian or extraspecial, and of order $p^{2}$ or $p$ are abelian.
+This is well-known, see e.g.
+[HUN74, Section 2.6] or [23] for chronological history.
+The result follows from combining Theorem 7.4, Theorem 7.6, and Theorem 7.8.
+∎
 Remark 7.10.
-We informally say some words about noetherianity to close, to hopefully conjure hope within the reader that for any finite $p$-group $G$, $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian.
-First, one can show $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian when $G$ satisfies the following property: the indecomposable representative of each irreducible endotrivial $C$ in $\mathcal{B}(G)$ has at most one indecomposable module in every homological degree.
-This holds for instance (and possibly only when) $G$ is abelian, where the endotrivials correspond to subgroups $N\leq G$ such that $G/N$ is cyclic (see [Mil24, Section 6]).
-With this property, the proof follows essentially identically as [BG25, Lemma 12.12], with that the maps $\iota^{1}_{C}$ and $\iota^{G}_{C}$ replacing the maps $b_{N}$ and $a_{N}$ respectively for each irreducible endotrivial $C\in\mathcal{B}(G)$.
-One crux of the matter is that if an irreducible endotrivial has a homological degree with two or more indecomposables, the exact sequence [BG25, 12.14] sends a term $f\in\operatorname{Hom}_{\mathscr{K}(G)}(k,v\otimes u^{\otimes q})$ to a direct sum of morphisms $(f_{1},\dots,f_{n})$ with $f_{i}\in\operatorname{Hom}_{\mathscr{K}(H_{i})}(k,\operatorname{Res}^{G}_{H_{i}}(v)[q])$, where the subgroups $H_{i}\leq G$ are the corresponding stabilizers of the indecomposable permutation modules.
-It is unclear if one can inductively lift this sum of morphisms to a morphism in $\operatorname{Hom}_{\mathscr{K}(G)}(1,v\otimes u^{\otimes q})$, polynomial in morphisms $k\to C[s]$ with $C\in\mathcal{B}(G)$, such that its image in every $\operatorname{Hom}_{\mathscr{K}(H_{i})}(k,\operatorname{Res}^{G}_{H_{i}}(v)[q])$ corresponds to the image of $f$.
-Additionally, the lengths of the irreducible endotrivials $C$ can be arbitrarily long, but an inductive argument may be used to resolve this issue.
-A potential roadmap to proving noetherianity modulo nilpotents is to verify that the image of the map $\prod\hat{\Psi}:\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\prod_{H\in s_{p}(G)/G}\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H})$ is finite, as the kernel of $\prod\hat{\Psi}$ is nilpotent (Proposition 5.6).
-One can further reduce this to the elementary abelian case by postcomposing by restriction to all elementary abelian subquotients of $G$, as the restriction $\operatorname{H}^{\scriptscriptstyle\bullet}(G)\to\prod_{E\in\operatorname{elemab}}(G)\operatorname{H}^{\scriptscriptstyle\bullet}(E)$ detects nilpotents as well (see [Ben98b, Proposition 5.2.2]).
-It is not too unreasonable to expect this to hold - for instance [Eve61, Theorem 7.1] states that for any subgroup $H\leq G$, $\operatorname{H}^{\scriptscriptstyle\bullet}(H)$ is finite in $\operatorname{H}^{\scriptscriptstyle\bullet}(G)$ via restriction.
-Hence $\operatorname{H}^{\scriptscriptstyle\bullet}(H)^{G}$ (i.e.
-the image of restriction, see [Ben98a, Proposition 3.8.2]), is finite in $\operatorname{H}^{\scriptscriptstyle\bullet}(G)$.
-For this reason, it is quite pertinent and interesting to determine the image of $\hat{\Psi}^{H}:\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H})$ for any subgroup $H\leq G$.
-We note for instance that one can show the image of $\prod\hat{\Psi}$ is finite when $p=2$ and $G=Q_{8}$ exploiting the fact that every subgroup of $Q_{8}$ is normal.
-However, it is not clear if an analogue of [Bal10a, Theorem 7.3] under the hypothesis that the cohomology ring is noetherian modulo nilpotents can be utilized.
+The identification $\ker(\operatorname{Res}^{G}_{H})=\operatorname{im}(t^{G}_{H})$ does some heavy lifting in the two proofs: it is a seemingly necessary condition to perform the ‘copying’ argument.
+This does not hold in full generality for subgroups of a $p$-group for morphisms of chain complexes of index at least $p^{2}$; simple counterexamples are easy to construct.
+However, if one considers the irreducible endotrivials $C\in\mathcal{B}(D_{2^{n}})$ with $n\geq 3$ (in particular, the three-term complexes), surprisingly, the same equality can be shown to hold for all subgroups, not just those of index 2.
+This seems to be a nice quirk of working up to homotopy.
+It is not clear to us if such an equality holds for all effective endotrivials for arbitrary groups, or if such an equality is necessary for Noetherianity of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ to hold.
+More generally, the proposition we’d like to show is as follows.
+Let $C$ be an effective endotrivial with $n+1$ terms, let $d_{i}$ be a transition map for $0<i\leq n$, let $\hat{C}$ be formed from $C$ by truncation with $\hat{C}_{j}=0$ for $j>i$, and let $\phi$ denote the obvious morphism $\hat{C}\to C_{i}[i]$.
+We write $C_{i}=\bigoplus_{l=1}^{m}k[G/H_{l}]$.
+Then we expect the induced homomorphism
+$\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes\hat{C})\xrightarrow{\phi^{*}}\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes C_{i}[i])\cong\bigoplus_{l=1}^{m}\operatorname{Hom}_{\mathscr{K}(H_{l})}(k,\operatorname{Res}^{G}_{H_{l}}D[i])$
+has image consisting of $G$-stable elements. Such a result need not occur for all complexes arising from the orbit category; this appears to be a special property of effective endotrivials.
+If indeed this holds, a combination of the recursive argument in Theorem 7.4 on the length of the endotrivial $C$, the ‘copying’ argument in Theorem 7.8 relying on $G$-stability, and Hilbert’s basis theorem to construct a finite $G$-stable generating set, should suffice to prove Noetherianity.
+We leave the full argument to the reader.
+Unfortunately, despite the progress we have made elucidating the structure of effective endotrivials, such a result at present still appears far out of reach.